### v1

# Online Fair Allocations with Binary Valuations and Beyond

Abstract
In an online fair allocation problem, a sequence of indivisible items arrives online and needs to be allocated to offline agents immediately and irrevocably. In our paper, we study the online allocation of either goods or chores. We adapt popular fairness notions, including envy-freeness up to one item (EF1) and maximin share fairness (MMS), and use utilitarian social welfare (USW) to measure efficiency. For both settings of items, we present a series of positive results regarding the existence of fair and efficient allocations with widely studied classes of binary and bivalued valuation/cost functions. Additionally, we complement our results by constructing some counterexamples to establish our results as among the best guarantees possible.
1 Introduction
How to allocate a set of indivisible items fairly among a set of agents is a hot topic in various research communities, including algorithmic game theory and artificial intelligence. For decades, there have been tremendous ongoing efforts to design efficient algorithms for finding fair allocations regarding envy-free up to one item (EF1), maximin share fairness (MMS), and some other fairness; detailed progress and results can be found in [1]. Most prior work focuses on offline setting, where the set of items is known upfront. In many practical situations, decisions about the allocation of items must be made online without or with limited knowledge about the future. A paradigmatic example is the food banks problem [2]. In this scenario, food comes online without prior knowledge of future food, and the food is perishable; it must be allocated upon arrival to one of the agents.
In this paper, we study the online problem of allocating a set of items that arrive one by one to a set of agents that are known in advance. When an item arrives, its values to the agents are revealed, and an immediate and irrevocable decision must be made about which agent to assign the item. There has been a line of work about online fair allocation, where they consider either additive binary valuations [3] or normalized valuations [4, 5, 6]. Few non-trivial results about the approximation fairness guarantee of online goods allocation beyond additive binary valuations and online chores allocation are known when there is no prior information about the items. In our paper, we would like to fill this gap and consider the online allocation model without any information about future items, where we discuss the goods and chores settings, respectively. Our primary research question is
Can we design online algorithms that are fair and efficient, for general valuation functions beyond additive binary, in the online allocation problem?
1.1 Our Contributions
Valuations $N$ Fairness/ Efficiency LB UB
[…]
USW $\times$ (Prop 2)
(a) Goods setting
Costs $N$ Fairness/ Efficiency LB UB
[…]
(b) Chores setting
Table 1: A summary of our main results. "$\times$" indicates inapproximability, $N$ indicates the number of agents. "LB" indicates lower bound, and "UB" indicates upper bound.
There are $n$ agents, and each indivisible item comes online. We study the goods and chores settings, and introduce non-wastefulness (NW) and completeness constraints to the goods and chores settings, respectively. For different classes of valuation or cost functions, including submodular/supermodular binary, additive bi-valued, and so on, our main results are summarized in Table 1.
In the goods setting, we first show the inapproximability of EF1 and MMS with general additive valuations, even for only two agents with additive tri-valued valuations. This indicates that we can only focus our efforts on restricted valuation functions.
  • For submodular binary valuation functions, we propose the Marginal-Greedy Algorithm (Algorithm 2) to compute allocations that simultaneously satisfy NW, $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW. Further, we show that no deterministic algorithm can guarantee an NW allocation that is $\alpha$-EF1, $\alpha$-MMS, or $\alpha$-max-USW for any $\alpha>\frac{1}{2}$.
  • For two agents with additive bi-valued valuation functions, we design the Adapted Envy-Graph Procedures (Algorithm 3) to compute NW, $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS allocations. Further, we show that no deterministic algorithm can guarantee a NW allocation that is $\alpha$-EF1 for any $\alpha>\frac{1}{2}$ or $\beta$-MMS for any $\beta>\frac{1}{3}$, and no deterministic algorithm can achieve NW, $\alpha$-EF1 or $\alpha$-MMS, and $\gamma$-max-USW for any $\alpha,\gamma>0$.
  • When one agent has the additive bi-valued valuation function and the remaining agents have identical additive binary valuation functions, we design the Adapted-Picking Algorithm (Algorithm 6) to compute an allocation that simultaneously satisfies NW, EF1, and MMS. Further, we show that no deterministic algorithm can compute an NW allocation that is EF1 or MMS guarantee $\alpha$-max-USW for any $\alpha>0$.
In the chores setting, we first demonstrate the inapproximability of EF1 for additive tri-valued cost functions, which hinders our ability to explore more general cost functions. Next, we investigate the restricted cost functions, such as additive binary cost functions.
  • For additive binary cost functions, we propose the Compelled-Greedy Algorithm (Algorithm 7) to compute a complete allocation that simultaneously satisfies EF1, MMS, and min-USC.
  • For supermodular binary cost functions, we prove that no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1, $\alpha$-MMS, or $\alpha$-min-USC for any $\alpha\geq 1$.
  • When two agents have additive bi-valued cost functions, we design the Adapted Chores Envy-Graph Procedures (Algorithm 8) to compute an allocation that satisfies completeness, $2$-EF1, and $\frac{5}{3}$-MMS. Further, we show that no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any $\alpha<2$ or $\beta$-MMS for any $\beta<\frac{3}{2}$, and no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 or $\beta$-MMS, and $\gamma$-min-USC for any $\alpha,\gamma\geq 1$ and $\beta<2$.
  • When one agent has the additive bi-valued cost function and the remaining agents have additive binary cost functions, we design the Adapted-Chores-Picking Algorithm (Algorithm 11) to compute an allocation that simultaneously satisfies completeness, EF1 and MMS. Further, no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1 or MMS guarantee $\beta$-min-USC for any $\alpha,\beta\geq 1$.
1.2 Related Work
Online Fair Allocation.
In the majority of the literature, there are two kinds of online arrivals: The former kind is that agents arrive over time [7, 8, 9, 10], and the latter is that items arrive over time, like [3], which is also the topic of our paper. Our work is mainly related to the study of [3] and [11], where items arrive one by one without any future information known in advance. Aleksandrov et al. [3] consider the food bank problem when all agents only have the additive binary valuation functions and design two simple mechanisms to guarantee ex-ante envy-freeness. In our paper, we study beyond additive binary valuation functions in the goods setting. Hosseini et al. [11] study the class fairness in online matching for indivisible and divisible items, and present a series of positive results. The model of indivisible items in their paper can be seen as a special case of our model with submodular binary valuation functions in our paper, which will be discussed later. In other words, we study a more general setting and derive the approximation results that strengthen their results.
Some work studies about online fair allocation with some prior information of future items, like the normalized valuations [5, 6], the rank of a new arrival item [12], and other information [13, 14, 15]. Zhou et al. [5] study the online allocation of indivisible goods or chores with normalized valuations and design several algorithms only to guarantee the final allocation that approximates maximin share, which is in contrast to our setting, where our online algorithm ensures the approximation ratio in any round. Benade et al. [6] study the trade-off between fairness and efficiency in allocating the indivisible items under different types of adversaries with normalized valuations. Benade et al. [12] study the allocation of indivisible items with partial information, i.e., the rank of a new arrival item with respect to the past items under different assumptions of agents’ valuation distributions. He et al. [13] study the minimum adjustments (the reallocation of previously allocated items) to ensure envy-freeness up to one item for online indivisible items allocation with or without the information where the values of future items are observed. Elkind et al. [14] and Cookson et al.[15] study the temporal allocation of online indivisible items where the arriving information of items, like values of items, is known and give several positive results. With some known information, we also explore online fair allocation, and the relevant results are presented in Appendix B.
Offline Fair Allocation.
When agents have additive valuation functions, there is a rich body of literature that discusses EF1 or MMS. Please see the survey of [1] for an overview. Here, we list some papers most relevant to our work about offline fair division with submodular binary valuation functions or supermodular binary cost functions. When agents have submodular binary valuations in the goods setting, Benabbou et al. [16] study the compatibility of fairness and efficiency and show that a maximizing utilitarian social welfare allocation that is EF1 can be efficiently computed, and Barman et al. [17] consider the fairness notion - MMS and prove that a maximizing utilitarian social welfare allocation that satisfies MMS can be computed in polynomial time. When agents have supermodular binary cost functions in the chores setting, Barman et al. [18] show that an EF1 or MMS allocation that minimizes the social cost can be computed in polynomial time.
2 Preliminaries
For each natural number $s\in\mathbb{N}$, let $[s]=\{1,\ldots,s\}$. Let $T=\{e_{1},\ldots,e_{t}\}$ be the set of indivisible items arriving one by one and $N=\{1,\ldots,n\}$ be the set of offline agents. When an online item $e_{t}$ arrives, its incident values or costs to agents are revealed. When items are goods, each agent is endowed with a valuation function $v_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$. When items are chores, each agent is endowed with a cost function $c_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$. We assume that $v_{i}(\emptyset)=0$ ($c_{i}(\emptyset)=0$), and monotone, $v_{i}(X)\leq v_{i}(Y)$ ($c_{i}(X)\leq c_{i}(Y)$) for all $X\subseteq Y\subseteq M$. For any item $e\in T$ and any subset $X\subseteq T$, we denote by $\Delta^{i}_{X}(e)=v_{i}(X\cup\{e\})-v_{i}(X)$ or $c_{i}(X\cup\{e\})-c_{i}(X)$ the marginal value or cost of an item $e\in M$ in the set $X$. $v_{i}$ is additive if the values of any set of items $S$ is the sum of the value of each item in $S$, that is, $v_{i}(X)=\sum_{e\in X}v_{i}(e)$. $v_{i}$ is submodular if $v_{i}(X)+v_{i}(Y)\geq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$. $v_{i}$ is supermodular if $v_{i}(X)+v_{i}(Y)\leq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$. There are similar arguments for $c_{i}$, and we omit them. Let $T^{k}$ be the set of items that have arrived by round $k\in[t]$. We denote an allocation by $A^{k}=(A_{1}^{k},\ldots,A_{n}^{k})$, where $A_{i}^{k}\subseteq T^{k}$ is the bundle of items that are allocated to agent $i$, and $A_{i}^{k}\cap A_{j}^{k}=\emptyset$ for any $i,j\in N$. Collectively, $(T,N,(v_{i})_{i\in N})$ forms an instance of online goods fair allocation and $(T,N,(c_{i})_{i\in N})$ forms an instance of online chores fair allocation. For a singleton set $\{e\}$, we will use $v_{i}(e)$ ($c_{i}(e)$) as a shorthand for $v_{i}(\{e\})$ ($c_{i}(\{e\})$).
Deterministic online algorithms
When one item arrives, the values (costs) of this item for all agents are revealed. The online algorithm must make an immediate and irrevocable decision to allocate the item to one agent or throw it away deterministically.
Our goal is to design the deterministic algorithms to compute allocations guaranteeing the desired approximation fairness or efficiency (approximate EF1, MMS, USW, USC) (Definitions 1, 2, 3, 8, 9, 10) and constraint (non-wastefulness for goods (Definition 4) and completeness for chores (Definition 7)) at the end of each round. In contrast, our impossibility results will hold even if the desired guarantees are required to hold only at the end. Let us demonstrate the deterministic online allocation procedure with the following example.
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $6$ $4$ $5$ 9 3
agent 2 $5$ $8$ $10$ 2 5
agent 3 $12$ 6 6 8 4
Table 2: Example of The Deterministic Online Allocation.
Example.
Let us consider the following example to demonstrate the deterministic online allocation procedure. Suppose that there are three agents who have additive valuation functions, and the arriving items are goods. The value of each item is shown in Table 2. Now we have a deterministic online algorithm named $\mathsf{OALG}$. In round $t=1$, $\mathsf{OALG}$ allocates $e_{1}$ to agent 1 after the values of this item are revealed. For allocation $A^{1}$, it is EF1, MMS and $\frac{1}{2}$-max-USW. In round $t=2$, by $\mathsf{OALG}$, $e_{2}$ is allocated to agent 2, and allocation $A^{2}$ is EF1, MMS, and $\frac{7}{10}$-max-USW. In round $t=3$, by $\mathsf{OALG}$, agent 3 picks item $e_{3}$ and allocation $A^{3}$ is EF1, MMS, and $\frac{2}{3}$-max-USW. In round $t=4$, by $\mathsf{OALG}$, agent 1 picks item $e_{4}$ and allocation $A^{4}$ is $\frac{3}{4}$-EF1, $\frac{3}{4}$-MMS, and $\frac{29}{39}$-max-USW. In round $t=5$, by $\mathsf{OALG}$, agent 1 picks item $e_{5}$ and allocation $A^{5}$ is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{8}{11}$-max-USW. Then, no further item will arrive and we say that $\mathsf{OALG}$ achieves $\min\{\frac{1}{2},\frac{3}{4},1\}=\frac{1}{2}$-EF1, $\min\{\frac{1}{2},\frac{3}{4},1\}=\frac{1}{2}$-MMS, and $\min\{\frac{1}{2},\frac{2}{3},\frac{7}{10},\frac{8}{11},\frac{29}{39}\}=\frac{1}{2}$-max-USW.
3 Allocation of Goods
In this section, we focus on the setting of goods. Here are some definitions of fairness and efficiency notions.
Definition 1 (Envy-freeness up to One Good).
For any constant $\alpha\in[0,1]$, an allocation $A^{k}$ is $\alpha$-approximate envy-free up to one good ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i^{\prime}}^{k}\neq\emptyset$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot v_{i}(A_{i^{\prime}}^{k}\setminus\{e\})$ for some $e\in A_{i^{\prime}}^{k}$.
Definition 2 (Maximin Share Fairness).
For any agent $i$, her maximin share $\mathsf{MMS}_{i}^{k}$ by round $t$ is defined as:
$\mathsf{MMS}_{i}^{k}=\max_{B^{k}\in\Pi_{n}(T^{k})}\min_{i^{\prime}\in N}v_{i}(B_{i^{\prime}}^{k}),$
where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$. For any constant $\alpha\in[0,1]$, an allocation $A^{k}$ is $\alpha$-approximate maximin share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
Definition 3 (Utilitarian Social Welfare).
The utilitarian social welfare $USW$ of allocation $A^{k}$ is given by $USW(A^{k})=\sum_{i\in N}v_{i}(A_{i}^{k})$. For any constant $\alpha\in[0,1]$, an allocation $A^{k}$ is $\alpha$-max-USW if $USW(A^{k})\geq\alpha\cdot\max_{B^{k}\in\Pi_{n}(T^{k})}USW(B^{k})$ holds, where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
In addition to the above definitions, we also pay attention to the following restriction, which is added in the goods setting.
Definition 4 (Non-Wastefulness).
An allocation $A^{k}$ is non-wasteful (NW) if (1) $\Delta_{A_{i}^{k}\setminus\{e\}}(e)>0$ for any $i\in N$ and any good $e\in A^{k}_{i}$. (2) $\Delta_{A_{i}^{k}}({e}^{{}^{\prime}})=0$ for any discarded good $e^{\prime}\in T^{k}$.
When agents have additive valuation functions, for any arriving good $e$, as long as it is allocated to one agent who values it positively, an NW allocation can be found. Next, We explore the existence of submodular valuation functions.
Claim 1.
For submodular binary valuation functions, an NW allocation always exists.
Proof.
Given an instance $(T,N,(v_{i})_{i\in N})$, Algorithm 1 can compute an NW allocation in polynomial time. When item $e_{k}$ ($k\in[t]$) arrives, if there exists an agent $j\in N$ such that $\Delta_{e_{k}}^{j}(A_{j}^{k-1})=1$, $e_{k}$ will be allocated. We assume that $e_{k}$ is allocated to agent $i$. We will show that $\Delta_{e_{k}}^{i}(A_{i}^{r}\setminus\{e_{r}\})=1$ holds for any round $r>k$ after some new items are added to agent $i$’s bundle. Without loss of generality, we assume that $e_{s}$ is an arbitrarily newly added item in agent $i$’s bundle, where $s>k$, it holds that $\Delta_{e_{s}}^{i}(A_{i}^{s-1})=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1})=1$. Then, we have
$\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{r}\})$ $=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
$=v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
$\geq v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\setminus\{e_{k}\})-v_{i}(e_{s})$
$=v_{i}(A_{i}^{s-1})-v_{i}(A^{s-1}\setminus\{e_{k}\})$
$=\Delta_{e_{k}}^{i}(A_{i}^{s-1}\setminus\{e_{k}\}),$
where the inequality holds for submodularity. Following this induction by repeatedly utilizing submodularity, we have that $\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{k}\})\geq\cdots\geq\Delta_{e_{k}}^{i}(A_{i}^{k}\setminus\{e_{k}\})=\Delta_{e_{k}}^{i}(A_{i}^{k-1})=1$. Since $v_{i}$ is a submodular binary valuation function, it holds that $\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{k}\})\leq 1$. Thus, we can conclude that $\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{k}\})=1$. ∎
Claim 2.
For submodular valuation functions, there exists an instance that no deterministic online algorithm can compute an NW allocation, even for two agents with identical valuation functions.
Proof.
We construct an instance with two agents, $v_{1}=v_{2}=v$. When first item $e_{1}$ arrives, $v(e_{1})=3$. W.l.o.g., we allocate $e_{1}$ to agent 1. When the second item $e_{2}$ arrives, $v(e_{2})=3,v(e_{1},e_{2})=5$. Case (1): we allocate $e_{2}$ to agent 2, for the third item $e_{3}$,$v(e_{3})=5,v(e_{1},e_{3})=v(e_{2},e_{3})=v(e_{1},e_{2},e_{3})=5$. We have that $v(e_{1},e_{3})-v(e_{3})=v(e_{2},e_{3})-v(e_{3})=0$. So, no matter who is assigned $e_{3}$, it cannot satisfy NW. Case (2): we allocate $e_{2}$ to agent 1, for the third item $e_{3}$ with $v(e_{3})=5,v(e_{1},e_{3})=v(e_{2},e_{3})=v(e_{1},e_{2},e_{3})=5$. To ensure NW, we must allocate $e_{3}$ to agent 2. For the fourth item $e_{4}$, $v(e_{4})=v(e_{1},e_{4})=v(e_{2},e_{4})=v(e_{3},e_{4})=v(e_{1},e_{2},e_{4})=v(e_{1},e_{3},e_{4})=v(e_{2},e_{3},e_{4})=v(e_{1},e_{2},e_{3},e_{4})=6$. We have that $v(e_{1},e_{2},e_{4})-v(e_{2},e_{4})=v(e_{3},e_{4})-v(e_{4})=0$. So, no matter who is assigned $e_{4}$, it cannot satisfy NW. ∎
Input: An instance $(T,N,(v_{i})_{i\in N})$ with submodular binary valuation functions
Output: A non-wasteful allocation $A$
1 Initialize: Let $A=(\emptyset,\ldots,\emptyset)$;
2 when item $e\in T$ arrives do
3
4      for $i=1$ to $n$ do
5
6            if $\Delta_{e}^{i}(A_{i})>0$ then
7
8                  Let $A_{i}=A_{i}\cup\{e\}$;
9                  break;
10             end if
11
12       end for
13
14
15
16return $A$;
Algorithm 1 Greedy Algorithm
3.1 The Impossibilities of Approximate EF1 and MMS
In this part, we present a strong negative result regarding the approximations of EF1 and MMS for general additive valuations. The valuation function $v_{i}$ is tri-valued if $v_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
Theorem 1.
For general additive functions, no deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any constant $\alpha>0$, even for two agents with additive tri-valued valuation functions.
Proof.
Consider an online instance with two agents. For any agent $i\in N$ and item $e\in T$, $v_{i}(e)\in\{\epsilon,1,\frac{1}{\epsilon}\}$, where $10<\epsilon<$ 1. The items arrive online, and the valuation of each item is shown in Table 3. When $t=1$, the arriving item has the value of 1 for both agents. Without loss of generality, agent 1 picks it. When $t=2$, the item that has the value of $\frac{1}{\epsilon}$ for agent 1 and $\epsilon$ for agent 2 arrives. In order to avoid 00-EF1 or 00-MMS, this item must be allocated to agent 2. In round $t=3$, the arriving item has the value of $\frac{1}{\epsilon}$ for both agents. In this case, if this item is allocated to agent 1, for agent 2, the approximation ratio of EF1 is $\epsilon$, and the approximation ratio of MMS is $\frac{\epsilon}{1+\epsilon}$; if this item is allocated to agent 2, for agent 1, the approximation ratio of EF1 is $\frac{1}{\frac{1}{\epsilon}}=\epsilon$ and the approximation ratio of MMS is $\frac{1}{1+\frac{1}{\epsilon}}=\frac{\epsilon}{1+\epsilon}$. Therefore, for any constant $\alpha>0$, let $\epsilon<\alpha$, and we can conclude that no algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS for any given $\alpha>0$. This result can be generalized to more than two agents, and we defer the proof to Proposition 5.
$t=1$ $t=2$ $t=3$
Agent $1$ $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Agent 2 $1$ $\epsilon$ $\frac{1}{\epsilon}$
Table 3: The impossibility results of EF1 and MMS for additive tri-valued valuations
By Theorem 1, we cannot make further progress in the approximation of EF1 and MMS for tri-valued valuation functions and beyond in the online allocation problem. This means that we can only focus our efforts on bi-values. In the following parts, we focus on two classes of valuation functions, including submodular binary and additive bi-valued valuation functions, and study whether there are positive results about the approximation of EF1 or MMS.
3.2 Submodular Binary Valuations
This part focuses on submodular binary valuation functions, where for each agent $i\in N$, her valuation function $v_{i}$ is submodular and the marginal value of any item is 0 or 1. For a subclass of submodular binary valuation functions based on maximum (unweighted) bipartite matching, which is also called assignment [19] or OXS [20] valuations, Hosseini et al.[11] show that online allocations satisfying $\frac{1}{2}$-EF1 and $\frac{1}{2}$-MMS exist. In their model, there are some groups, each of which includes some members, and each member can only receive one item at most. When an item arrives, they greedily match the item to some member who likes the item and has never received any items. Leaving the online matching model, we consider the general online allocation problem, greedily assigning each arriving item to the agent that brings the maximum marginal utility. Deriving from the greedy technique of the submodular optimization problem and the sequential allocating technique to guarantee fairness in [11], we propose an algorithm named Marginal-Greedy Algorithm (Algorithm 2) for general binary valuation functions. We prove that input submodular binary valuations into Algorithm 2, we can get an NW allocation that is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW. Additionally, we conjecture that the Algorithm 2 will also perform well when we input subadditive binary valuation functions into Algorithm 2, and leave this conjecture as an open question.
Input: An instance $(T,N,(v_{i})_{i\in N})$ with general binay valuation functions
Output: A non-wasteful approximate EF1 and approximate MMS allocation $A$
1 Initialize: let $\pi=(\pi_{1},\cdots,\pi_{n})$ be an arbitrary order of $n$ agents;
2Let $A=(\emptyset,\ldots,\emptyset)$;
3 when item $e\in T$ arrives do
4
5      for $i=1$ to $n$ do
6
7            if $\Delta_{A_{\pi_{i}}}^{\pi_{i}}(e)>0$ then
8
9                  Let $A_{\pi_{i}}=A_{\pi_{i}}\cup\{e\}$;
10                  $\pi=(\pi_{1},\cdots,\pi_{i-1},\pi_{i+1},\cdots,\pi_{n},\pi_{i})$;
11                  break;
12             end if
13
14       end for
15
16
17
18return $A$;
Algorithm 2 Marginal-Greedy Algorithm
In the analysis of results about submodular binary valuation functions, we skillfully use submodularity and the properties of the matroid. Then, we introduce the relationship of submodularity and matroid: a function is submodular binary if and only if it is a rank function of a matroid.
Definition 5.
A set system $\mathcal{M}=(E,\mathcal{F})$, where $E$ is a ground set and $\mathcal{F}$ is a family of subset of $E$, is a matroid if it satisfies the following properties: (1) $\emptyset\in\mathcal{F}$; (2) if $D\in\mathcal{F}$ and $C\subseteq D$, then $C\in\mathcal{F}$; (3) if $C,D\in\mathcal{F}$ and $|D|>|C|$, then there exists $x\in D\setminus C$ such that $C+x\in\mathcal{F}$. The rank function of matroid $\mathcal{M}$, denoted by $r(\cdot)$, is defined by: $r:2^{T}\rightarrow R$, $r(X)=\max\{|Y|:Y\subseteq T,Y\in\mathcal{F}\}$.
Theorem 2.
For the deterministic allocation of indivisible goods with submodular binary valuation functions, the Marginal-Greedy Algorithm (Algorithm 2) computes an NW allocation that satisfies $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS and $\frac{1}{2}$-max-USW.
Proof.
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, where $v_{i}$ is submodular and binary for all $i\in N$. Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ be the set of items that have already arrived, denoted by ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ and $A^{k}_{0}$ the allocation and the set of all unallocated items at the end of round $k$ respectively. In the algorithm procedure, the arriving item is allocated to an agent who thinks that the marginal utility of this item is $1$ if such an agent exists. By Claim 1, $A^{k}_{i}$ is non-wasteful. For each submodular binary valuation function $v_{i}$, there exists a matroid such that $v_{i}$ is the rank function of this matroid. let $\mathcal{M}_{i}=(T,\mathcal{F}_{i})$ be the corresponding matroid of $v_{i}$. We have $v_{i}(A^{k}_{i})=|A^{k}_{i}|$, and $A^{k}_{i}\in\mathcal{F}_{i}$.
Firstly, we show that ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ is $\frac{1}{2}$-EF1. Consider any two agents $i,j\in N$, $i\neq j$. If $v_{i}(A^{k}_{j})>v_{i}(A^{k}_{i})$, by the property (3) of the matroid, there must exist an item $e\in A^{k}_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$. Let $B_{j}=\{e:e\in A^{k}_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$ and $B_{j}^{*}\in\arg\max\{|S|:S\subseteq B_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$. We will show $|B_{j}|\leq|A^{k}_{i}|+1$. Assume that $|B_{j}|\geq|A^{k}_{i}|+2$, $j$ has priority higher than $i$ at least $|B_{j}|\geq|A^{k}_{i}|+2$ times, which contradicts the order in the algorithm. Combining the property (3) of the matroid, we have $v_{i}(A^{k}_{j})=|A^{k}_{i}|+|B_{j}^{*}|\leq|A^{k}_{i}|+|B_{j}|\leq 2|A^{k}_{i}|+1$. Thus, ${A^{k}}$ is $\frac{1}{2}$-EF1.
Secondly, we show that ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ is $\frac{1}{2}$-MMS. Consider a MMS partition $X=(X_{1},\cdots,X_{n})$ for agent $i$ of item set $T$, we assume that the share of agent $i$ in ${A^{k}}$ is less than half of the maximin share of $i$, i.e. $v_{i}(A^{k}_{i})\leq\frac{1}{2}v_{i}(X_{j})$ for any $j\in N$. By the property (3) of the matroid, there exists an item $e\in X_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$. Let $C_{j}=\{e:e\in X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$, $C_{j}^{*}\in\arg\max\{|S|:S\subseteq C_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$. We have that $v_{i}(X_{j})=v_{i}(A^{k}_{i}\cup C_{j}^{*})=v_{i}(A^{k}_{i})+v_{i}(C_{j}^{*})=|A^{k}_{i}|+|C_{j}^{*}|$. Since $v_{i}(A^{k}_{i})=|A^{k}_{i}|\leq\frac{1}{2}v_{i}(X_{j})$, $|C_{j}|\geq|C_{j}^{*}|>|A^{k}_{i}|+1$. Then, $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$. We have $|\cup_{j\in N}C_{j}|=|\{e:e\in\cup_{j\in N}X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\{e:e\in T,A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\cup_{j\in N}B_{j}|\leq(n-1)(|A^{k}_{i}|+1)$, it contradicts the inequality $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$. Thus, ${A^{k}}$ is $\frac{1}{2}$-MMS.
Finally, we show that ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ is $\frac{1}{2}$-USW. We make a slight adjustment to the Algorithm 2: when a new item arrives, if the marginal utility is 0 for all agents, then assign the item to agent $1$ without updating the agents’ order. Then, we will show that ${A^{k^{\prime}}}=(A^{k}_{1}\cup A^{k}_{0},\cdots,A^{k}_{n})$ is the output of the adjusted algorithm. We only need to prove that in each round $s\in[k]$, if the marginal utility of the arrived item $e_{s}$ to $A^{s-1}_{1}$ is $1$, $v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1$, by the monotonicity of $v_{i}$, $v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0}\cup\{e_{s}\})\geq v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1=v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0})+1$. Then, the marginal utility to $A^{s-1}_{1}\cup A^{s-1}_{0}$ is also $1$. So, ${A^{k^{\prime}}}=(A^{k}_{1}\cup A^{t}_{0},\cdots,A^{k}_{n})$ is the output of the adjusted algorithm. Now, we will show that ${A^{k^{\prime}}}$ is $\frac{1}{2}$-USW. The remaining proof part is based on the idea from [20]. Consider the first item $e_{1}$, we assume that $e_{1}$ is allocated to agent $i^{*}$ during the execution of the algorithm and $v_{i^{*}}(e_{1})=1$, $A_{i^{*}}^{1}=\{e_{1}\}$ and $A_{i}^{1}=\emptyset$ for $i\neq i^{*}$. Let ${Y}=(Y_{1},\cdots,Y_{n})$ be an allocation achieving the optimal USW of instance $I=(T=\{e_{1},\cdots,e_{k}\},N,(v_{i})_{i\in N})$, and we assume that $e_{1}$ is allocated to agent $i^{o}$ in $Y$, $e_{1}\in Y_{i^{o}}$. Next, we construct a new instance ${I^{(1)}}=(T^{\prime},N,(v_{i}^{(1)})_{i\in N})$, where $v_{i^{*}}^{(1)}={\Delta^{i^{*}}_{e_{1}}}$ and $v_{i}^{(1)}=v_{i}$ for all $i\neq i^{*}$. In $I^{(1)}$, $T^{\prime}=T\setminus\{e_{1}\}$, that is, the first item arriving is $e_{2}$. Input the instance $I^{(1)}$ into the adjusted algorithm. Let initialized order of $N$ be $\pi=(\pi_{1},\cdots,\pi_{n},\pi_{i^{*}})$, the initialized allocation be ${D^{0}}=(D^{0}_{1},\cdots,D^{0}_{n})$ , where $D_{i^{*}}=\{e_{1}\}$ and $D_{i}=\emptyset$ for all $i\neq i^{*}$. Assume ${D^{k}}=(D^{k}_{1},\cdots,D^{k}_{n})$ is the allocation at the end of round $k$. We have ${D^{k}}={A^{k^{\prime}}}$, then $USW({D^{k}})=USW({A^{k^{\prime}}})=USW({A^{k}})$. Now, we costruct an allocation ${S}=(S_{1},\cdots,S_{n})$ for instance $I^{\prime}$ as follows: $S_{i^{o}}=Y_{i^{o}}\setminus\{e_{1}\}$, $S_{i}=Y_{i}$ for all $i\neq{i^{o}}$. It holds $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})=v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$. By the monotonicity of $v_{i^{*}}$, $v_{i^{*}}^{(1)}(S_{i^{*}})=v_{i^{*}}(Y_{i^{*}}\cup\{e_{1}\})-v_{i^{*}}(\{e_{1}\})\geq v_{i^{*}}(Y_{i^{*}})-v_{i^{*}}(\{e_{1}\})$. $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$, where the first inequality holds because of the submodularity and the second inequality holds because of $v_{i^{o}}(\{e_{1}\})\leq v_{i^{*}}(\{e_{1}\})$, Let ${Y^{(1)}}=(Y_{1}^{(1)},\cdots,Y_{n}^{(1)})$ be an allocation achieving the optimal USW of instance $I^{(1)}$. It holds that $USW({Y^{(1)}})\geq USW(S)=\cup_{i\in N}v_{i}^{(1)}(S_{i})=\cup_{i\in N,i\neq i^{*},i\neq i^{o}}v_{i}^{(1)}(S_{i})+v_{i^{o}}^{(1)}(S_{i^{o}})+v_{i^{*}}^{(1)}(S_{i^{*}})\geq USW(Y)-2v_{i^{*}}(\{e_{1}\})$. By lemma 1 in [20], $v_{i^{*}}^{(1)}$ is also submodular valuation. So, by constructing the new instances $I^{(2)},\cdots,I^{(k-1)}$ and inducting each round based on the items that arrive, we have $USW(A^{k})=USW(A^{k^{\prime}})=USW(D^{k})=USW(Y^{k})\geq\frac{1}{2}USW(Y).$
Theorem 3.
For submodular binary valuation functions, no deterministic online algorithm can achieve $(\frac{1}{2}+\epsilon)$-EF1 or $(\frac{1}{2}+\epsilon)$-MMS, and non-wastefulness simultaneously for any $\epsilon>0$.
Proof.
Figure 1: An instance with $\frac{1}{2}$-EF1 and $\frac{1}{2}$-MMS allocations
Consider an online allocation instance, where there are two agents $\{1,2\}$ with identical valuations $v$, and the online items belong to several categories $C_{1},\cdots,C_{l}$. In Figure 1, categories are distinguished by colours, and edges of different colours represent items of different categories. Denote a partition matroid $\mathcal{M}=(T,\mathcal{F})$, =$\mathcal{F}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~{}k\in[l]\}$. The valuation $v$ is the rank function of the matroid $\mathcal{M}$; for any item set $S$, the value of $S$ is the number of items of the maximal independent set of $S$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are $1$. Without loss of generality, we allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, to avoid breaking fairness, we must allocate it to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are $1$. We will discuss the following two cases.
Case 1: Allocate $e_{3}$ to agent $2$, shown in Figure 1(1). We must allocate the fourth item, $e_{4}$, to agent $2$ due to the requirement of non-wastefulness. Then, $A^{4}_{1}=\{e_{1}\}$,$A^{4}_{2}=\{e_{2},e_{3},e_{4}\}$. We have that, $v(A^{4}_{1})=\frac{1}{2}v(A^{4}_{2}\setminus\{e\})$ for any $e\in A^{4}_{2}$, and $v(A^{4}_{1})=\frac{1}{2}\mathsf{MMS}_{1}$.
Case 2: Allocate $e_{3}$ to agent $1$, shown in Figure 1(2); we must allocate the fourth item, $e_{4}$, to agent $1$ due to the requirement of non-wastefulness. Then, $A^{4}_{1}=\{e_{1},e_{3},e_{4}\}$,$A^{4}_{2}=\{e_{2}\}$. We have that, $v(A^{4}_{2})=\frac{1}{2}v(A^{4}_{1}\setminus\{e\})$ for any $e\in A^{4}_{1}$, and $v(A^{4}_{2})=\frac{1}{2}\mathsf{MMS}_{2}$. ∎
Theorem 4.
For submodular binary valuation functions, no deterministic online algorithm can achieve $(\frac{1}{2}+\epsilon)$-max-USW and non-wastefulness for any $\epsilon>0$.
Proof.
Consider an online allocation instance with two agents $\{1,2\}$. For the first arrived item $e_{1}$, the marginal utilities for agents $1$ and $2$ are both $1$, $v_{1}(\{e_{1}\})=v_{2}(\{e_{1}\})=1$. Without loss of generality, we allocate $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, $v_{1}(\{e_{2}\})=v_{1}(\{e_{1},e_{2}\})=1$,$v_{2}(\{e_{2}\})=0$, the marginal utilities for agents $1$ and $2$ are both 00. No matter which agent gets $e_{2}$, $USW(A^{2})=1$. However, the max-USW is $2$ by allocating $e_{1}$ to agent $2$ and $e_{1}$ to agent $1$. ∎
There is a side result of our algorithm, where the Marginal-Greedy Algorithm (Algorithm 2) can also work for additive binary valuation functions to compute an NW allocation that is EF1, MMS, and max-USW, where for each agent $i\in N$, $v_{i}$ is additive and $v_{i}(e)\in\{0,1\}$ for any $e\in T$.
3.3 Additive Bi-valued Valuations
In this part, we consider additive bi-valued valuation functions. For any agent $i\in N$, $v_{i}$ is bi-valued if $v_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$. Someone may wonder whether a non-wasteful EF1 or MMS allocation can be found with additive bi-valued valuation functions because the additive binary valuation function can be seen as a special additive bi-valued valuation function. However, we find that even for two agents with additive bi-valued valuations, a non-wasteful EF1 or MMS allocation cannot be guaranteed.
3.3.1 Two Agents
We study the setting of additive bi-valued valuations for two agents and establish the best possible result for the approximation of EF1 and MMS simultaneously. In order to achieve this goal, we propose the Adapted Envy-Graph Procedures (Algorithm 3). Note that there are two technical difficulties for us when designing this approximation algorithm. One is that we do not know the value of future items. The other is that we cannot do any reallocation.
The main idea of our algorithm is that we take the Adapted Envy-Graph Procedures as the base algorithm to allocate items. If we follow the base algorithm, there will be an envy cycle between two agents, and we temporarily terminate it. Then, we consider implementing the Preliminary-Breaking-Cycle Algorithm (Algorithm 4) to eliminate the potential envy cycle manually. Then, there are two possible results of Algorithm 4. One is that we have eliminated the potential envy cycle successfully. The other is that there is still an envy cycle, and we execute the Deep-Breaking-Cycle Algorithm (Algorithm 5) to eliminate it further. In whichever case, when there is no envy cycle and the allocation is EF1, we go back to the base algorithm. The illustration of Algorithms 3, 4 and 5 is shown in Figure 2. Before presenting the theorem, we first define an envy graph.
Figure 2: The illustration of Algorithms 3, 4 and 5.
Definition 6 (Envy Graph for Goods).
Given an allocation $A^{k}$, in an envy graph G, each vertex corresponds to an agent, and each directed edge $(i,j)$ represents that agent $i$ envies agent $j$, where we have $v_{i}(A_{i}^{k})<v_{i}(A_{j}^{k})$.
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive bi-valued valuation functions
Output: An non-wasteful approximate EF1 and approximate MMS allocation $A$
1
2Let $A=(\emptyset,\ldots,\emptyset)$ and build an envy graph $G$ for two agents;
3 when item $e\in T$ arrives do
4       if  there is no envy between the two agents  then
5             if $\forall k\in N$, $v_{k}(e)=a_{k}$ or $\forall k\in N$, $v_{k}(e)=b_{k}$  then
6                   Arbitrarily choose one agent $i\in N$;
7            else
8                   Choosing the agent $i\in N$ who has the valuation of $b_{i}$;
9             end if
10
11      else
12             Let $i\in N$ be an unenvied agent and $j=2-i$;
13            if $v_{i}(A_{i}\cup\{e\})<v_{i}(A_{j})$ then
14                   Terminate this algorithm and implement the Preliminary-Breaking-Cycle algorithm (Algorithm 4);
15
16       end if
17      $A_{i}=A_{i}\cup\{e\}$;
18
return $A$;
Algorithm 3 Adapted Envy-Graph Procedures
Theorem 5.
For the deterministic allocation of indivisible goods for two agents with additive bi-valued valuations, the Adapted Envy-Graph Procedures (Algorithm 3) computes an NW allocation that satisfies $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS.
Before we show Theorem 5, we first give the following key observations and lemmas.
Observation 1.
When Algorithm 4 is executed, for the first arriving item $e$, we have $v_{i}(e)=a_{i}$.
Proof.
The reason why Algorithm 4 is executed is that for the arriving item $e$, if we add it to agent $i$’s bundle, there will be an envy cycle between two agents. i.e., $v_{i}(A_{i}\cup\{e\})<v_{i}(A_{j})$ and $v_{j}(A_{j})<v_{i}(A_{i}\cup\{e\})$. Suppose for the contradiction that $v_{i}(e)=b_{i}$. Since the input allocation $A$ is EF1, for agent $i$, we have $v_{i}(A_{j})-v_{i}(A_{i})\leq b_{i}$. Thus, allocating item $e$ to agent $i$ eliminates the envy and cannot lead to an envy cycle, which is a contradiction. Therefore, for the arriving item $e$, it holds that $v_{i}(e)=a_{i}$. ∎
Observation 2.
When Algorithm 4 is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
Proof.
Suppose that $|A_{i}|=0$. In that case, we have $|A_{j}|=1$. Otherwise, there would be an envy cycle. After allocating $e$, there will be an envy cycle, i.e., $v_{i}(e)=a_{i}<v_{i}(A_{j})=b_{i}$ and $v_{j}(A_{j})<v_{j}(e)$. If the second inequality holds, we have $v_{j}(A_{j})=a_{j}$ but $v_{i}(A_{j})=b_{i}$, which contradicts the line 3 in Algorithm 3. Suppose that $|A_{j}|=0$. In that case, we have $|A_{i}|=0$, which contradicts the fact that after allocating $e$, an envy cycle appears. ∎
Lemma 1.
Algorithm 4 computes an NW allocation that is EF1 with or without an envy cycle, or $\frac{1}{2}$-EF1.
Proof.
Before Algorithm 4 is implemented, the output allocation $A=(A_{1},A_{2})$ by Algorithm 3 is EF1. According to the algorithm description, it is executed in at most two rounds.
$\lambda=1$. For the first arriving item $e_{1}$, by Observation 1, we have $v_{i}(e_{1})=a_{i}$. If $\lvert A_{j}\rvert=1$, we have $v_{i}(A_{j})=b_{i}$ since agent $i$’s bundle is EF1. Then, agent $j$ picks this item. For agent $i$, her bundle is EF1 since we have $\lvert A_{i}\lvert\geq 1$ and $v_{i}(e_{1})=a_{i}$. For agent $j$, her new bundle is EF since she does not envy agent $i$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. If $\lvert A_{j}\rvert>1$, let us consider two cases about the valuation of item $e_{1}$.
  • If $v_{j}(e_{1})=b_{j}$, item $e_{1}$ is allocated to agent $j$ and we set $\mu=1$. In this case, for agent $j$, her new bundle is EF since $v_{j}(A_{j}\cup\{e_{1}\})>v_{j}(A_{j})\geq v_{j}(A_{i})$ and for agent $i$, her bundle is $\frac{1}{2}$-EF1, i.e, $v_{i}(A_{i})\geq\frac{1}{2}v_{i}(A_{j}\cup\{e_{1}\}\setminus\{g\})$ holds for $g\in A_{j}\cup\{e_{1}\}$, since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(A_{i})\geq a_{i}=v_{i}(e_{1})$. Thus, this allocation is $\frac{1}{2}$-EF1.
  • If $v_{j}(e_{1})=a_{j}$, item $e_{1}$ is allocated to agent $i$ and we set $\mu=2$. In this case, for agent $i$, her new bundle is still EF1 since she is an unenvied agent in Algorithm 3 and the input allocation $A$ is EF1, and for agent $j$, her bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ but $v_{j}(A_{j})<v_{j}(A_{j}\cup\{e_{1}\})$. Thus, this allocation is EF1 with an envy cycle.
$\lambda=2$. The second item $e_{2}$ arrives. Let us consider two cases about $\mu$.
  • If $\mu=1$, agent $i$ picks it. In this case, for agent $j$, her bundle is still EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{1})=b_{j}\geq v_{j}(e_{2})$. For agent $i$, her new bundle is EF1 since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(e_{2})\geq v_{i}(e_{1})=a_{i}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • If $\mu=2$, let us consider two cases. When $v_{i}(e_{2})=a_{i}$, agent $j$ picks it. For agent $i$, her bundle is still EF1 since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(e_{1})=v_{i}(e_{2})=a_{i}$. For agent $j$, her new bundle is EF because $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{2})=v_{j}(e_{1})=a_{j}$. Hence, the allocation is EF1 without an envy cycle, and the algorithm is terminated. When $v_{i}(e_{2})=b_{i}$, it is allocated to agent $i$. For agent $i$, her new bundle is EF since it holds that $v_{i}(A_{i}\cup\{e_{1},e_{2}\})=v_{i}(A_{i})+a_{i}+b_{i}\geq v_{i}(A_{j}\setminus\{g\})+v_{i}(g)+a_{i}$ for some $g\in A_{j}$ where the inequality follows from the fact that $v_{i}(g)\in\{a_{i},b_{i}\}$. For agent $j$, her bundle is $\frac{1}{2}$-EF1, i.e., $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\cup\{e_{1},e_{2}\}\setminus\{g\})$ holds for some $e\in A_{i}\cup\{e_{1},e_{2}\}$, since $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(A_{j})\geq v_{j}(e_{1})=\min\{v_{j}(e_{1}),v_{j}(e_{2})\}$. Thus, this allocation is $\frac{1}{2}$-EF1 and the algorithm is terminated.
Combining the above cases, it is not hard to see that our claim is correct. ∎
Observation 3.
When Algorithm 5 is executed, there are two properties about input allocation $A$:
  • $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$;
  • Let $e_{1}$ and $e_{2}$ denote two new allocated items in Algorithm 4, we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$, where $e\in\arg\max_{g\in\{e_{1},e_{2}\}}v_{j}(g)$.
Proof.
When Algorithm 5 is ready to be executed, Algorithm 4 is terminated for $\lambda=\mu=2$. Then, by the proof of Algorithm 4, it is not hard to see that the statements are correct. ∎
Lemma 2.
Algorithm 5 computes an NW allocation that is EF1 with or without an envy cycle, or $\frac{1}{2}$-EF1.
Proof.
By Observation 3, the input allocation $A=(A_{1},A_{2})$ is $\frac{1}{2}$-EF1, where $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$ for some $e\in A_{i}$. Unlike Algorithm 4, this algorithm may be implemented in many rounds. Let $g_{1}$ and $g_{2}$ denote the two online items that are allocated to agent $i$ in Algorithm 4 and we have $\min\{v_{j}(g_{1}),v_{j}(g_{2})\}=a_{j}$.
$\lambda=1$. For the arriving item $e_{1}$ when $\lambda=1$, let us consider four cases of agents’ valuations.
  • When $v_{i}(e_{1})=a_{i}$ and $v_{j}(e_{1})=b_{j}$, agent $j$ picks it and we set $\mu=1$. For agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 without an envy cycle.
  • When $v_{i}(e_{1})=b_{i}$ and $v_{j}(e_{1})=b_{j}$, it is allocated to agent $j$ and we set $\mu=2$. For agent $i$, her bundle is EF1 since after removing $e_{1}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$. For agent $j$, her new bundle is also EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 with an envy cycle.
  • When $v_{i}(e_{1})=a_{i}$ and $v_{j}(e_{1})=a_{j}$, it is allocated to agent $j$ and we set $\mu=3$. For agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=a_{i}=\min\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 without an ency cycle.
[…]
$\lambda=2$. For the arriving item $e_{2}$ when $\lambda=2$, let us consider four cases about $\mu$.
  • When $\mu=1$, item $e_{2}$ is directly allocated to agent $j$. For agent $i$, her bundle is EF1 since after removing $e_{2}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j}\cup\{e_{1}\})=v_{i}(A_{j})+a_{i}$. For agent $j$, her new bundle is EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{1})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{2})\geq v_{j}(g_{1})=a_{j}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • When $\mu=2$, we need to consider two subcases. If $v_{i}(e_{2})=b_{i}$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{2})=b_{i}=v_{i}(e_{1})$. For agent $j$, her bundle is $\frac{1}{2}$-EF1 since we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{j}(e_{1})=b_{j}\geq v_{j}(e_{2})$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda=1$ and continue executing this algorithm. If $v_{i}(e_{2})=a_{i}$, it is allocated to agent $j$. For agent $i$, her bundle is EF1 since after removing $e_{1}$ from $A_{j}$,we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{2}\})$. For agent $j$, her new bundle is trivially EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{1})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{2})\geq a_{i}=v_{j}(g_{1})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • When $\mu=3$, let us also consider two subcases. If $v_{j}(e_{2})=b_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF1 since after removing $e_{2}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j}\cup\{e_{1}\})=v_{i}(A_{j})+a_{i}$. For agent $j$, her new bundle is EF since $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{2})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{1})=a_{j}=v_{j}(e_{2})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. If $v_{j}(e_{2})=a_{j}$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{2})=a_{i}=v_{i}(e_{1})$. For agent $j$, her bundle is $\frac{1}{2}$-EF1 since we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{i}$ and $v_{j}(e_{1})=a_{j}=v_{j}(e_{2})$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda=1$ and continue executing this algorithm.
  • When $\mu=4$, it is directly allocated to agent $j$. For agent $i$, her bundle is EF since $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{1})=b_{i}\geq v_{i}(e_{2})$. For agent $j$, her new bundle is $\frac{1}{2}$-EF1 since $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{i}$ and $v_{j}(e_{2})\geq v_{j}(e_{1})=a_{j}$. Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3. Additionally, we set $\lambda=1$ and continue executing this algorithm.
Combining the above cases, our claim holds. ∎
Proof of Theorem 5.
In the main body of Algorithm 3, we follow a similar way of offline envy graph procedures to allocate items. The precondition of Algorithm 3 is that there is no envy cycle between two agents. Initially, the empty allocation is EF1 without an envy cycle. When there is no envy between two agents, we allocate the arriving item to one agent according to different conditions. When there exists an envy between two agents, we have two cases. One is that if we allocate it to the unenvied agent, the envy cycle does not appear. We will follow this procedure. The other is that if we allocate it to the unenvied agent, the envy cycle will appear, and we terminate Algorithm 3. In the online setting, we cannot swap bundles to eliminate the envy cycle. Thus, to enable the Algorithm 3 to be executed normally, Algorithms 4 and 5 help us to manually break the possible envy cycle, where, by Lemmas 1 and 2, we guarantee that $\frac{1}{2}$-EF1 always holds. When the allocation is EF1 and there is no envy cycle, we continue executing Algorithm 3. By the induction, we can conclude that the output allocation of Algorithm 3 is always EF1 without an envy cycle. Combining Lemmas 1, 2, and the description of Algorithm 3, it can be seen that we can guarantee an NW allocation that is $\frac{1}{2}$-EF1. As for MMS, by the implication between EF1 and MMS [21], i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{\alpha}{(n-1)\alpha+1}$-MMS, where $n$ is the number of agents, it is clear that the output allocation is $\frac{\frac{1}{2}}{\frac{1}{2}+1}=\frac{1}{3}$-MMS. ∎
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive bi-valued valuation functions
Output: An non-wasteful approximate EF1 and approximate MMS allocation $A$
1
2Initialize: $\lambda=1$ and $\mu=0$;
3 when item $e\in T$ arrives do
4
5      if $\lambda=1$ then
6
7            if $\lvert A_{j}\rvert=1$ then
8
9                  $A_{j}=A_{j}\cup\{e\}$;
10                  Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
11            else
12
13                  if $v_{j}(e)=b_{j}$ then
14
15                        $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
16                  else
17                         $A_{i}=A_{i}\cup\{e\}$ and $\mu=2$;
18                   end if
19                  $\lambda=2$;
20             end if
21
22      else
23
24            if $\mu=1$  then
25
26                  $A_{i}=A_{i}\cup\{e\}$;
27            else
28
29                  if $v_{i}(e)=a_{i}$ then
30
31                        $A_{j}=A_{j}\cup\{e\}$;
32                  else
33
34                        $A_{i}=A_{i}\cup\{e\}$;
35                        Terminate this algorithm and execute the deep-breaking-cycle algorithm (Algorithm 5);
36                   end if
37
38             end if
39
40            Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
41       end if
42
43
return $A$;
Algorithm 4 Preliminary-Breaking-Cycle Algorithm
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive bi-valued valuation functions
Output: An non-wasteful approximate EF1 and approximate MMS allocation $A$
1 Initialize: $\lambda=1$ and $\mu=0$;
2 when item $e\in T$ arrives do
3
4      if $\lambda=1$ then
5
6            if $v_{i}(e)=a_{i}$ and $v_{j}(e)=b_{j}$ then
7                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
8             else if $v_{i}(e)=b_{i}$ and $v_{j}(e)=b_{j}$ then
9
10                  $A_{j}=A_{j}\cup\{e\}$ and $\mu=2$;
11             else if  $v_{i}(e)=a_{i}$ and $v_{j}(e)=a_{j}$  then
12                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=3$;
13             else
14                   $A_{i}=A_{i}\cup\{e\}$ and $\mu=4$;
15             end if
16            $\lambda=2$;
17      else
18             if $\mu=1$ then
19                   $A_{j}=A_{j}\cup\{e\}$;
20             else if  $\mu=2$  then
21                   if $v_{i}(e)=b_{i}$ then
22
23                        $A_{i}=A_{i}\cup\{e\}$;
24                        $\lambda=1$;
25                  else
26                         $A_{j}=A_{j}\cup\{e\}$;
27                   end if
28
29             else if  $\mu=3$  then
30                   if $v_{j}(e)=b_{j}$ then
31
32                        $A_{j}=A_{j}\cup\{e\}$;
33                  else
34
35                        $A_{i}=A_{i}\cup\{e\}$;
36                        $\lambda=1$;
37                   end if
38
39             else
40
41                  $A_{j}=A_{j}\cup\{e\}$;
42                  $\lambda=1$;
43             end if
44
45            if $\lambda=2$ then
46                   Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
47
48       end if
49
50
51
52return $A$;
Algorithm 5 Deep-Breaking-Cycle Algorithm
Then, we present the impossible results to show that our approximation guarantees for EF1 and MMS are tight, and there is no compatibility between EF1 or MMS, and USW.
Theorem 6.
For two agents with additive bi-valued valuation functions, no deterministic online algorithm can compute an NW allocation that is $(\frac{1}{2}+\epsilon)$-EF1 or $(\frac{1}{3}+\epsilon)$-MMS for any $\epsilon>0$.
Proof.
$t=1$ $t=2$ $t=3$ $t=4$
agent $1$ $\epsilon$ $1$ $\epsilon$ 1
agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1
Case 1
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $\epsilon$ $1$ $\epsilon$ $\epsilon$ $1$
agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1 $1$
Case 2-1
$t=1$ $t=2$ $t=3$ $t=4$
agent $1$ $\epsilon$ $1$ $\epsilon$ $\epsilon$
agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1
Case 2-2
Table 4: The impossibility results of EF1 and MMS for additive bi-valued valuations for two agents.
Consider an online allocation instance with two agents $\{1,2\}$. For each agent $i\in\{1,2\}$, $v_{i}(e)\in\{\epsilon,1\}$, where $10<\epsilon<$ 1. The value of each arrived item is shown in Table 4. When $t=1$, the item with the value of $\epsilon$ for both agents arrives. Without loss of generality, we allocate $e_{1}$ to agent 1. When $t=2$, the arriving item has the value of 1 for agent 1 and $\epsilon$ for agent 2. To avoid 0-EF1 and 0-MMS, it should be allocated to agent 2. Then, when $t=3$, the arriving item has the value of $\epsilon$ for both agents. Let us consider two different choices of this item.
Case 1. The arriving item in round $t=3$ is allocated to agent 1. Then, in round $t=4$, the item that has the value of 1 for both agents arrives. Since we consider non-wastefulness, one agent must pick it. If it is allocated to agent 1, for agent 1, her bundle is EF1 and MMS to her, and for agent 2, her bundle is $\frac{\epsilon}{1+\epsilon}$-EF1 and $\frac{1}{3}$-MMS. If it is allocated to agent 2, for agent 1, her bundle is $2\epsilon$-EF1 and $\frac{2\epsilon}{1+\epsilon}$-MMS, and for agent 2, her bundle is EF1 and MMS. Therefore, we can derive that the upper bound of EF1 is $2\epsilon$ and MMS is $\frac{1}{3}$ respectively.
Case 2. The arriving item in round $t=3$ is allocated to agent 2. Then, in round $t=4$, the item has the value of $\epsilon$ for agent 1 and 1 for agent 2. Let us consider two choices for allocating this item.
  • If it is allocated to agent 1, this allocation is EF1 and $\frac{2}{3}$-MMS. At last, in round $t=5$, the arriving item has the value of 1 for both agents. To guarantee non-wastefulness, it must be allocated to some agent. It is not hard to check that no matter which agent picks it, the final allocation is $\frac{2\epsilon}{1+\epsilon}$-EF1 and $\frac{2\epsilon}{1+\epsilon}$-MMS.
  • If it is allocated to agent 2, there is no further arriving item. For agent 1, her bundle is $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS, and for agent 2, her bundle is EF and MMS. Thus, this allocation is $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS.
Therefore, we can derive that the upper bound of EF1 is $\frac{1}{2}$ and MMS is $\frac{1}{3}$ respectively. Combining the above two cases, it is clear that no algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\beta$-MMS for any $\alpha>\frac{1}{2}$ and $\beta>\frac{1}{3}$. ∎
Proposition 1.
For the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS, and $\beta$-max-USW for any $\alpha,\beta>0$, even for two agents with additive bi-valued valuation functions.
Proof.
Consider an online allocation instance with two agents $\{1,2\}$. Assume that, for any item $e$, $v_{1}(e)\in\{1,\frac{1}{\epsilon}\}$ and $v_{2}(e)\in\{\epsilon^{2},\epsilon\}$, where $10<\epsilon<$ 1. When $t=1$, the first item $e_{1}$ with the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is EF1, MMS, and $\epsilon$-max-USW. If not, agent $1$ picks it, and this allocation is EF1, MMS, and max-USW. Then, when $t=2$, the second arriving item $e_{2}$ has the value of 1 for agent 1 and $\epsilon^{2}$ for agent 2. If it is allocated to agent 1, this allocation is 0-EF1, 0-MMS, and max-USW. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\epsilon^{2}+1}{1+\frac{1}{\epsilon}}=\frac{\epsilon^{3}+\epsilon}{1+\epsilon}$-max-USW. For any $\beta>0$, let $\epsilon<\beta^{2}$, we can derive that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1/MMS and $\beta$-max-USW for any $\alpha,\beta>0$
3.3.2 The Combination with Additive Binary Valuations
In this part, we consider the combination of additive bi-valued and additive binary valuation functions and study the setting where only one agent (agent $n$) has the additive bi-valued valuation function, and the remaining agents have identical additive binary valuation functions. We show that an NW allocation that is EF1 and MMS always exists and can be computed by the Adapted-Picking Algorithm (Algorithm 6). The high level idea of our algorithm is that we allocate the appearing items that all agents value at non-zero values evenly with respect to the requirement that agent $n$ has the top priority of picking the item with the value of $b_{n}$ and the lowest priority of picking the item with the value of $a_{n}$. In that case, agent $n$ takes advantage of the option of items, and then we can guarantee that EF1 and MMS hold simultaneously.
Theorem 7.
For the deterministic allocation of indivisible goods for $(n-1)$ agents with identical additive binary valuation functions and one agent with the additive bi-valued valuation function, the Adapted-Picking Algorithm (Algorithm 6) computes an NW allocation that satisfies EF1 and MMS.
Input: An instance $(T,N,(v_{i})_{i\in N})$ with identical $(n-1)$ additive binary valuation functions $(v_{1},\ldots,v_{n})$ and one additive bi-valued valuation function $(v_{n})$
Output: An non-wasteful EF1 and MMS allocation $A$
1
2Let $A=(\emptyset,\ldots,\emptyset)$;
3 when item $e\in T$ arrives do
4
5      Let $i_{min}\in\arg\min_{i\in[n-1]}|A_{i}|$;
6      if $v_{i_{min}}(e)=0$ then
7
8            $A_{n}=A_{n}\cup\{e\}$;
9      else
10
11            if $v_{n}(e)=b_{n}$ then
12
13                  if |$|\{g~{}|~{}g\in A_{n}~{}and~{}v_{i_{min}}(g)>0\}|\leq\min_{i\in[n-1]}|A_{i}|$ then
14                         $A_{n}=A_{n}\cup\{e\}$;
15                  else
16
17                        $A_{i_{min}}=A_{i_{min}}\cup\{e\}$;
18                   end if
19
20            else
21                   if |$|\{g~{}|~{}g\in A_{n}~{}and~{}v_{i_{min}}(g)>0\}|\geq\min_{i\in[n-1]}|A_{i}|$ then
22
23                        $A_{i_{min}}=A_{i_{min}}\cup\{e\}$;
24                  else
25                         $A_{n}=A_{n}\cup\{e\}$;
26                   end if
27
28             end if
29
30       end if
31
32
return $A$;
Algorithm 6 Adapted-Picking Algorithm
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$ldots
agent $1$ $1$ $1$ 00 1 1 $\ldots$
agent 2 $a_{2}$ $a_{2}$ $a_{2}$ $b_{2}$ $a_{2}$ldots
Table 5: A simple illustration for two agents of Algorithm 6.
Proof.
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$. Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 6. Firstly, we show that the allocation in any round is EF1 by induction. For the base case, the empty allocation is trivially EF1. For the induction step, after allocating item $e_{k}$, the allocation $A^{k}$ is EF1. Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k}$ is still EF1. When item $e_{k+1}$ arrives, let us consider several cases about the allocation of this item.
  • If $v_{i_{min}}(e_{k+1})=0$, it is allocated to agent $n$. For any agent $i\in[n-1]$, she values it at zero, so her bundle is still EF1 to her. For agent $n$, the value of her bundle increases, so her bundle is also EF1 to her.
  • If $v_{n}(e_{k+1})=b_{n}$, let us consider two subcases. When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|\leq\min_{i\in[n-1]}|A_{i}^{k}|$, agent $n$ picks it. For agent $n$, her new bundle is EF. That is because she has the priority of picking the item with the value of $b_{n}$ and before allocating item $e_{k+1}$, if she envies other agents $j\in[n-1]$, in her perspective, the gap between the value of agent $j$’s bundle and her bundle is at most $a_{n}$, i.e.,$v_{n}(A_{j})-v_{n}(A_{n})\leq a_{n}$. Now, fix an arbitrary agent $i\in[n-1]$. Since agent $i$ has an identical binary valuation function and she only picks the item with the non-zero value, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{n}^{k})=v_{i}(A_{n}^{k+1}\setminus\{e_{k+1}\})$, implying her bundle is also EF1. When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|>\min_{i\in[n-1]}|A$ , $i_{min}$ picks it. Clearly, for agent $i_{min}$, her new bundle is still EF1. For any other agent $i\in N\setminus\{i_{min},n\}$, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{i_{min}}^{k})=v_{i}(A_{i_{min}}^{k+1}\setminus\{e_{k+1}\})$, implying that her bundle is still EF1. For agent $n$, based on the description of our algorithm, she must already pick one item that has the value of $b_{n}$, and in her perspective, the number of items in her bundle is at least one more than that of agent $i_{min}$’s bundle, which means that agent $n$ does not envy agent $i_{min}$ and in her perspective, the value of her bundle is at least $b_{n}$ greater than agent $i_{min}$’s bundle. Therefore, allocating $e_{k+1}$ to agent $i_{min}$ will not make agent $n$ envy agent $i_{min}$.
  • If $v_{n}(e_{k+1})=a_{n}$, let us consider two subcases. When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|\geq\min_{i\in[n-1]}|A_{i}^{k}|$, agent $i_{min}$ picks it. It is trivial that for agent $i_{min}$, her new bundle is EF1. For any other agent $i\in N\setminus\{i_{min},n\}$, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{i_{min}}^{k})=v_{i}(A_{i_{min}}^{k+1}\setminus\{e_{k+1}\})$, implying that her bundle is still EF1. For agent $n$, before allocating $e_{k+1}$, she does not envy agent $i_{min}$ based on the fact that the number of items in her bundle is at least the same as that of agent $i_{min}$’s bundle and our algorithm description. When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|<\min_{i\in[n-1]}|A_{i}^{k}|$, agent $n$ picks it. For agent $n$, her new bundle is actually EF since she has the priority of picking the item with the value of $b_{n}$ and before allocating $e_{k+1}$, she envies other agent $j\in[n-1]$; in her perspective, the gap between the value of agent $j$’s bundle and her bundle is at most $a_{n}$, i.e.,$v_{n}(A_{j})-v_{n}(A_{n})\leq a_{n}$. Fix an arbitrary agent $i\in[n-1]$. Since agent $i$ has the same binary valuation function and she only picks the item with the non-zero value, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{n}^{k})=v_{i}(A_{n}^{k+1}\setminus\{e_{k+1}\})$, implying EF1 also holds.
Combining the above cases, it can be concluded that in round $k+1$, the allocation is EF1. This completes the induction and establishes the correctness of our claim.
Secondly, we will show that $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ is MMS. Pick an arbitrary agent $i\in[n-1]$. For agent $i$, assume that the number of items with the value of one that arrive between round $t=1$ and $k$ is $s_{i}^{k}$. By the description of our algorithm, items are almost evenly allocated to agents with identical binary valuations. Thus, it is trivial that $v_{i}(A_{i})=|v_{i}(A_{i})|\geq\mathsf{MMS}_{i}=\lfloor\frac{s_{i}^{k}}{n}\rfloor$. For agent $n$, if she picks item $e_{k}$, by the above proof of EF1, her bundle is EF, implying that her bundle is also MMS; If she does not pick item $e_{k}$, by the above proof of EF1, the value of her bundle is at most $a_{n}$ less than that of the envied agent’s bundle. Considering that she has the priority of picking the item with the value of $b_{n}$ and the number of items in her bundle is at most one less than that of the other agent’s bundle, it is not hard to check that her bundle is still MMS. ∎
Then, we complement the above result by constructing the following example to show the incompatibility of EF1 or MMS, and USW in the combination of additive binary and additive bi-valued valuation functions.
Proposition 2.
For the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation that is EF1 or MMS, and $\alpha$-max-USW for any $\alpha>0$, even for two agents where one agent has the additive binary valuation function, and the other agent has the additive bi-valued valuation function.
Proof.
Consider an online allocation instance with two agents $\{1,2\}$. Assume that, for any item $e$, $v_{1}(e)\in\{0,1\}$ and $v_{2}(e)\in\{\epsilon^{2},\epsilon\}$, where $10<\epsilon<$ 1. When $t=1$, the first item $e_{1}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\epsilon$-max-USW. If not, agent $1$ picks it, and this allocation is EF1, MMS, and max-USW. Then, when $t=2$, the second arriving item $e_{2}$ has the value of 1 for agent 1 and $\epsilon^{2}$ for agent 2. If it is allocated to agent 1, this allocation is 0-EF1, 0-MMS, and max-USW. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\epsilon^{2}+1}{2}\approx\frac{1}{2}$-max-USW. When $t=3$, the item $e_{3}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives. If it is allocated to agent 1, this allocation is $\epsilon$-EF1, $\epsilon$-MMS, and $\frac{2+\epsilon^{2}}{3}\approx\frac{2}{3}$-max-USW. If it is allocated to agent 2, this allocation is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1+\epsilon^{2}+\epsilon}{3}\approx\frac{1}{3}$-max-USW. In other words, no allocation is EF1 or MMS. Combining the above cases, it can be seen that if an online algorithm outputs an NW allocation that is EF1 or MMS in the setting of binary and bi-valued valuations, it may not guarantee $\alpha$-max-USW for any given $\alpha>0$. ∎
4 Allocation of Chores
In this section, we study the setting of indivisible chores. Because “online” chores can be regarded as daily compulsory tasks or duties, it is unreasonable to throw them away. In other words, for each arriving chore, there must be some agent taking it. Therefore, we first introduce the following constraint, and then give the fairness and efficiency definitions in the chores setting.
Definition 7 (Completeness for Chores).
An allocation $A^{k}$ is complete if, for any chore $e$ that has arrived by round $k$, it is not thrown away.
Definition 8 (Envy-freeness up to One Chore).
For any $\alpha\in[1,+\infty)$, an allocation $A^{k}$ is $\alpha$-approximate envy-free up to one chore ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i}^{k}\neq\emptyset$, it holds that $c_{i}(A_{i}^{k}\setminus\{e\})\leq\alpha\cdot c_{i}(A_{i^{\prime}}^{k})$ for some $e\in A_{i}^{k}$.
Definition 9 (Minimax Share Fairness).
For any agent $i$, her minimax share $\mathsf{MMS}_{i}^{k}$ by round $k$ is defined as:
$\mathsf{MMS}_{i}^{k}=\min_{B^{k}\in\Pi_{n}(T^{k})}\max_{i^{\prime}\in N}c_{i}(B_{i^{\prime}}^{k}),$
where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$. For any $\alpha\in[1,+\infty)$, an allocation $A^{k}$ is $\alpha$-approximate minimiax share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $c_{i}(A_{i}^{k})\leq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
Definition 10 (Utilitarian Social Cost).
The utilitarian social cost $USC$ of allocation $A^{k}$ is given by $USC(A^{k})=\sum_{i\in N}c_{i}(A_{i}^{k})$. For any $\alpha\in[1,+\infty)$, an allocation $A^{k}$ is $\alpha$-min-USC if it holds that $USC(A^{k})\leq\alpha\cdot\min_{B^{k}\in\Pi_{n}(T^{k})}USC(B^{k})$, where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
4.1 The Impossibility of Approximate EF1
In this part, we show the following negative result about the approximation of EF1 for tri-valued cost functions. For any agent $i\in N$, her cost function $c_{i}$ is additive tri-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
Theorem 8.
For general additive cost functions, no deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$, even for two agents with additive tri-valued cost functions.
Proof.
$t=1$ $t=2$ $t=3$
agent $1$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Table 6: The impossibility result of EF1 for additive tri-valued cost functions.
Consider an online allocation instance with two agents $\{1,2\}$. Assume that, for any agent $i\in\{1,2\}$ and item $e$, $v_{i}(e)\in\{\epsilon,1,\frac{1}{\epsilon}\}$, where $10<\epsilon<$ 1. The cost of each arrived item is shown in Table 6. When $t=1$, the arriving item has a cost of 1 for both agents. Without loss of generality, agent 1 picks it. When $t=2$, the item that has the cost of $\epsilon$ for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. In order to avoid the approximation ratio of EF1 being infinite, this item must be allocated to agent 2. At last, in round $t=3$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. In this case, if this item is allocated to agent 1, for agent 2, the approximation ratio of EF1 is $\frac{1}{\epsilon}$; if this item is allocated to agent 2, for agent 1, the approximation ratio of EF1 is $\frac{1}{\epsilon}$. Therefore, for any $\alpha>0$, let $\epsilon<\frac{1}{\alpha}$, and we can conclude that no algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any given $\alpha>0$. This impossibility result can be generalized to more than two agents and we defer the proof to Proposition 6. ∎
4.2 Additive/Supermodular Binary Costs
In this part, we consider (additive) binary and supermodular binary cost functions in the chores setting.
4.2.1 Additive Binary Costs
For additive binary cost functions, where for any agent $i\in N$, $c_{i}$ is additive and $c_{i}(e)\in\{0,1\}$ for any $e\in T$, we propose the Compelled-Greedy Algorithm (Algorithm 7) to compute a complete allocation that satisfies EF1, MMS, and min-USC.
Input: An instance $(T,N,(c)_{i\in N})$ with additive binary cost functions
Output: An complete EF1 allocation $A$
1 Initialize: let $\pi=(\pi_{1},\cdots,\pi_{n})$ be an arbitrary order of $n$ agents;
2Let $A=(\emptyset,\ldots,\emptyset)$;
3 when item $e\in T$ arrives do
4
5      for $i=1$ to $n$ do
6
7            if $v_{\pi_{i}}(\{e\})=0$ then
8                   Let $A_{\pi_{i}}=A_{\pi_{i}}\cup\{e\}$;
9                  break;
10             end if
11
12       end for
13      $A_{\pi_{1}}=A_{\pi_{1}}\cup\{e\}$;
14      $\pi=(\pi_{2},\cdots,\pi_{i-1},\pi_{i+1},\cdots,\pi_{n},\pi_{1})$;
15
16
17return $A$;
Algorithm 7 Compelled-Greedy Algorithm
Theorem 9.
For additive binary cost functions, the Compelled-Greedy Algorithm (Algorithm 7) computes a complete allocation that satisfies EF1, MMS, and min-USC.
Proof.
Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, where $c_{i}$ is additive and binary for all $i\in N$. Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ be the set of items that have already arrived, denoted by ${\bf A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $k$. For every item $e\in T$, if there exists one agent such that the cost of $e$ to some agent is 00, the algorithm must allocate $e$ to some agent $a$ with $v_{a}(e)=0$. Otherwise, the costs of all agents are $1$. Regardless of how it is allocated, the increment brought to utilitarian social welfare is $1$, so $A^{k}$ is a min-USC allocation.
Next, we show that ${\bf A^{k}}$ is EF1. Assume that there exists $i,j\in N$ such that, for any item $e\in A^{k}_{i}$,
$c_{i}(A^{k}_{j})\leq c_{i}(A^{k}_{i}\setminus\{e\}).$ (1)
There must exist an item $e^{i}\in A^{k}_{i}$ such that $c_{i}(A^{k}_{i})-c_{i}(A^{k}_{i}\setminus\{e\})=1$. Without loss of generality, we assume $e^{i}$ is the last item with $1$ cost, and it arrives at $\ell$-th round. For any agent, the cost of $e^{i}$ is $1$, and agent $i$’s priority is higher than $j$. Due to the algorithm’s order in each round, the number of items with a $1$ cost of agent $j$ must be at least as large as the number of items with a $1$ cost of agent $j$. It contradicts inequality (1).
Finally, we show that ${\bf A^{k}}$ is MMS. For any agent $j\in N$ and any item $e\in A^{k}_{j}$, if the cost of $c_{j}(\{e\})$ is $1$, $c_{i}(\{e\})=1$ for all $i\in N$. So, $c_{i}(A^{r}_{j})\geq c_{j}(A^{r}_{j})$ for any $r$-th round. $e^{i}$ is the last item with $1$ cost and it arrives at $\ell$-th round, let $T_{\ell}$ be the set arrived after $\ell$-th round. In the $\ell-1$-th round, $c_{i}(A^{\ell}_{i})=c_{i}(A^{\ell-1}_{i})+1\leq\frac{\sum_{j\in N}c_{i}(A^{\ell-1}_{j})}{n}+1=\frac{c_{i}(T_{\ell-1})}{n}+1$. Since $c_{i}(A^{\ell}_{i})$ is an integer, we have $c_{i}(A^{k}_{i})=c_{i}(A^{\ell}_{i})\leq\lceil\frac{c_{i}(T_{\ell})}{n}\rceil\leq\lceil\frac{c_{i}(T)}{n}\rceil=\mathsf{MMS}_{i}$. ∎
4.2.2 Supermodular Binary Costs
For supermodular binary cost functions, where for any agent $i\in N$, $c_{i}$ is supermodular and the marginal cost of any item is 0 or 1, we first obtain a significant negative result regarding the approximation of fairness and efficiency. This contrasts starkly with the findings in offline settings, where it is possible to efficiently obtain an allocation satisfying EF1, MMS, or min-USC [18].
Lemma 3.
(Lemma 2 in [18]) A cost function $c$ is supermodular and binary if and only if the function $g$, $g(S)=|S|-c(S)$ for any set $S$, is a matroid rank function.
Theorem 10.
For supermodular binary cost functions, no deterministic online algorithms can achieve $\alpha$-EF1 and completeness for any $\alpha\geq 1$.
Proof.
Figure 3: An instance with $\frac{1}{0}$-EF1 allocations
Consider an online allocation instance, there are two agents $\{1,2\}$ with supermodular binary cost functions $c_{1}$ and $c_{2}$. Based on the relationship between the supermodular and binary functions and the mathematical rank function in Lemma 10, we construct cost functions for the instance by constructing a matroid rank function. For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $\{C^{i}_{1},\cdots,C^{i}_{l}\}$. In Figure 3, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents. For agent $i\in\{1,2\}$, consider the corresponding partition matroid $\mathcal{M}_{i}=\{T,\mathcal{F}_{i}\}$, $\mathcal{F}_{i}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~{}k\in[l]\}$. Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 00. W.l.o.g., we allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 00. We will discuss the following two cases.
Case 1: we allocate $e_{2}$ to agent $1$. For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 1-1: we allocate $e_{3}$ to agent $1$. When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. To not break the requirement of EF1, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{3},e_{6}\}$, $A^{6}_{2}=\{e_{4},e_{5}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{6}_{1}\setminus\{e\})$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{3}\}$, $A^{6}_{2}=\{e_{4},e_{5},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$. Case 1-2: we allocate $e_{3}$ to agent $2$. When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 1-2-1: we allocate $e_{4}$ to agent $1$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{4},e_{6}\}$, $A^{6}_{2}=\{e_{3},e_{5}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$ for any $e\in A^{5}_{1}$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{4}\}$, $A^{6}_{2}=\{e_{3},e_{5},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$ for any $e\in A^{6}_{2}$. Case 1-2-2: we allocate $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $1$. When the sixth item $e_{6}$ arrives, the marginal utility for agent $1$ and agent $2$ is $1$. If we allocate $e_{6}$ to agent $1$, $A^{6}_{1}=\{e_{1},e_{2},e_{5},e_{6}\}$, $A^{6}_{2}=\{e_{3},e_{4}\}$, then we have $c_{1}(A^{6}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$ for any $e\in A^{5}_{1}$. Else, we allocate $e_{6}$ to agent $2$, $A^{6}_{1}=\{e_{1},e_{2},e_{5}\}$, $A^{6}_{2}=\{e_{3},e_{4},e_{6}\}$, then we have $c_{2}(A^{6}_{1})=0<1=c_{2}(A^{6}_{2}\setminus\{e\})$ for any $e\in A^{6}_{2}$.
Case 2: we allocate $e_{2}$ to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 00. Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases. Case 2-1: we allocate $e_{3}$ to agent $1$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. If we allocate $e_{5}$ to agent $1$, $A^{5}_{1}=\{e_{1},e_{3},e_{5}\}$, $A^{5}_{2}=\{e_{2},e_{4}\}$, then we have $c_{1}(A^{5}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$. Else, we allocate $e_{5}$ to agent $2$, $A^{5}_{1}=\{e_{1},e_{3}\}$, $A^{5}_{2}=\{e_{2},e_{4},e_{5}\}$, then we have $c_{2}(A^{5}_{1})=0<1=c_{2}(A^{5}_{2}\setminus\{e\})$. Case 2-2: we allocate $e_{3}$ to agent $2$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. To avoid breaking the requirement of EF1, we must allocate item $e_{4}$ to agent $1$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. If we allocate $e_{5}$ to agent $1$, $A^{5}_{1}=\{e_{1},e_{4},e_{5}\}$, $A^{5}_{2}=\{e_{2},e_{3}\}$, then we have $c_{1}(A^{5}_{2})=0<1=c_{1}(A^{5}_{1}\setminus\{e\})$. Else, we allocate $e_{5}$ to agent $2$, $A^{5}_{1}=\{e_{1},e_{4}\}$, $A^{5}_{2}=\{e_{2},e_{3},e_{5}\}$, then we have $c_{2}(A^{5}_{1})=0<1=c_{2}(A^{5}_{2}\setminus\{e\})$. ∎
Theorem 11.
For supermodular binary cost functions, no deterministic online algorithms can achieve $\alpha$-MMS or $\alpha$-min-USC, and completeness for any $\alpha\geq 1$.
Proof.
Figure 4: An instance with $\frac{1}{0}$-MMS allocations
Consider an online allocation instance, there are two agents $\{1,2\}$ with supermodular binary cost functions $c_{1},c_{2}$. For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $C^{i}_{1},\cdots,C^{i}_{l}$. In Figure 2, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents. For agent $i\in\{1,2\}$, consider the corresponding partition matroid $\mathcal{M}_{i}=\{T,\mathcal{F}_{i}\}$, $\mathcal{F}_{i}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~{}k\in[l]\}$. Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 00. We allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 00. We will discuss the following two cases.
Case 1: We allocate $e_{2}$ to agent $1$. For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{3}$ to agent $2$. When the fourth item $e_{4}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$.
Case 2: We allocate $e_{2}$ to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 00. Continue to discuss in two sub-cases. Case 2-1: We allocate $e_{3}$ to agent $2$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$. Case 2-2: We allocate $e_{3}$ to agent $1$. For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are 00. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{4}$ to agent $2$. When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 00. In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{5}$ to agent $2$. When the sixth item $e_{6}$ arrives, the marginal utilities for agent $1$ and agent $2$ are $1$. No matter which agent is assigned, the cost of its allocation is $1$, but $\mathsf{MMS}_{1}(T)=\mathsf{MMS}_{2}(T)=0$, $\min USC(T)=0$. ∎
4.3 Additive Bi-valued Costs
In this part, we consider additive bi-valued cost functions.For any agent $i\in N$, her cost function $c_{i}$ is additive bi-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$.Similar to the results of goods setting, even for two agents, an EF1 or MMS allocation cannot be guaranteed.
4.3.1 Two Agents
First, we study the additive bi-valued cost functions for two agents and propose the Adapted Chores Envy-Graph Procedures (Algorithm 8) to return an allocation that achieves the tight approximation guarantee for EF1 and has a good approximation for MMS. Using techniques similar to the case of goods, we take the chores adapted envy-graphs procedures as the base algorithm, and then design two auxiliary algorithms (Algorithms 9 and 10) to remove the possible envy cycle by complicated case analysis to ensure the approximation of EF1 and MMS.
Definition 11 (Envy Graph for Chores).
Given an allocation $A^{k}$, in an envy graph, each vertex corresponds to an agent, and each directed edge $(i,j)$ represents that agent $i$ envies agent $j$, where we have $c_{i}(A_{i}^{k})>c_{i}(A_{j}^{k})$.
Input: An instance $(T,N,(c_{1},c_{2}))$ with additive bi-valued cost functions
Output: A complete approximate EF1 and MMS allocation $A$
1
2Let $A=(\emptyset,\ldots,\emptyset)$ and build an envy graph $G$ for two agents;
3 when item $e\in T$ arrivesdo
4
5      if there is no envy between the two agents then
6             if $\forall k\in N$, $c_{k}(e)=a_{k}$ or $\forall k\in N$, $c_{k}(e)=b_{k}$  then
7                   Arbitrarily choose one agent $i\in N$;
8            else
9                   Choosing the agent $i\in N$ who values it at $a_{i}$;
10             end if
11
12      else
13             Let $i\in N$ be an agent who does not envy the other agent and $j=2-i$;
14            if $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ then
15                   Terminate this algorithm and implement the chores preliminary-breaking-cycle algorithm (Algorithm 9);
16
17       end if
18      $A_{i}=A_{i}\cup\{e\}$;
19
return $A$;
Algorithm 8 Adapted Chores Envy-Graph Procedures
Input: An instance $(T,N,(c_{1},c_{2}))$ with additive bi-valued cost functions
Output: A complete approximate EF1 and MMS allocation $A$
1
2Initialize: $\lambda=1$ and $\mu=0$;
3 when item $e\in T$ arrives do
4
5      if $\lambda=1$ then
6
7            if $|A_{j}|=1$ then
8
9                  $A_{j}=A_{j}\cup\{e\}$;
10                  Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
11            else
12
13                  if $c_{i}(e)=b_{i}$ then
14
15                        $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
16                  else
17                         $A_{i}=A_{i}\cup\{e\}$ and $\mu=2$;
18                   end if
19                  $\lambda=2$;
20             end if
21
22      else
23
24            if $\mu=1$  then
25
26                  $A_{i}=A_{i}\cup\{e\}$;
27            else
28
29                  if $c_{j}(e)=a_{j}$ then
30
31                        $A_{j}=A_{j}\cup\{e\}$;
32                  else
33
34                        $A_{i}=A_{i}\cup\{e\}$;
35                        Terminate this algorithm and execute the chores deep-breaking-cycle algorithm (Algorithm 10);
36                   end if
37
38             end if
39
40       end if
41      Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
42
return $A$;
Algorithm 9 Chores Preliminary-Breaking-Cycle Algorithm
Input: An instance $(T,N,(c_{1},c_{2}))$ with additive bi-valued cost functions
Output: A complete approximate EF1 and MMS allocation $A$
1 Initialize: $\lambda=1$ and $\mu=0$;
2 when item $e\in T$ arrives do
3
4      if $\lambda=1$ then
5
6            if $c_{j}(e)=a_{j}$ and $c_{i}(e)=b_{i}$ then
7                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
8             else if $c_{j}(e)=b_{j}$ and $c_{i}(e)=b_{i}$ then
9
10                  $A_{j}=A_{j}\cup\{e\}$ and $\mu=2$;
11             else if  $c_{j}(e)=a_{j}$ and $c_{i}(e)=a_{i}$  then
12                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=3$;
13             else
14                   $A_{i}=A_{i}\cup\{e\}$ and $\mu=4$;
15             end if
16            $\lambda=2$;
17      else
18             if $\mu=1$ then
19                   $A_{j}=A_{j}\cup\{e\}$;
20             else if  $\mu=2$  then
21                   if $c_{j}(e)=b_{j}$ then
22
23                        $A_{i}=A_{i}\cup\{e\}$;
24                        $\lambda=1$;
25                  else
26                         $A_{j}=A_{j}\cup\{e\}$;
27                   end if
28
29             else if  $\mu=3$  then
30                   if $c_{i}(e)=b_{i}$ then
31
32                        $A_{j}=A_{j}\cup\{e\}$;
33                  else
34
35                        $A_{i}=A_{i}\cup\{e\}$;
36                        $\lambda=1$;
37                   end if
38
39             else
40
41                  $A_{j}=A_{j}\cup\{e\}$;
42                  $\lambda=1$;
43             end if
44
45            if $\lambda=2$ then
46                   Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
47
48       end if
49
50
51return $A$;
Algorithm 10 Chores Deep-Breaking-Cycle Algorithm
Theorem 12.
For the deterministic allocation of indivisible chores for two agents with additive bi-valued cost functions, the Adapted Chores Envy-Graph Procedures (Algorithm 8) computes a complete allocation that satisfies $2$-EF1 and $\frac{5}{3}$-MMS.
Similar to the goods setting, we first show the following key observations and lemmas before the proof of Theorem 12.
Observation 4.
When Algorithm 9 is executed, for the first arriving item $e$, we have $c_{j}(e)=a_{j}$.
Proof.
When Algorithm 9 is ready to be executed, it means that if the arriving item $e$ is allocated to agent $i$, there will be an envy cycle, i.e, $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ and $c_{i}(A_{i}\cup\{e\})>c_{i}(A_{j})$ For agent $j$, her bundle is EF1. Suppose that $c_{j}(e)=b_{j}$. Then, we have $c_{j}(A_{j})\leq c_{j}(A_{i})+c_{j}(f)\leq c_{j}(A_{i})+b_{j}=c_{j}(A_{i}\cup\{e\})$ for some $f\in A_{j}$, which is a contradiction. Therefore, it must hold that $v_{j}(e)=a_{j}$. ∎
Observation 5.
When Algorithm 9 is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
Proof.
Suppose that $|A_{i}|=0$. Then, we have $|A_{j}|\geq 1$ since agent $j$ envies agent $i$. After allocating item $e$ to agent $i$, they envies each other, i.e, $c_{i}(e)>c_{i}(A_{j})$ and $c_{j}(A_{j})>c_{j}(e)$. In that case, we derive that $c_{j}(A_{j})=b_{j}$ and $c_{i}(A_{j})=a_{i}$, which contradicts the description of our algorithm. Suppose that $|A_{j}|=0$. Then, we have $|A_{i}|=0$. Otherwise, agent $j$ will not envy agent $i$, after allocating item $e$ to agent $i$. In that case, the envy cycle will not appear. Thus, it hold that $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$. ∎
Lemma 4.
Algorithm 9 computes a complete allocation that is EF1 with or without an envy cycle, or 2-EF1.
Proof.
When Algorithm 9 is executed, the input allocation $A=(A_{1},A_{2})$ is EF1. According to the description of our algorithm, it is executed in at most two rounds. Then, let us consider different rounds.
$\lambda=1$. When the first item $e_{1}$ arrives, we have $c_{j}(e_{1})=a_{j}$ by Observation 4. When $|A_{j}|=1$, we have $c_{j}(A_{j})=b_{j}$ since she envies agent $i$. Then, this item is allocated to agent $j$. For agent $i$, her bundle is EF since she does not envy agent $j$. For agent $j$, her new bundle is EF1 since $|A_{j}|=1$ and $c_{j}(e_{1})=a_{j}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. When $|A_{j}|>1$, let us consider two subcases about the valuation of $e_{1}$.
  • If $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\mu=1$. For agent $j$, her new bundle is $2$-EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(A_{i})\geq c_{j}(e_{1})=a_{j}$. For agent $i$, her bundle is EF since she does not envy agent $j$. Thus, this allocation is $2$-EF1.
  • If $c_{j}(e_{1})=a_{i}$, agent $i$ picks it and we set $\mu=2$. For agent $i$, her new bundle is EF1 since after removing $e_{1}$ from her bundle, we have $c_{j}(A_{j})\leq c_{j}(A_{i})$. For agent $j$, her bundle is EF1. Thus, this allocation is EF1 with an envy cycle.
$\lambda=2$. When the second item $e_{2}$ arrives, let us consider two cases about $\mu$.
  • If $\mu=1$, agent $i$ picks it. For agent $i$, her new bundle is EF since we have $c_{i}(A_{i}\cup\{e_{2}\})\leq c_{i}(A_{i})+b_{i}\leq c_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her bundle is EF1 since we have $v_{j}(A_{j}\setminus\{f\})\leq v_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(e_{2})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • If $\mu=2$, let us consider two subcases about the cost of $e_{2}$. When $c_{j}(e_{2})=a_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i})\leq c_{i}(A_{j})$ and $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{2})=a_{j}\leq c_{j}(e_{1})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. When $c_{j}(e_{2})=b_{j}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i})\leq c_{i}(A_{j})$ and $c_{i}(e_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her bundle is EF1 since she does not pick any item in these two rounds. Thus, this allocation is 2-EF1, and the algorithm is terminated.
Combining the above cases together, it is not hard to see that the claim holds. ∎
Observation 6.
When Algorithm 10 is executed, there are two properties about allocation $A$:
  • $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$;
  • Let $e_{1}$ and $e_{2}$ denote the two new allocated items in Algorithm 9, we have $c_{i}(A_{i}\setminus\{e\}\leq 2c_{i}(A_{j})$, where $e\in\arg\max_{f\in\{e_{1},e_{2}\}}c_{i}(f)$.
Proof.
When Algorithm 10 is executed, in Algorithm 9 we have $\lambda=2$ and $\mu=2$. By the proof of Lemma 4, it is easy to check the statements are correct. ∎
Lemma 5.
Algorithm 10 computes a complete allocation that is EF1 with or without an envy cycle, or 2-EF1.
Proof.
When Algorithm 10 is executed, by Observation 6, the input allocation $A=(A_{1},A_{2})$ is 2-EF1, where $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$ and $c_{i}(A_{i}\setminus\{e\})\leq 2c_{i}(A_{j})$ for some $e\in A_{i}$. Let $f_{1}$ and $f_{2}$ denote the two online items that are allocated to agent $j$ in Algorithm 9 and we have $\min\{c_{j}(e_{1}),c_{j}(e_{2})\}=a_{j}$.
$\lambda=1$. When the item $e_{1}$ arrives at $\lambda=1$, let us consider four cases of agent’s costs functions.
  • When $c_{j}(e_{1})=a_{j}$ and $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\mu=1$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(f_{1})$. Thus, this allocation is EF1 with an envy cycle.
  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\mu=2$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=b_{j}\leq c_{j}(f_{2})$. Thus, this allocation is EF1 with an envy cycle.
  • When $c_{j}(e_{1})$ and $c_{i}(e_{1})=a_{i}$, agent $j$ picks it and we set $\mu=3$. For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$. For agent $j$, her new bundle is 2-EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{i}\leq c_{j}(f_{1})$. Thus, this allocation is EF1 with an envy cycle.
  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=a_{j}$, agent $i$ picks it and we set $\mu=4$. For agent $i$, her new bundle is still 2-EF1 since we have $\frac{c_{i}(A_{i}\setminus\{f\}\cup\{e_{1}\})}{c_{i}(A_{j})}\leq\frac{c_{i}(A_{j})+c_{i}(e_{1})}{c_{i}(A_{j})}\leq 2$ for some $f\in\{f_{1},f_{2}\}$ where the inequality follows from the fact that $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}$ and $|A_{j}|\geq 2$. Thus, this allocation is 2-EF1.
$\lambda=2$. When the item $e_{2}$ arrives at $\lambda=2$, let us consider four cases about $\mu$.
  • When $\mu=1$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{2})\leq c_{i}(e_{1})=b_{i}$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$, $c_{j}(e_{1})=a_{j}$ and $c_{j}(e_{2})\leq c_{j}(f_{2})=b_{j}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • When $\mu=2$, we need to consider two subcases. If $c_{j}(e_{2})=b_{j}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$ and $c_{j}(e_{2})\leq c_{i}(e_{1})=b_{i}$. For agent $j$, her bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{1})=b_{j}=c_{j}(e_{2})$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda=1$ and continue executing this algorithm. If $c_{j}(e_{2})=a_{j}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{2})\leq c_{i}(e_{1})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{j}(e_{1})=b_{j}=c_{j}(f_{2})$ and $c_{j}(e_{2})=a_{j}=c_{j}(f_{1})$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
  • When $\mu=3$, let us consider two subcases. If $c_{i}(e_{2})=b_{i}$, agent $j$ picks it. For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}=c_{i}(e_{1})$ and $c_{i}(f_{2})=b_{i}=c_{i}(e_{2})$. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{j}(e_{1})=a_{j}=c_{j}(f_{1})$ and $c_{j}(e_{2})\leq c_{j}(f_{2})=b_{i}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. If $c_{i}(e_{2})=a_{i}$, agent $i$ picks it. For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(e_{2})=a_{i}=c_{i}(e_{1})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(e_{2})$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda=1$ and continue executing this algorithm.
  • When $\mu=4$, agent $j$ picks it, For agent $i$, her bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$. For agent $j$, her new bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{2})\leq c_{j}(e_{1})=b_{j}$. Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6. Additionally, we set $\lambda=1$ and continue executing this algorithm.
Combining the above cases, it can be seen that our statement holds. ∎
Proof of Theorem 12.
In Algorithm 8, we follow a similar way of offline envy graph procedures to allocate items. The precondition of Algorithm 8 is that there is no envy cycle between two agents. Initially, the empty allocation is EF1 without an envy cycle. When there is no envy between two agents, we allocate the arriving item to one agent according to different conditions. When there exists an envy between two agents, we have two cases. One is that we allocate it to the unenvied agent when the envy cycle does not appear. The other is that if we allocate it to the unenvied agent, the envy cycle will appear, and we terminate Algorithm 8. Similar to the allocation of goods, we need to ensure that there is no envy cycle between two agents. Because we can not revoke the previous allocation in the online setting, Algorithms 9 and 10 help us eliminate the possible envy cycle. When the allocation $A$ satisfies the precondition and is EF1, we will continue to execute Algorithm 8. By the induction, it is not hard to see that the output allocation of Algorithm 8 is always EF1 without an envy cycle. Combining Lemmas 4, 5, and the description of Algorithm 8, we can conclude that the output complete allocation is $2$-EF1. For MMS, by the implication between EF1 and MMS [22], i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{n\alpha+n-1}{n-1+\alpha}$-MMS, where $n$ is the number of agents, it is easy to see that the output allocation is $\frac{5}{3}$-MMS. ∎
Theorem 13.
For two agents with additive bi-valued cost functions, no deterministic online algorithm can compute a complete allocation that is $(2-\epsilon)$-EF1 or $(\frac{3}{2}-\epsilon)$-MMS for any $\epsilon>0$.
Proof.
$t=1$ $t=2$ $t=3$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 1
$t=1$ $t=2$ $t=3$ $t=4$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$ $1$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 2.2
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$ $1$ $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 2.3
Table 7: The impossibility results of EF1 for bi-valued cost functions for two agents.
$t=1$ $t=2$
agent $1$ $1$ $1$
agent 2 $1$ $1$
Case 1
$t=1$ $t=2$ $t=3$
agent $1$ $1$ $1$ 2
agent 2 $1$ $1$ $2$
Case 2
Table 8: The impossibility results of MMS for additive bi-valued cost functions for two agents.
Consider an online allocation instance with two agents $\{1,2\}$. For any agent $i\in\{1,2\}$ and any item $e$, $c_{i}(e)\in\{1,\frac{1}{\epsilon}\}$, where $10<\epsilon<$ 1. The cost of each arrived item is shown in Table 9. To guarantee the completeness of chores allocation, we should guarantee that each arriving item is allocated to one agent.
The impossibility of EF1.
When $t=1$, the arriving item has a cost of 1 for both agents. Without loss of generality, assume that agent 1 picks it. When $t=2$, the item that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. To avoid the approximation ratio of EF1 being infinity, it is allocated to agent 2. Then, it is easy to check that the allocation is EF1 and MMS. In round $t=3$, the item that has the cost of $\frac{1}{\epsilon}$ for both agents arrives. Then, let us consider two cases.
Case 1. If it is allocated to agent 2, no further item will appear. Thus, the allocation is $\frac{1}{\epsilon}$-EF1.
Case 2. If it is allocated to agent 1, it is clear that this allocation is EF1. Next, when $t=4$, the arriving item has the cost of $1$ for agent 1 and $\frac{1}{\epsilon}$ for agent 2. Then, let us consider two subcases. If agent 1 picks it, no further item will appear. So this allocation is trvially $(1+\frac{1}{\epsilon})$-EF1. If agent $2$ picks it, the allocation is EF1. At last, in round $t=5$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. If it is allocated to agent 1, the allocation is $(\frac{1}{2\epsilon}+\frac{1}{2})$-EF1. If it is allocated to agent 2, the allocation is $\frac{2}{1+\epsilon}$-EF1.
Combining the above cases together, we can derive that the lower bound of EF1 is 2.
The impossibility of MMS.
The constructed instance is below. Let $a_{1}=a_{2}=1$, $b_{1}=b_{2}=2$. When $t=1$, the arriving item has the cost of 1. Assume that it is allocated to agent 1. Then, in round $t=2$, the item that has the cost of 1 for both agents arrives. If it is allocated to agent 1, no further allocation will appear. Thus, this allocation is 2-MMS. If it is allocated to agent 2, this allocation is MMS. In round $t=3$, the arriving item has the cost of 2 for both agents. No matter which agent picks it, the final allocation is $\frac{3}{2}$-MMS. Combining the above cases together, we can derive that the lower bound of MMS is $\frac{3}{2}$. ∎
Proposition 3.
For the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation that is $\alpha$-EF1 or $\beta$-MMS, and $\gamma$-min-USC for any $\alpha,\gamma\geq 1$ and $\beta<2$, even for two agents with additive bi-valued cost functions.
Proof.
Consider an online allocation instance with two agents $\{1,2\}$, $c_{1}\in\{1,2\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $10<\epsilon<$ 1. When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\frac{1}{\epsilon}\rightarrow\infty$-min-USC. If not, agent $1$ picks it, and this allocation is EF1, MMS, and min-USC. Then, when $t=2$, the second arriving item $e_{2}$ has the cost of 1 for agent 1 and $\frac{1}{\epsilon^{2}}$ for agent 2. If it is allocated to agent 1, this allocation is $\infty$-EF1, 2-MMS, and min-USC. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+\epsilon}{\epsilon+1}\rightarrow\infty$-min-USC. Combining the above cases, it can be seen that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha\geq 1$ or $\beta$-MMS for any $\beta<2$, and $\gamma$-min-USC for any given $\gamma\geq 1$ when two agents have the bi-valued cost functions. ∎
4.3.2 The Combination with Additive Binary Costs
Next, we consider the setting where one agent (agent $n$) has the additive bi-valued cost function, and the remaining agents have additive binary cost functions and propose the Adapted-Chores-Picking Algorithm (Algorithm 11) to compute a complete allocation that satisfies EF1 and MMS. The intuition of this algorithm is similar to the goods setting. In Table 9, we illustrate how our algorithm works for three agents. Regarding efficiency, we show that EF1 or MMS is incompatible with USC in this setting.
-Input: An instance $(T,N,(c_{i})_{i\in N})$ with $(n-1)$ additive binary cost functions $(c_{1}.\ldots,c_{n-1})$ and one additive bi-valued cost function $(c_{n})$.
Output: A complete EF1 and MMS allocation $A$
1
2Let $A=(\emptyset,\ldots,\emptyset)$;
3 when item $e\in T$ arrives do
4
5      Let $i_{min}\in\arg\min_{i\in[n-1]}|\{f~{}|f\in A_{i}~{}and~{}c_{i}(f)>0\}|$;
6      if $\exists i^{*}\in N$ such that $c_{i^{*}}(e)=0$ then
7
8            $A_{i^{*}}=A_{i^{*}}\cup\{e\}$;
9      else
10
11            if $c_{n}(e)=a_{n}$ then
12
13                  if |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\geq|A_{n}|$ then
14                         $A_{n}=A_{n}\cup\{e\}$;
15                  else
16
17                        $A_{i_{min}}=A_{i_{min}}\cup\{e\}$;
18                   end if
19
20            else
21                   if |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\leq|A_{n}|$ then
22
23                        $A_{i_{min}}=A_{i_{min}}\cup\{e\}$;
24                  else
25                         $A_{n}=A_{n}\cup\{e\}$;
26                   end if
27
28             end if
29
30       end if
31
32
return $A$;
Algorithm 11 Adapted-Chores-Picking Algorithm
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$ldots
agent $1$ $1$ 00 $1$ 1 $1$ldots
agent $2$ $1$ 1 1 $1$ 1 $\ldots$
agent 3 $a_{3}$ $a_{3}$ $a_{3}$ $b_{3}$ $b_{3}$ldots
Table 9: A simple illustration for three agents of Algorithm 11.
Theorem 14.
For the deterministic allocation of indivisible chores for $(n-1)$ agents with additive binary cost functions and one agent with the additive bi-valued cost function, the Adapted-Chores-Picking Algorithm (Algorithm 11) computes a complete allocation that satisfies EF1 and MMS.
Proof.
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 11.
Firstly, we show that $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ is EF1 by induction. For the base case, the empty allocation is trivially EF1. For the induction step, after the round $k$, the allocation $A^{k}$ is EF1. We show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is also EF1.
  • If there exists an agent $i\in N$ such that $c_{i}(e_{k+1})=0$, agent $i$ picks it. It is clear that this allocation is still EF1.
  • If $c_{n}(e_{k+1})=a_{n}$, then let us consider two subcases. When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\geq|A_{n}|$, agent $n$ picks it. For agent $i\in[n-1]$, her bundle is still EF1. For agent $n$, her bundle is EF1 since she has the priority of picking the item with the cost of $a_{n}$ and the number of items in her bundle is no more than that of items with non-zero costs in agent $i_{min}$’s bundle. When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|<|A_{n}|$, agent $i_{min}$ picks it. For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1. For agent $i_{min}$, her bundle is EF1 since she has the binary cost valuation and the lowest number of items with non-zero costs.
  • If $v_{n}(e_{k+1})=b_{n}$, the let us consider two subcases. When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\leq|A_{n}|$, agent $i_{min}$ picks it. For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1. For agent $i_{min}$, her bundle is EF1 since she has the binary cost valuation and the lowest number of items with non-zero costs. For any agent $i\in N\setminus\{i_{min}\}$, her bundle remains untouched and is still EF1. When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|>|A_{n}|$, agent $n$ picks it. For agent $i\in[n-1]$, her bundle is trivially EF1. For agent $n$, her bundle is EF since she has the priority of picking the item with the cost of $a_{n}$, and in this case, the extra one item in another agent’s bundle must have the cost of $b_{n}$ for agent $n$.
Combining the above cases, it can be seen that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is still EF1.
Next, we will show that $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ is MMS. For arbitrary agent $i\in N$, if her bundle is EF, we are done. If not, then let us consider two cases. When $i\in[n-1]$, she has a binary valuation function, and by our algorithm description, the number of items with non-zero costs in her bundle is at most one more than that of the other agent’s bundle. Thus, it is not hard to see that her bundle is MMS. When $i=n$, currently, the number of items in her bundle is one more than that of the other agent’s bundle, and the extra item has the cost of $a_{n}$. After removing this item, the cost of her bundle is no more than that of the other agent’s bundle in her perspective, so we can conclude that her bundle is MMS. ∎
Proposition 4.
For the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation that is $\alpha$-EF1 or MMS, and $\beta$-min-USC for any $\alpha,\beta\geq 1$, even for two agents where one agent has the additive binary cost function and the other agent has the additive bi-valued cost function.
Proof.
Without loss of generality, we assume that $c_{1}\in\{0,1\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $10<\epsilon<$ 1. When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives. If it is allocated to agent 2, no further items arrive. Hence, this allocation is clearly EF1, MMS, and $\frac{1}{\epsilon}\rightarrow\infty$-min-USC. If not, agent $1$ picks it, and this allocation is EF1, MMS, and min-USC. Then, when $t=2$, the second arriving item $e_{2}$ has the cost of 1 for agent 1 and $\frac{1}{\epsilon^{2}}$ for agent 2. If it is allocated to agent 1, this allocation is $\infty$-EF1, 2-MMS, and min-USC. If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+1}{2}=\frac{1}{2\epsilon^{2}}+\frac{1}{2}\rightarrow\infty$-min-USC. Combining the above cases, it can be seen that no online algorithm outputs a complete allocation that is $\alpha$-EF1 or MMS in the setting of binary and bi-valued cost functions, and guarantees $\beta$-min-USC for any given $\alpha,\beta\geq 1$. ∎
5 Conclusion and Future Work
In our paper, we study the fair online allocation of indivisible items and present a series of possible and impossible results regarding the existence and approximation results in fairness and efficiency with various valuation or cost functions, which give a picture of what can be achieved. For future work, one interesting direction is to fill the gap of the approximation of EF1 and MMS for bi-valued valuation or cost functions when there are more than three agents. Additionally, based on our results in other online settings, it would also be interesting to study the items with deadline setting in a more complicated valuation or cost function setting.
Appendix A Missing Proofs
A.1 Section 3
A.1.1 Proof of Proposition 5
Proposition 5.
For general additive functions, no deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any $\alpha>0$ when there are more than two agents with additive tri-valued valuation functions.
Proof.
$t=1$ $t=2$ldots $t=n-2$ $t=n-1$ $t=n$ $t=n+1$
agent $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
agent $n-2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $n-1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
More than two agents
Table 10: The impossibility results of EF1 and MMS for additive tri-valued valuations.
We further extend the impossibility results of approximate EF1 and MMS for $n\geq 3$. Consider the following instance with $n\geq 3$ agents, where the value of each item is shown in Table 10. In the first $(n-2)$ rounds, to avoid 0-EF1 or 0-MMS, each item is allocated to the agent with the empty bundle. For the item that arrives before round $t=n-2$, it has the value of $\frac{1}{\epsilon}$ for any agent. Without loss of generality, we assume that agent $i$ picks the item in the round $t=i$, where $i\leq n$. In round $t=n-1$, the arriving item has the value of 1 for any agent, and we assume that it is allocated to agent $n-1$. In round $t=n$, the arriving item has the value of $\epsilon$ for agent $i\in N\setminus\{n-1\}$ and $\frac{1}{\epsilon}$ for agent $n-1$, and it should be allocated to agent $n$ to avoid 0-EF1 or 0-MMS. Finally, in the round $t=n+1$, the item with the value of $\frac{1}{\epsilon}$ for each agent arrives. Due to non-wastefulness, this item must be allocated to one agent. If this item is allocated to agent $i\in N\setminus\{n-1,n\}$, it is not hard to see that the allocation is $\epsilon^{2}$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If this item is allocated to agent $n-1$, this allocation is $\epsilon$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS. If this item is allocated to agent $n$, we can derive that the competitive ratio of EF1 is $\epsilon$ and MMS is $\epsilon$. Combining the above cases together, we can conclude that the upper bound of EF1 is $\epsilon$ and MMS is $\epsilon$. For given $\alpha>0$, let $\epsilon<\alpha$, we can conclude that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS for any given $\alpha>0$. ∎
A.2 Section 4
A.2.1 Proposition 6
Proposition 6.
For general additive cost functions, no deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$ for more than two agents with additive tri-valued cost functions.
Proof.
$t=1$ $t=2$ldots $t=n-2$ $t=n-1$ $t=n$ $t=n+1$
agent $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
agent $n-2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $n-1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
More than two agents
Table 11: The impossibility result of EF1 for additive tri-valued cost functions.
We extend the impossibility result of approximate EF1 for chores to $n\geq 3$ agents. Consider the following instance with $n\geq 3$ agents, where the cost of each item is shown in Table 11. Before each agent’s bundle has one agent, we should allocate the arriving item to the agent with an empty bundle to avoid the approximation ratio of EF1 being infinite. For $t\leq n-2$, each arriving item has the cost of $\frac{1}{\epsilon}$ for each agent. Without loss of generality, assume that agent $i\in[n-2]$ picks the item that arrives at round $i$. When $t=n-1$, the arriving item has the cost of 1 for all agents. Assume that it is allocated to agent $n-1$. When $t=n$, the arriving item has the cost of $\frac{1}{\epsilon}$ for agent $n$ and $\epsilon$ for the remaining agents. It is clear that agent $n$ picks it. In the last round $t=n+1$, the item that has the cost of $\frac{1}{\epsilon}$ for all agents arrives. If it is allocated to agent $i\in[n-2]$, we can derive that the approximation ratio of EF1 is $\frac{1}{\epsilon}$. If it is allocated to agent $n-1$ or $n$, the allocation is also $\frac{1}{\epsilon}$-EF1. Thus, the allocation in this round is $\frac{1}{\epsilon}$-EF1. Given any $\alpha>1$, let $\epsilon>\frac{1}{\alpha}$, we can conclude that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha>1$. ∎
Appendix B Other Online Settings
Online algorithms face two roadblocks: unknown future information and the inability to revoke previous decisions. A natural question is how to avoid the worst-case scenario in online algorithms. In this section, we study two special settings. One is that each arriving item can wait for one period and does not need to be allocated immediately, which can be found in the online matching literature [23]. In this setting, we prove that an EF1 allocation for two agents with bi-valued valuation/cost functions exists. The other one is knowing some information in advance, which originates from the online MMS allocation with normalized valuations [5]. In this setting, we show that the Marginal-Greedy Algorithm computes a non-wasteful EF1 allocation when items are goods, and the Round-Robin algorithm computes a complete EF1 allocation when items are chores.
B.1 Items with Deadline
In this part, each item has a deadline $d$ after its arrival. After $d$ units of time, we must decide how to allocate it. For two agents with bi-valued valuation functions in the goods setting, we design the Adapted Matching Algorithm (Algorithm 12) to compute an NW allocation that satisfies EF1. For two agents with bi-valued cost functions in the chores setting, we design the Adapted Chores Matching Algorithm (Algorithm 13) to compute a complete allocation that satisfies EF1. The common intuition of the above two algorithms is that we find a perfect matching between two agents and two items after two consecutive items arrive so that we can guarantee that there is no envy cycle between two agents and EF1 always holds.
Theorem 15.
For the deterministic allocation of indivisible goods for two agents with bi-valued valuation functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Matching Algorithm (Algorithm 12) computes an NW allocation that satisfies EF1.
Lemma 6.
In Algorithm 12, for the allocation $A$ after allocating two consecutive online items $e_{1}$ and $e_{2}$, we have the following two properties for any agent $i\in N$:
  • If $v_{i}(A_{i})<v_{i}(A_{j})$, then $v_{i}(A_{i})-v_{i}(A_{j})=a_{i}-b_{i}$;
  • If $v_{i}(A_{i})>v_{i}(A_{j})$, then $v_{i}(A_{i})-v_{i}(A_{j})\geq b_{i}-a_{i}$.
Proof.
We show the claim by induction. For the base case, the statement trivially holds. For the induction step, after allocating items $e_{k-1}$ and $e_{k}$, the two properties hold. Without loss of generality, let $A$ denote the allocation after allocating items $e_{k-1}$ and $e_{k}$. Then, we show that allocating items $e_{k+1}$ and $e_{k+2}$, our claim also holds. Fix an arbitrary agent $i\in N$. Now, consider two cases about her.
  • $v_{i}(A_{i})\geq v_{i}(A_{j})$. Because one agent picks exactly one item, it is clear that after allocating items $e_{k+1}$ and $e_{k+2}$, the claim holds since $v_{i}(e_{k+1}),v_{i}(e_{k+2})\in\{a_{i},b_{i}\}$.
  • $v_{i}(A_{i})<v_{i}(A_{j})$. In this case, according to our algorithm description, she picks one item with a value that is greater than or equal to that of the other one. Suppose that she picks $e_{k+1}$. If $v_{i}(e_{k+1})=v_{i}(e_{k+2})$, then we have $v_{i}(A_{i}\cup\{e_{k+1}\})-v_{i}(A_{j}\cup\{e_{k+2}\})=a_{i}-b_{i}+v_{i}(e_{k+1})-v_{i}(e_{k+2})=a_{i}-b_{i}$. If $v_{i}(e_{k+1})>v_{i}(e_{k+2})$, then we have $v_{i}(A_{i}\cup\{e_{k+1}\})-v_{i}(A_{j}\cup\{e_{k+2}\})=a_{i}-b_{i}+v_{i}(e_{k+1})-v_{i}(e_{k+2})=a_{i}-b_{i}+b_{i}-a_{i}=0$.
This completes the induction and establishes the correctness of our claim. ∎
Proof of Theorem 15.
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 12. We will show that the returned allocation $A$ is always EF1 by induction. For the base case, the empty bundle is trivially EF1 without an envy cycle. For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $A^{t}$ is EF1 without an envy cycle. Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is EF1 without an envy cycle. Let us consider three cases about item $e_{k+1}$.
  • $\lambda=1$ and item $e_{k+1}$ is the last arriving item. When envy exists between two agents, assume that agent $i$ is the agent who envies agent $j=2-i$. If she picks $e_{k+1}$, we are done. If not, agent $j$ picks it and we get $v_{i}(e_{k+1})=a_{i}$. For agent $j$, she still does not envy agent $i$, and her new bundle is EF. For agent $i$, by Lemma 6, we have $v_{i}(A_{i}^{k})-v_{i}(A_{j}^{k})=a_{i}-b_{i}$. Then, we have $v_{i}(A_{i}^{k})-v_{i}(A_{j}^{k}\cup\{e_{k+1}\}\setminus\{g\})=a_{i}-b_{i}-a_{i}+b_{i}=0$ where $A_{i}^{k}$ is EF1 and there must exists an item $g\in A_{j}$ such that $v_{i}(g)=b_{i}$, implying the allocation $A^{k+1}$ is EF1 without an envy cycle. When no envy exists between two agents, allocating $e_{k+1}$ to any agent will not violate EF1.
  • $\lambda=2$ and item $e_{k+1}$ needs to be allocated firstly. When envy exists between two agents, assume that agent $i$ is the agent who envies agent $j=2-i$. If agent $i$ picks $e_{k+1}$, we are done. If not, for agent $i$, her bundle is still EF1 since we have $v_{i}(A_{i})-v_{i}(A_{j}\cup\{e_{k+1}\}\setminus\{g\})=a_{i}-b_{i}-a_{i}+b_{i}=0$ where there must exist $g\in A_{j}$ such that $v_{i}(g)=b_{i}$. For agent $j$, she does not envy agent $i$ and by Lemma 6, we have $v_{j}(A_{j})-v_{j}(A_{i})\geq b_{j}-a_{j}$. Thus, her bundle is EF. When no envy exists between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
  • $\lambda=2$ and item $e_{k+1}$ needs to be allocated secondly. When envy exists between two agents, assume that agent $i$ is the agent who envies agent $j=2-i$. If she picks it, we are done. If not, it means that item $e_{k}$ is allocated to agent $i$ and $v_{i}(e_{k})\geq v_{i}(e_{k+1})$. For agent $i$, considering that $A_{i}^{k-1}$ is EF1, we can derive that her bundle is still EF1. For agent $j$, it is trivial that her new bundle is EF. When no envy exists between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
Combining the above cases together, we can conclude that the allocation $A^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
Theorem 16.
For the deterministic allocation of indivisible chores for two agents with bi-valued cost functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Chores Matching algorithm (Algorithm 13) computes a complete allocation that satisfies EF1.
Lemma 7.
In Algorithm 13, for the allocation $A$ after allocating two consecutive online items $e_{1}$ and $e_{2}$, we have the following two properties for any agent $i\in N$:
  • If $c_{i}(A_{i})<c_{i}(A_{j})$, then $c_{i}(A_{i})-c_{i}(A_{j})\leq a_{i}-b_{i}$;
  • If $c_{i}(A_{i})>c_{i}(A_{j})$, then $c_{i}(A_{i})-c_{i}(A_{j})=b_{i}-a_{i}$.
Proof.
We show the claim by induction. For the base case, the statement trivially holds. For the induction step, after allocating items $e_{k-1}$ and $e_{k}$, the two properties hold. Let $A$ denote such an allocation. Then, we show that after allocating items $e_{k+1}$ and $e_{k+2}$, our claim also holds. Fix an arbitrary agent $i\in N$. Now, consider two cases about her.
  • $c_{i}(A_{i})<c_{i}(A_{j})$. After allocating two items $e_{k+1}$ and $e_{k+2}$, the two properties hold since agent $i$ only picks one item and has a bi-valued cost function i.e., $c_{i}(e_{k+1}),c_{i}(e_{k+2})\in\{a_{i},b_{i}\}$.
  • $c_{i}(A_{i})>c_{i}(A_{j})$. By our algorithm description, she picks one item with a cost that is lower than or equal to the other one. Suppose that she picks $e_{k+1}$. If $c_{i}(e_{1})=c_{i}(e_{k+2})$, we have $c_{i}(A_{i}\cup\{e_{k+1}\})-c_{i}(A_{j}\cup\{e_{k+2}\})=b_{i}-a_{i}$. If $c_{i}(e_{1})<c_{i}(e_{2})$, we have $c_{i}(A_{i}\cup\{e_{k+1}\})-c_{i}(A_{j}\cup\{e_{k+2}\})=b_{i}-a_{i}+a_{i}-b_{i}=0$.
This completes the induction and establishes the correctness of our claim. ∎
Proof of Theorem 16.
Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 13. We will show that the output allocation $A$ is always EF1 by induction. For the base cases, the empty bundle is EF1 without an envy cycle. For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $A^{t}$ is EF1 without an envy cycle. Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is EF1 without an envy cycle. Let us consider three cases about item $e_{k+1}$.
  • $\lambda=1$ and item $e_{k+1}$ is the last arriving item. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent $j=2-i$. If she picks it, we are done. If agent $j$ picks it, it means that $c_{j}(e_{k+1})=a_{j}$. For agent $i$, she does not envy agent $j$, so her bundle is still EF. For agent $j$, by Lemma 7, we have $c_{j}(A_{j}^{k})-c_{j}(A_{i}^{k})=b_{j}-a_{j}$. Then, we have $c_{j}(A_{j}^{k}\setminus\{f\}\cup\{e_{k+1}\})-c_{j}(A_{i})=b_{j}-a_{j}-b_{j}+a_{j}=0$ where $A_{j}^{k}$ is EF1 and there must exist an item $f$ such that $c_{j}(f)=b_{j}$, implying that allocation $A^{k+1}$ is EF1 without an envy cycle. When there is no envy between two agents, allocating $e_{k+1}$ to any agent will not violate EF1.
  • $\lambda=2$ and item $e_{k+1}$ needs to allocated firstly. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent $j=2-i$. If agent $i$ picks it, we are done. If not, agent $j$ picks it. For agent $i$, she does not envy agent $j$, so her bundle is EF. For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}^{k})-c_{j}(A_{i}^{k})=b_{j}-a_{j}$ and $c_{j}(e_{k+1})=a_{j}$ by the condition of our algorithm. Thus, this allocation is EF1 without an envy cycle. When there is no envy between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
  • $\lambda=2$ and item $e_{k+2}$ needs to be allocated secondly. When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent $j=2-i$. If agent $i$ picks it, we are done. If not, agent $j$ picks it. In that case, item $e_{k}$ is allocated to agent $i$. For agent $i$, she does not envy agent $j$, so her bundle is EF. For agent $j$, we can still derive that her new bundle is EF1 since by our assumption, we have $A^{k-1}_{j}$ is EF1, and $c_{j}(e_{k+2})\leq c_{j}(e_{k+1})$ by the condition of our algorithm. Thus, this allocation is EF1 without an envy cycle. When there is no envy between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
Combining the above cases together, we can conclude that the allocation $A^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
B.2 Partial Information in Advance
In this part, agents’ preferences over items are the same, and the arriving order of items is known. With this information, we can obtain a non-wasteful EF1 allocation by the Marginal-Greedy (Algorithm 2) in the goods setting and a complete EF1 allocation by the Round-Robin algorithm in the chores setting.
Definition 12.
An online goods/chores instance is monotone, if for any agent $i\in N$, we have $f_{i}(e_{1})\geq\cdots\geq f_{i}(e_{T})$ or for any agent $i\in N$, we have $f_{i}(e_{1})\leq\cdots\leq f_{i}(e_{T})$, where $f_{i}\in\{v_{i},c_{i}\}$
Theorem 17.
For the deterministic allocation of indivisible goods from a monotone instance, the Marginal-Greedy Algorithm (Algorithm 2) computes an NW allocation that is EF1.
Proof.
Without loss of generality, we assume that the picking sequence is $1,\ldots,n$. First, let us consider the case of $v_{i}(e_{1})\leq\ldots\leq v_{i}(e_{T})$ for any $i\in N$. Note that, for any agent $i\in N$, if she values some items at zero, these items will arrive before the items with non-zero values. In this phase, she does not pick anything until the item with non-zero values arrives. Additionally, when these zero items are allocated to other agents, from the perspective of agent $i$, the value of the other agent’s bundle remains unchanged. Thus, without loss of generality, we assume that $v_{i}(e_{1})\neq 0$ for any agent $i\in N$. Fix an arbitrary round $k$ and an arbitrary agent $i\in N$. Let $A^{k}$ denote the allocation after the allocation of $e_{k}$. Suppose that she picks item $e_{k}$. For agent $i$, we have $v_{i}(A^{k}_{i})\geq v_{i}(A^{k}_{j}\setminus\{e\})$, where $e$ is the last item added to agent $j$’s bundle, since for any $j<i$, agent $i$ is the latter one to pick the item and for any $j>i$, $v_{i}(e_{s})\geq v_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and $s-t=n-i+j$. For agent $j\neq i$, we have $v_{j}(A_{j}^{k})\geq v_{j}(A_{i}^{k}\setminus\{e\})$, where $e$ is the last item added to agent $i$’s bundle, since for any $j>i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and $s-t=j-i$, and $j<i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and $s-t=n-j+i$. Thus, the allocation $A^{k}$ is EF1. As for the case of $v_{i}(e_{1})\geq\ldots\geq v_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it. ∎
Theorem 18.
For the deterministic allocation of indivisible chores from a monotone instance, the Round-Robin Algorithm computes a complete allocation that is EF1.
Proof.
Without loss of generality, we assume that the picking sequence is $1,\ldots,n$. First, let us consider the case of $c_{i}(e_{1})\leq\ldots\leq c_{i}(e_{T})$ for any $i\in N$. Fix an arbitrary round $k$ and an arbitrary agent $i\in N$. Let $A^{k}$ denote the allocation after the allocation of $e_{k}$. Suppose that she picks item $e_{k}$. For agent $i$, we have $c_{i}(A^{k}_{i}\setminus\{e_{k}\})=c_{i}(A^{k-1}_{i})\leq c_{i}(A^{k-1}_{j})=c_{i}(A^{k}_{j})$ since for any $j>i$, agent $i$ has the priority of picking the item and for any $j<i$, $c_{i}(e_{s})\leq c_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and $t-s=n-i+j$. For agent $j\neq i$, we have $c_{j}(A_{j}^{k})\leq c_{j}(A_{i}^{k})$ for any $j<i$ agent $j$ has the priority of picking items and $j>i$, $c_{j}(e_{s})\leq c_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and $t-s=n-j+i$. Therefore, the allocation $A^{k}$ is EF1. As for the case of $c_{i}(e_{1})\geq\ldots\geq c_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it.∎
In the monotone instance, each agent has the same preference over all items, and the arriving order of items follows the nonincreasing/nondecreasing order in terms of value or cost. Then, if we allow the arriving order to be arbitrary and keep the same preference, we have the following proposition.
Proposition 7.
For an online goods/chores instance where agents only have identical preferences, there exists an instance for which no EF1 allocation can be found, even for two agents with bi-valued valuation or cost functions.
Proof.
$t=1$ $t=2$ $t=3$ $t=4$
agent $1$ $1$ $1$ $\epsilon$ $1$
agent 2 $1$ $\epsilon$ $\epsilon$ $1$
Case 1
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $1$ $1$ $\epsilon$ $1$ $1$
agent 2 $1$ $\epsilon$ 1 $\epsilon$ $1$
Case 2
Table 12: The nonexistence of EF1 for two agents that have the same preference over goods.
$t=1$ $t=2$ $t=3$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 1
$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
agent $1$ $1$ $1$ $\frac{1}{\epsilon}$ 1 $\frac{1}{\epsilon}$
agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
Case 2
Table 13: The nonexistence of EF1 for two agents that have the same preference over chores.
First, we consider the online goods instance. Without loss of generality, we assume that $v_{1},v_{2}\in\{\epsilon,1\}$, where $10<\epsilon<$ 1 The value of each item is shown in Table 12. When $t=1$, the arriving item has the value of $1$ for both agents. Without loss of generality, it is allocated to agent 1. For the second round $t=2$, the item that has the value of $1$ for agent 1 and $\epsilon$ for agent 2. To guarantee EF1, agent 2 picks it. In round $t=3$, the arriving item has the value of $\epsilon$ for both agents. Then, let us consider two cases related to allocating it.
  • If agent 2 picks it, this allocation is still EF1. In the last round $t=4$, the arriving item has the value of $1$ for both agents. In this case, it is not hard to check that no matter which agent picks it, the resulting allocation is not EF1.
  • If agent 1 picks it, this allocation is still EF1. Then, in round $t=4$, the arriving item has the value of $1$ for agent 1 and $\epsilon$ for agent 2. To guarantee that EF1 is not violated, it should be allocated to agent 2. In this last round $t=5$, the item has the value of $1$ for both agents. In this case, it is trivial that as long as one agent picks it, the final allocation is not EF1.
Next, let us consider the online chores instance. We assume that $c_{1},c_{2}\in\{1,\frac{1}{\epsilon}\}$ where $10<\epsilon<$ 1. The cost of each item is shown in Table 13. When $t=1$, the arriving item has the cost of 1 for both agents, and we assume that agent 1 picks it. In round $t=2$, the item that has the cost of 1 for agent $1$ and $\frac{1}{\epsilon}$ for agent 2. To maintain EF1, agent 2 picks it. For the next round $t=3$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. Then, let us consider two cases related to allocating it.
  • If it is allocated to agent 2, no further items appear. In this case, it is clear that this allocation is not EF1.
  • If it is allocated to agent 1, in the following round $t=4$, the arriving item has the cost of 1 for agent $1$ and $\frac{1}{\epsilon}$ for agent 2. It should be allocated to agent 2 to ensure that the property of EF1 still holds. In the last round $t=5$, this arriving item has the cost of $\frac{1}{\epsilon}$ for both agents. In this case, it is easy to verify that no allocation is EF1 when it is allocated to one agent.
Combining the above cases, we can conclude that our statement holds. ∎
Input: An instance $(T,N,(v_{1},v_{2}))$ with bi-valued valuation functions and deadline $d=1$
Output: An non-wasteful EF1 allocation $A$
1
2Let $A=(\emptyset,\ldots,\emptyset)$;
3Initialize $\lambda=1$;
4 when item $e\in T$ arrives do
5
6      if $\lambda=1$  then
7
8            Let this item wait for $d$ period and use $e_{1}$ to denote it.
9            $\lambda=2$;
10      else
11
12            Let $e_{2}$ denote the new arriving item;
13            if  there exists an envy between the two agents  then
14
15                  Let $i\in N$ be the agent who envies the other one and $j=2-i$;
16                  if  $v_{i}(e_{1})=b_{i}$  then
17                         $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
18                   else if $v_{i}(e_{2})=b_{i}$ then
19                         $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
20                   else if $v_{j}(e_{1})=b_{j}$ then
21                         $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
22                   else
23                         $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
24                   end if
25
26            else
27                   if  $v_{i}(e_{1})>v_{i}(e_{2})$  then
28                         $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
29                   else if  $v_{i}(e_{1})<v_{i}(e_{2})$  then
30                         $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
31                   else
32                         if  $v_{j}(e_{1})>v_{j}(e_{2})$  then
33                               $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
34                        else
35                               $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
36                         end if
37
38                   end if
39
40             end if
41
42            $\lambda=1$;
43       end if
44
45
46
47if  $\lambda=1$ and no further items appear  then
48       if there is no envy between two agents then
49
50            Allocate item $e_{1}$ to an arbitrarily agent;
51      else
52
53            Let $i\in N$ be the agent who envies the other one and $j=2-i$;
54            if $v_{i}(e_{1})=b_{i}$ then
55                   $A_{i}=A_{i}\cup\{e_{1}\}$;
56            else
57                   $A_{j}=A_{j}\cup\{e_{1}\}$;
58             end if
59
60       end if
61
62
return $A$;
Algorithm 12 Adapted Matching Algorithm
Input: An instance $(T,N,(c_{1},c_{2}))$ with bi-valued cost functions and deadline $d=1$
Output: A complete EF1 allocation $A$
1
2Let $A=(\emptyset,\ldots,\emptyset)$;
3Initialize $\lambda=1$;
4 when item $e\in T$ arrives do
5
6      if $\lambda=1$  then
7
8            Let this item wait for $d$ period and use $e_{1}$ to denote it.
9            $\lambda=2$;
10      else
11
12            Let $e_{2}$ denote the new arriving item;
13            if  there exists an envy between the two agents  then
14
15                  Let $i\in N$ be the agent who does not envy the other one, and $j=2-i$;
16                  if  $c_{j}(e_{1})=b_{j}$  then
17                         $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
18                   else if $c_{j}(e_{2})=b_{j}$ then
19                         $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
20                   else if $c_{i}(e_{1})=b_{i}$ then
21                         $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
22                   else
23                         $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
24                   end if
25
26            else
27                   if  $c_{j}(e_{1})>c_{j}(e_{2})$  then
28                         $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
29                   else if  $c_{j}(e_{1})<c_{j}(e_{2})$  then
30                         $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
31                   else
32                         if  $c_{i}(e_{1})>c_{i}(e_{2})$  then
33                               $A_{j}=A_{j}\cup\{e_{1}\}$ and $A_{i}=A_{i}\cup\{e_{2}\}$;
34                        else
35                               $A_{i}=A_{i}\cup\{e_{1}\}$ and $A_{j}=A_{j}\cup\{e_{2}\}$;
36                         end if
37
38                   end if
39
40             end if
41
42            $\lambda=1$;
43       end if
44
45
46
47if  $\lambda=1$ and no further items appear  then
48       if there is no envy between two agents then
49
50            Allocate item $e_{1}$ to an arbitrarily agent;
51       end if
52
53      Let $i\in N$ be the agent who does not envy the other one and $j=2-i$;
54      if $c_{j}(e_{1})=b_{j}$ then
55             $A_{i}=A_{i}\cup\{e_{1}\}$;
56      else
57             $A_{j}=A_{j}\cup\{e_{1}\}$;
58       end if
59
60
return $A$;
Algorithm 13 Adapted Chores Matching Algorithm

### v2

@@
-Online Fair Allocations with Binary Valuations and Beyond
+Yuanyuan Wang1    Tianze Wei2
 Abstract
@@
 In our paper, we study the online allocation of either goods or chores.
-We adapt popular fairness notions, including envy-freeness up to one item (EF1) and maximin share fairness (MMS), and use utilitarian social welfare (USW) to measure efficiency.
-For both settings of items, we present a series of positive results regarding the existence of fair and efficient allocations with widely studied classes of binary and bivalued valuation/cost functions.
-Additionally, we complement our results by constructing some counterexamples to establish our results as among the best guarantees possible.
+We employ popular fairness notions, including envy-freeness up to one item (EF1) and maximin share fairness (MMS) to capture fairness, and utilitarian social welfare (USW) to measure efficiency.
+For both settings of items, we present a series of positive results regarding the existence of fair and efficient allocations with widely studied classes of additive binary and personalized bi-valued valuation/cost functions.
+Furthermore, we complement our results by constructing counterexamples to establish our results as among the best guarantees possible.
 1 Introduction
-How to allocate a set of indivisible items fairly among a set of agents is a hot topic in various research communities, including algorithmic game theory and artificial intelligence.
-For decades, there have been tremendous ongoing efforts to design efficient algorithms for finding fair allocations regarding envy-free up to one item (EF1), maximin share fairness (MMS), and some other fairness; detailed progress and results can be found in [1].
-Most prior work focuses on offline setting, where the set of items is known upfront.
-In many practical situations, decisions about the allocation of items must be made online without or with limited knowledge about the future.
-A paradigmatic example is the food banks problem [2].
-In this scenario, food comes online without prior knowledge of future food, and the food is perishable; it must be allocated upon arrival to one of the agents.
-In this paper, we study the online problem of allocating a set of items that arrive one by one to a set of agents that are known in advance.
-When an item arrives, its values to the agents are revealed, and an immediate and irrevocable decision must be made about which agent to assign the item.
-There has been a line of work about online fair allocation, where they consider either additive binary valuations [3] or normalized valuations [4, 5, 6].
-Few non-trivial results about the approximation fairness guarantee of online goods allocation beyond additive binary valuations and online chores allocation are known when there is no prior information about the items.
-In our paper, we would like to fill this gap and consider the online allocation model without any information about future items, where we discuss the goods and chores settings, respectively.
-Our primary research question is
-Can we design online algorithms that are fair and efficient, for general valuation functions beyond additive binary, in the online allocation problem?
+The fair division of indivisible items is a prominent topic in algorithmic game theory and artificial intelligence, with practical applications Budish (2011); Goldman and Procaccia (2015).
+Most prior work has focused on offline settings, where the information of all items is known a priori.
+In practice, items are often online, and waiting for all resources to appear can significantly impede progress.
+We must proactively allocate resources in a timely manner for better outcomes.
+A paradigmatic example is the food banks problem Lee et al.
+(2019), where food arrives online without prior knowledge of future items.
+Since the food is perishable, it must be allocated upon arrival to some agent.
+Motivated by these real-world scenarios, we define and discuss the following problem:
+(Online Fair Division) There are $n$ agents that are fixed and known in advance.
+The items arrive online (the number of items is unknown).
+Upon the arrival of an item, its values to all agents are revealed, and we must allocate it to some agent or discard it immediately and irrevocably.
+At all times, the allocation among agents must be fair.
+If the online algorithm knows additional information about all future items beforehand, it somewhat reduces the uncertainty of the online model, allowing for more effective decision-making and improved outcomes.
+Zhou et al.
+(2023); Benadè et al.
+(2024) study the normalized valuations, i.e., online algorithms know the total value of future items.
+Benadè et al.
+(2022) studies the online fair allocation, foreseeing the rank of a new arrival item with respect to the past items under different assumptions of agents’ valuation distributions.
+And there are other studies about additional information He et al.
+(2019); Elkind et al.
+(2024); Cookson et al.
+(2024).
+For an independent interest, we also examine the online setting with additional information, with relevant results presented in the Appendix.
+In numerous practical applications, such as vaccine allocation Rey et al.
+(2023), the online algorithm operates without prior knowledge of future items.
+Aleksandrov et al.
+(2015) studies the online fair division with additive binary valuations and shows that an envy-free up to one item (EF1) allocation exists.
+However, there are a few non-trivial approximation fairness results of online goods allocation beyond additive binary valuations and online chores allocation.
+In our paper, we aim to address this gap and study the online allocation model without any information about future items.
+We discuss the goods and chores settings, respectively.
+Our primary research question is:
+Can we design deterministic online algorithms that are fair and efficient, for valuation functions beyond additive binary, in the online allocation problem?
 1.1 Our Contributions
-Valuations $N$ Fairness/ Efficiency LB UB
+We study two settings: goods and chores, in the online fair indivisible items allocation model.
+In the goods setting, for different classes of valuations, including binary and additive personalized bi-valued, our main results are summarized in Table 1.
+We introduce the non-wastefulness (NW) notion in this setting.
+First, we show the inapproximability of EF1 and MMS with general additive valuations, even for only two agents with additive personal tri-valued valuations.
+Then, we study some restricted valuation functions, such as submodular binary valuation functions.
+Valuations
+Number
+of agents
+Fairness/
+Efficiency
+LB UB
+Additive
+(personalized tri-valued)
+[…]
+[…]
+Submodular
+binary
+$n$ EF1
+$\frac{1}{2}$
+(Thm 2)
+$\frac{1}{2}$
+(Thm 3)
+[…]
+MMS
+$\frac{1}{2}$
+(Thm 2)
+$\frac{1}{2}$
+(Thm 3)
+[…]
+USW
+$\frac{1}{2}$
+(Thm 2)
+$\frac{1}{2}$
+(Thm 3)
+[…]
+Additive personalized
+bi-valued
+22 EF1
+$\frac{1}{2}$
+(Thm 4)
+$\frac{1}{2}$
+(Thm 5)
+[…]
+MMS
+$\frac{1}{3}$
+(Thm 4)
+$\frac{1}{3}$
+(Thm 5)
+[…]
+[…]
+Identical additive
+binary
++ personalized bi-valued
+[…]
+[…]
+USW $\times$ (Prop 3)
+Table 1: Goods setting. LB: lower bound. UB: upper bound. $\times"$ indicates inapproximability.
+  • For submodular binary valuation functions, we propose the Marginal-Greedy Algorithm (Algorithm 2) to compute an allocation that simultaneously satisfies NW, $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW, and we show that these approximation ratios are all tight.
+  • For two agents with additive personalized bi-valued valuation functions, we design the Adapted Envy-Graph Procedure (Algorithm 3) to compute an NW allocation satisfying $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS simultaneously.
+Further, we show that no deterministic algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\beta$-MMS for any $\alpha>\frac{1}{2}$, $\beta>\frac{1}{3}$, and no deterministic algorithm can compute an NW allocation where $\alpha$-EF1 (or $\alpha$-MMS) is compatible with $\beta$-max-USW for any $\alpha,\beta>0$.
+  • When one agent has the additive personalized bi-valued valuation function and the remaining agents have identical additive binary valuation functions, we design the Adapted-Picking Algorithm (Algorithm 7) to compute an allocation satisfying NW, EF1, and MMS.
+Further, we show that no deterministic algorithm can compute an NW allocation that is EF1 (or MMS) and $\alpha$-max-USW for any $\alpha>0$.
+Costs
+Number
+of agents
+Fairness/
+Efficiency
+LB UB
+Additive
+(personalized tri-valued)
+[…]
+Additive
+binary
 […]
@@
 […]
+Supermodular
+binary
 […]
@@
 […]
+Additive personalized
+bi-valued
+22 EF1
+22
+(Thm 11)
+22
+(Thm 10)
+[…]
+MMS
+$\frac{3}{2}$
+(Thm 11)
+$\frac{5}{3}$
+(Thm 10)
+[…]
+[…]
+Additive binary
++ personalized bi-valued
 […]
@@
 […]
-[…]
-USW $\times$ (Prop 2)
-(a) Goods setting
-Costs $N$ Fairness/ Efficiency LB UB
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-(b) Chores setting
-Table 1: A summary of our main results. "$\times$" indicates inapproximability, $N$ indicates the number of agents. "LB" indicates lower bound, and "UB" indicates upper bound.
-There are $n$ agents, and each indivisible item comes online.
-We study the goods and chores settings, and introduce non-wastefulness (NW) and completeness constraints to the goods and chores settings, respectively.
-For different classes of valuation or cost functions, including submodular/supermodular binary, additive bi-valued, and so on, our main results are summarized in Table 1.
-In the goods setting, we first show the inapproximability of EF1 and MMS with general additive valuations, even for only two agents with additive tri-valued valuations.
-This indicates that we can only focus our efforts on restricted valuation functions.
-  • For submodular binary valuation functions, we propose the Marginal-Greedy Algorithm (Algorithm 2) to compute allocations that simultaneously satisfy NW, $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW.
-Further, we show that no deterministic algorithm can guarantee an NW allocation that is $\alpha$-EF1, $\alpha$-MMS, or $\alpha$-max-USW for any $\alpha>\frac{1}{2}$.
-  • For two agents with additive bi-valued valuation functions, we design the Adapted Envy-Graph Procedures (Algorithm 3) to compute NW, $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS allocations.
-Further, we show that no deterministic algorithm can guarantee a NW allocation that is $\alpha$-EF1 for any $\alpha>\frac{1}{2}$ or $\beta$-MMS for any $\beta>\frac{1}{3}$, and no deterministic algorithm can achieve NW, $\alpha$-EF1 or $\alpha$-MMS, and $\gamma$-max-USW for any $\alpha,\gamma>0$.
-  • When one agent has the additive bi-valued valuation function and the remaining agents have identical additive binary valuation functions, we design the Adapted-Picking Algorithm (Algorithm 6) to compute an allocation that simultaneously satisfies NW, EF1, and MMS.
-Further, we show that no deterministic algorithm can compute an NW allocation that is EF1 or MMS guarantee $\alpha$-max-USW for any $\alpha>0$.
-In the chores setting, we first demonstrate the inapproximability of EF1 for additive tri-valued cost functions, which hinders our ability to explore more general cost functions.
+Table 2: Chores setting. LB: lower bound. UB: upper bound. $\times"$ indicates inapproximability.
+In the chores setting, we introduce the completeness constraints.
+We first show the inapproximability of EF1 even for two agents with additive personalized tri-valued cost functions, which blocks our way of studying more general cost functions.
 Next, we investigate the restricted cost functions, such as additive binary cost functions.
   • For additive binary cost functions, we propose the Compelled-Greedy Algorithm (Algorithm 8) to compute a complete allocation that simultaneously satisfies EF1, MMS, and min-USC.
-  • For supermodular binary cost functions, we prove that no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1, $\alpha$-MMS, or $\alpha$-min-USC for any $\alpha\geq 1$.
-  • When two agents have additive bi-valued cost functions, we design the Adapted Chores Envy-Graph Procedures (Algorithm 8) to compute an allocation that satisfies completeness, $2$-EF1, and $\frac{5}{3}$-MMS.
-Further, we show that no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any $\alpha<2$ or $\beta$-MMS for any $\beta<\frac{3}{2}$, and no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 or $\beta$-MMS, and $\gamma$-min-USC for any $\alpha,\gamma\geq 1$ and $\beta<2$.
-  • When one agent has the additive bi-valued cost function and the remaining agents have additive binary cost functions, we design the Adapted-Chores-Picking Algorithm (Algorithm 11) to compute an allocation that simultaneously satisfies completeness, EF1 and MMS.
-Further, no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1 or MMS guarantee $\beta$-min-USC for any $\alpha,\beta\geq 1$.
+For supermodular binary cost functions, we prove that no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1 (or $\alpha$-MMS or $\alpha$-min-USC) for any $\alpha\geq 1$.
+  • When two agents have additive personalized bi-valued cost functions, we design the Adapted Chores Envy-Graph Procedure (Algorithm 9) to compute an allocation that satisfies completeness, $2$-EF1, and $\frac{5}{3}$-MMS.
+Further, we show that no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any $\alpha<2$ or $\beta$-MMS for any $\beta<\frac{3}{2}$, and no deterministic algorithm can guarantee a complete allocation that is $\alpha$-EF1 (or $\beta$-MMS) and $\gamma$-min-USC for any $\alpha,\gamma\geq 1$ and $\beta<2$.
+  • When one agent has the additive personalized bi-valued cost function and the remaining agents have additive binary cost functions, we design the Adapted-Chores-Picking Algorithm (Algorithm 13) to compute an allocation satisfying completeness, EF1, and MMS.
+Further, no deterministic algorithm can compute a complete allocation that is $\alpha$-EF1 (or MMS) and $\beta$-min-USC for any $\alpha,\beta\geq 1$.
 1.2 Related Work
 Online Fair Allocation.
-In the majority of the literature, there are two kinds of online arrivals: The former kind is that agents arrive over time [7, 8, 9, 10], and the latter is that items arrive over time, like [3], which is also the topic of our paper.
-Our work is mainly related to the study of [3] and [11], where items arrive one by one without any future information known in advance.
+In the majority of the literature, there are two kinds of online arrivals: The former kind is that agents arrive over time Kash et al.
+(2014); Friedman et al.
+(2015, 2017); Li et al.
+(2018), and the latter is that items arrive over time, like Aleksandrov et al.
+(2015), which is also the topic of our paper.
+Our work is mainly related to the study of Aleksandrov et al.
+(2015) and Hosseini et al.
+(2024), where items arrive one by one without any future information known in advance.
 Aleksandrov et al.
-[3] consider the food bank problem when all agents only have the additive binary valuation functions and design two simple mechanisms to guarantee ex-ante envy-freeness.
+(2015) consider the food bank problem when all agents only have the additive binary valuation functions and design two simple mechanisms to guarantee ex-ante envy-freeness.
 In our paper, we study beyond additive binary valuation functions in the goods setting.
 Hosseini et al.
-[11] study the class fairness in online matching for indivisible and divisible items, and present a series of positive results.
+(2024) study the class fairness in online matching for indivisible and divisible items, and present a series of positive results.
 The model of indivisible items in their paper can be seen as a special case of our model with submodular binary valuation functions in our paper, which will be discussed later.
 In other words, we study a more general setting and derive the approximation results that strengthen their results.
-Some work studies about online fair allocation with some prior information of future items, like the normalized valuations [5, 6], the rank of a new arrival item [12], and other information [13, 14, 15].
-Zhou et al.
-[5] study the online allocation of indivisible goods or chores with normalized valuations and design several algorithms only to guarantee the final allocation that approximates maximin share, which is in contrast to our setting, where our online algorithm ensures the approximation ratio in any round.
-Benade et al.
-[6] study the trade-off between fairness and efficiency in allocating the indivisible items under different types of adversaries with normalized valuations.
-Benade et al.
-[12] study the allocation of indivisible items with partial information, i.e., the rank of a new arrival item with respect to the past items under different assumptions of agents’ valuation distributions.
-He et al.
-[13] study the minimum adjustments (the reallocation of previously allocated items) to ensure envy-freeness up to one item for online indivisible items allocation with or without the information where the values of future items are observed.
-Elkind et al.
-[14] and Cookson et al.[15] study the temporal allocation of online indivisible items where the arriving information of items, like values of items, is known and give several positive results.
-With some known information, we also explore online fair allocation, and the relevant results are presented in Appendix B.
 Offline Fair Allocation.
 When agents have additive valuation functions, there is a rich body of literature that discusses EF1 or MMS.
-Please see the survey of [1] for an overview.
-Here, we list some papers most relevant to our work about offline fair division with submodular binary valuation functions or supermodular binary cost functions.
+Please refer to the survey by Amanatidis et al.
+(2023) for an overview.
+Here, we highlight some papers most relevant to our work on offline fair division with submodular binary valuation functions or supermodular binary cost functions.
 When agents have submodular binary valuations in the goods setting, Benabbou et al.
-[16] study the compatibility of fairness and efficiency and show that a maximizing utilitarian social welfare allocation that is EF1 can be efficiently computed, and Barman et al.
-[17] consider the fairness notion - MMS and prove that a maximizing utilitarian social welfare allocation that satisfies MMS can be computed in polynomial time.
+(2021) study the compatibility of fairness and efficiency and show that a maximum utilitarian social welfare allocation that is EF1 can be efficiently computed, and Barman and Verma (2021) consider the fairness notion - MMS and prove that a maximum utilitarian social welfare allocation that satisfies MMS can be computed in polynomial time.
 When agents have supermodular binary cost functions in the chores setting, Barman et al.
-[18] show that an EF1 or MMS allocation that minimizes the social cost can be computed in polynomial time.
+(2023) show that an EF1 or MMS allocation minimizing the social cost can be computed in polynomial time.
 2 Preliminaries
-For each natural number $s\in\mathbb{N}$, let $[s]=\{1,\ldots,s\}$.
+For each natural number $s\in\mathbb{N}$, let $s]=\{1,\ldots,s\}$.
 Let $T=\{e_{1},\ldots,e_{t}\}$ be the set of indivisible items arriving one by one and $N=\{1,\ldots,n\}$ be the set of offline agents.
-When an online item $e_{t}$ arrives, its incident values or costs to agents are revealed.
+When an online item $e_{t}$ arrives, its values or costs to all agents are revealed.
 When items are goods, each agent is endowed with a valuation function $v_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$.
 When items are chores, each agent is endowed with a cost function $c_{i}:2^{T}\rightarrow\mathbb{R}_{\geq 0}$.
-We assume that $v_{i}(\emptyset)=0$ ($c_{i}(\emptyset)=0$), and monotone, $v_{i}(X)\leq v_{i}(Y)$ ($c_{i}(X)\leq c_{i}(Y)$) for all $X\subseteq Y\subseteq M$.
-For any item $e\in T$ and any subset $X\subseteq T$, we denote by $\Delta^{i}_{X}(e)=v_{i}(X\cup\{e\})-v_{i}(X)$ or $c_{i}(X\cup\{e\})-c_{i}(X)$ the marginal value or cost of an item $e\in M$ in the set $X$.
-$v_{i}$ is additive if the values of any set of items $S$ is the sum of the value of each item in $S$, that is, $v_{i}(X)=\sum_{e\in X}v_{i}(e)$.
+We assume that $v_{i}(\emptyset)=0$ ($c_{i}(\emptyset)=0$), and monotone, $v_{i}(X)\leq v_{i}(Y)$ ($c_{i}(X)\leq c_{i}(Y)$) for all $X\subseteq Y\subseteq T$.
+For any item $e\in T$ and any subset $X\subseteq T$, we denote by $\Delta^{i}_{X}(e)=v_{i}(X\cup\{e\})-v_{i}(X)$ or $c_{i}(X\cup\{e\})-c_{i}(X)$ the marginal value or cost of an item $e\in T$ in the set $X$.
+$v_{i}$ is additive if the value of any set of items $S$ is the sum of the value of each item in $S$, that is, $v_{i}(X)=\sum_{e\in X}v_{i}(e)$.
 $v_{i}$ is submodular if $v_{i}(X)+v_{i}(Y)\geq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$.
 $v_{i}$ is supermodular if $v_{i}(X)+v_{i}(Y)\leq v_{i}(X\cup Y)+v_{i}(X\cap Y)$ for any two sets $X,Y\subseteq T$.
-There are similar arguments for $c_{i}$, and we omit them.
+There are similar arguments for $c_{i}$, and we omit it.
 Let $T^{k}$ be the set of items that have arrived by round $k\in[t]$.
-We denote an allocation by $A^{k}=(A_{1}^{k},\ldots,A_{n}^{k})$, where $A_{i}^{k}\subseteq T^{k}$ is the bundle of items that are allocated to agent $i$, and $A_{i}^{k}\cap A_{j}^{k}=\emptyset$ for any $i,j\in N$.
+We denote an allocation by ${\bf A}^{k}=(A_{1}^{k},\ldots,A_{n}^{k})$, where $A_{i}^{k}\subseteq T^{k}$ is the bundle of items that are allocated to agent $i$, and $A_{i}^{k}\cap A_{j}^{k}=\emptyset$ for any $i,j\in N$.
+Let $\Pi_{n}(T^{k})$ be the set of all allocations of $T^{k}$.
 Collectively, $(T,N,(v_{i})_{i\in N})$ forms an instance of online goods fair allocation and $(T,N,(c_{i})_{i\in N})$ forms an instance of online chores fair allocation.
 For a singleton set $\{e\}$, we will use $v_{i}(e)$ ($c_{i}(e)$) as a shorthand for $v_{i}(\{e\})$ ($c_{i}(\{e\})$).
-Deterministic online algorithms
-When one item arrives, the values (costs) of this item for all agents are revealed.
-The online algorithm must make an immediate and irrevocable decision to allocate the item to one agent or throw it away deterministically.
-Our goal is to design the deterministic algorithms to compute allocations guaranteeing the desired approximation fairness or efficiency (approximate EF1, MMS, USW, USC) (Definitions 1, 2, 3, 8, 9, 10) and constraint (non-wastefulness for goods (Definition 4) and completeness for chores (Definition 7)) at the end of each round.
+Deterministic online algorithms.
+In the online setting, the items in $T$ arrive one by one in an arbitrary order, where $|T|$ is unknown.
+When an item arrives, the values (costs) of this item for all agents are revealed.
+The online algorithm must make an immediate and irrevocable decision to allocate the item to one agent or discard it deterministically.
+Our objective is to design deterministic algorithms to compute allocations guaranteeing the desired approximation fairness or efficiency, including EF1, MMS, and USW/USC (Definitions 1, 2, 3, 8, 9, and 10), as well as constraints, including non-wastefulness (Definition 4) and completeness (Definition 7), at the end of each round.
 In contrast, our impossibility results will hold even if the desired guarantees are required to hold only at the end.
-Let us demonstrate the deterministic online allocation procedure with the following example.
-$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
-agent $1$ $6$ $4$ $5$ 9 3
-agent 2 $5$ $8$ $10$ 2 5
-agent 3 $12$ 6 6 8 4
+For a better understanding of our online model, let us demonstrate the deterministic online allocation procedure through the following example.
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
+agent 11 6 88 10 66
+agent 2 13 8 5 0
 Table 3: Example of the deterministic online allocation.
 Example.
-Let us consider the following example to demonstrate the deterministic online allocation procedure.
-Suppose that there are three agents who have additive valuation functions, and the arriving items are goods.
-The value of each item is shown in Table 2.
+Suppose that there are two agents who have additive valuation functions, and the arriving items are goods.
+The value of each item is shown in Table 3, and note that the value of each item is only revealed after it arrives.
 Now we have a deterministic online algorithm named $\mathsf{OALG}$.
-In round $t=1$, $\mathsf{OALG}$ allocates $e_{1}$ to agent 1 after the values of this item are revealed.
-For allocation $A^{1}$, it is EF1, MMS and $\frac{1}{2}$-max-USW.
-In round $t=2$, by $\mathsf{OALG}$, $e_{2}$ is allocated to agent 2, and allocation $A^{2}$ is EF1, MMS, and $\frac{7}{10}$-max-USW.
-In round $t=3$, by $\mathsf{OALG}$, agent 3 picks item $e_{3}$ and allocation $A^{3}$ is EF1, MMS, and $\frac{2}{3}$-max-USW.
-In round $t=4$, by $\mathsf{OALG}$, agent 1 picks item $e_{4}$ and allocation $A^{4}$ is $\frac{3}{4}$-EF1, $\frac{3}{4}$-MMS, and $\frac{29}{39}$-max-USW.
-In round $t=5$, by $\mathsf{OALG}$, agent 1 picks item $e_{5}$ and allocation $A^{5}$ is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{8}{11}$-max-USW.
-Then, no further item will arrive and we say that $\mathsf{OALG}$ achieves $\min\{\frac{1}{2},\frac{3}{4},1\}=\frac{1}{2}$-EF1, $\min\{\frac{1}{2},\frac{3}{4},1\}=\frac{1}{2}$-MMS, and $\min\{\frac{1}{2},\frac{2}{3},\frac{7}{10},\frac{8}{11},\frac{29}{39}\}=\frac{1}{2}$-max-USW.
+In round $t=1$, $\mathsf{OALG}$ allocates $e_{1}$ to agent 1, and the allocation ${\bf A}^{1}$ is EF1, MMS, and $\frac{6}{13}$-max-USW.
+In round $t=2$, $\mathsf{OALG}$ allocates $e_{2}$ to agent 2, and the allocation ${\bf A}^{2}$ is EF1, MMS, and $\frac{2}{3}$-max-USW.
+In round $t=3$, $\mathsf{OALG}$ allocate $e_{3}$ to agent 2, and the allocation $\textbf{A}^{3}$ is $\frac{3}{4}$-EF1, $\frac{3}{5}$-MMS, and $\frac{19}{31}$-max-USW.
+In round $t=4$, $\mathsf{OALG}$ discards $e_{4}$, and the allocation $\textbf{A}^{4}$ is $\frac{3}{4}$-EF1, $\frac{3}{7}$-MMS, and $\frac{19}{37}$-max-USW.
+Then, no further item arrives and we say that $\mathsf{OALG}$ achieves $\min\{\frac{3}{4},1\}=\frac{3}{4}$-EF1, $\min\{\frac{3}{7},\frac{3}{5},1\}=\frac{3}{7}$-MMS, and $\min\{\frac{6}{13},\frac{19}{37},\frac{19}{31},\frac{2}{3}\}$$\frac{6}{13}-$max-USW.
 3 Allocation of Goods
-In this section, we focus on the setting of goods. Here are some definitions of fairness and efficiency notions.
+Throughout this section, we focus on the goods setting. We first introduce some definitions of fairness and efficiency notions that are used in this section.
 Definition 1 (Envy-freeness up to One Good).
-For any constant $\alpha\in[0,1]$, an allocation $A^{k}$ is $\alpha$-approximate envy-free up to one good ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i^{\prime}}^{k}\neq\emptyset$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot v_{i}(A_{i^{\prime}}^{k}\setminus\{e\})$ for some $e\in A_{i^{\prime}}^{k}$.
+For any $\alpha\in[0,1]$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate envy-free up to one good ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i^{\prime}}^{k}\neq\emptyset$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot v_{i}(A_{i^{\prime}}^{k}\setminus\{e\})$ for some $e\in A_{i^{\prime}}^{k}$.
 Definition 2 (Maximin Share Fairness).
@@
 $\mathsf{MMS}_{i}^{k}=\max_{B^{k}\in\Pi_{n}(T^{k})}\min_{i^{\prime}\in N}v_{i}(B_{i^{\prime}}^{k}),$
-where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
-For any constant $\alpha\in[0,1]$, an allocation $A^{k}$ is $\alpha$-approximate maximin share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
+For any $\alpha\in[0,1]$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate maximin share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $v_{i}(A_{i}^{k})\geq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
 Definition 3 (Utilitarian Social Welfare).
-The utilitarian social welfare $USW$ of allocation $A^{k}$ is given by $USW(A^{k})=\sum_{i\in N}v_{i}(A_{i}^{k})$.
-For any constant $\alpha\in[0,1]$, an allocation $A^{k}$ is $\alpha$-max-USW if $USW(A^{k})\geq\alpha\cdot\max_{B^{k}\in\Pi_{n}(T^{k})}USW(B^{k})$ holds, where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
-In addition to the above definitions, we also pay attention to the following restriction, which is added in the goods setting.
+The utilitarian social welfare of allocation ${\bf A}^{k}$ is given by $USW({\bf A}^{k})=\sum_{i\in N}v_{i}(A_{i}^{k})$.
+For any $\alpha\in[0,1]$, an allocation ${\bf A}^{k}$ is $\alpha$-max-USW if $USW({\bf A}^{k})\geq\alpha\cdot\max_{{\bf B}^{k}\in\Pi_{n}(T^{k})}USW({\bf B}^{k})$.
+Besides the above definitions, we also pay attention to the following restriction, which is added in the goods setting.
 Definition 4 (Non-Wastefulness).
-An allocation $A^{k}$ is non-wasteful (NW) if (1) $\Delta_{A_{i}^{k}\setminus\{e\}}(e)>0$ for any $i\in N$ and any good $e\in A^{k}_{i}$.
-(2) $\Delta_{A_{i}^{k}}({e}^{{}^{\prime}})=0$ for any discarded good $e^{\prime}\in T^{k}$.
-When agents have additive valuation functions, for any arriving good $e$, as long as it is allocated to one agent who values it positively, an NW allocation can be found.
-Next, We explore the existence of submodular valuation functions.
-Claim 1.
+An allocation ${\bf A}^{k}$ is non-wasteful (NW) if (1) $\Delta_{A_{i}^{k}\setminus\{e\}}^{i}(e)>0$ for any agent $i\in N$ and any good $e\in A^{k}_{i}$, and (2) $\Delta_{A_{i}^{k}}^{i}({e}^{{}^{\prime}})=0$ for any agent $i\in N$ and any discarded good $e^{\prime}\in T^{k}$.
+Note that imposing NW in the goods setting is necessary, as our online model permits item discarding (see our example for illustration).
+Without this constraint, achieving EF1 would be trivial by simply discarding all online items.
+For agents with additive valuation functions, an NW allocation can always be found as long as each arriving item is allocated to at least one agent who values it positively.
+Below, we present a positive result where an NW allocation always exists for submodular binary valuation functions that we study in the subsequent parts, and complement it with a negative result for general submodular valuation functions.
+Lemma 1.
 For submodular binary valuation functions, an NW allocation always exists.
 Proof.
-Given an instance $(T,N,(v_{i})_{i\in N})$, Algorithm 1 can compute an NW allocation in polynomial time.
-When item $e_{k}$ ($k\in[t]$) arrives, if there exists an agent $j\in N$ such that $\Delta_{e_{k}}^{j}(A_{j}^{k-1})=1$, $e_{k}$ will be allocated.
-We assume that $e_{k}$ is allocated to agent $i$.
-We will show that $\Delta_{e_{k}}^{i}(A_{i}^{r}\setminus\{e_{r}\})=1$ holds for any round $r>k$ after some new items are added to agent $i$’s bundle.
-Without loss of generality, we assume that $e_{s}$ is an arbitrarily newly added item in agent $i$’s bundle, where $s>k$, it holds that $\Delta_{e_{s}}^{i}(A_{i}^{s-1})=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1})=1$.
-Then, we have
-$\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{r}\})$ $=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
-$=v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
-$\geq v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\setminus\{e_{k}\})-v_{i}(e_{s})$
-$=v_{i}(A_{i}^{s-1})-v_{i}(A^{s-1}\setminus\{e_{k}\})$
-$=\Delta_{e_{k}}^{i}(A_{i}^{s-1}\setminus\{e_{k}\}),$
-where the inequality holds for submodularity.
-Following this induction by repeatedly utilizing submodularity, we have that $\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{k}\})\geq\cdots\geq\Delta_{e_{k}}^{i}(A_{i}^{k}\setminus\{e_{k}\})=\Delta_{e_{k}}^{i}(A_{i}^{k-1})=1$.
-Since $v_{i}$ is a submodular binary valuation function, it holds that $\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{k}\})\leq 1$.
-Thus, we can conclude that $\Delta_{e_{k}}^{i}(A_{i}^{s}\setminus\{e_{k}\})=1$.
-∎
-Claim 2.
-For submodular valuation functions, there exists an instance that no deterministic online algorithm can compute an NW allocation, even for two agents with identical valuation functions.
-Proof.
-We construct an instance with two agents, $v_{1}=v_{2}=v$.
-When first item $e_{1}$ arrives, $v(e_{1})=3$.
-W.l.o.g., we allocate $e_{1}$ to agent 1.
-When the second item $e_{2}$ arrives, $v(e_{2})=3,v(e_{1},e_{2})=5$.
-Case (1): we allocate $e_{2}$ to agent 2, for the third item $e_{3}$,$v(e_{3})=5,v(e_{1},e_{3})=v(e_{2},e_{3})=v(e_{1},e_{2},e_{3})=5$.
-We have that $v(e_{1},e_{3})-v(e_{3})=v(e_{2},e_{3})-v(e_{3})=0$.
-So, no matter who is assigned $e_{3}$, it cannot satisfy NW.
-Case (2): we allocate $e_{2}$ to agent 1, for the third item $e_{3}$ with $v(e_{3})=5,v(e_{1},e_{3})=v(e_{2},e_{3})=v(e_{1},e_{2},e_{3})=5$.
-To ensure NW, we must allocate $e_{3}$ to agent 2.
-For the fourth item $e_{4}$, $v(e_{4})=v(e_{1},e_{4})=v(e_{2},e_{4})=v(e_{3},e_{4})=v(e_{1},e_{2},e_{4})=v(e_{1},e_{3},e_{4})=v(e_{2},e_{3},e_{4})=v(e_{1},e_{2},e_{3},e_{4})=6$.
-We have that $v(e_{1},e_{2},e_{4})-v(e_{2},e_{4})=v(e_{3},e_{4})-v(e_{4})=0$.
-So, no matter who is assigned $e_{4}$, it cannot satisfy NW.
-∎
 Input: An instance $(T,N,(v_{i})_{i\in N})$ with submodular binary valuation functions
-Output: A non-wasteful allocation $A$
-1 Initialize: Let $A=(\emptyset,\ldots,\emptyset)$;
+Output: A NW allocation $\bf A$}
+1 Initialize: Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
 2 when item $e\in T$ arrives do
@@
 5
-6            if $\Delta_{e}^{i}(A_{i})>0$ then
+6    if $\Delta_{A_{i}}^{i}({e})>0$ then
 7
@@
 15
-16return $A$;
+16return $\bf A$};
 Algorithm 1 Greedy Algorithm
+Given an instance $(T,N,(v_{i})_{i\in N})$, Algorithm 1 can compute an NW allocation.
+When item $e_{k}$ ($k\in[t]$) arrives, if there exists an agent $j\in N$ such that $\Delta_{A_{j}^{k-1}}^{j}({e_{k}})=1$, $e_{k}$ will be allocated.
+We assume that $e_{k}$ is allocated to agent $i$.
+We will show that $\Delta_{A_{i}^{r}\setminus\{e_{r}\}}^{i}({e_{k}})=1$ holds for any round $r>k$ after some new items are added to agent $i$’s bundle.
+Without loss of generality, we assume that $e_{s}$ is an arbitrarily newly added item in agent $i$’s bundle, where $s>k$, it holds that $\Delta_{A_{i}^{s-1}}^{i}({e_{s}})=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1})=1$.
+Then, we have
+$\Delta_{{A_{i}^{s}}\setminus\{e_{r}\}}^{i}({e_{k}})$ $=v_{i}(A_{i}^{s-1}\cup\{e_{s}\})-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
+$=v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\cup\{e_{s}\}\setminus\{e_{k}\})$
+$\geq v_{i}(A_{i}^{s-1})+1-v_{i}(A_{i}^{s-1}\setminus\{e_{k}\})-v_{i}(e_{s})$
+$=v_{i}(A_{i}^{s-1})-v_{i}(A^{s-1}\setminus\{e_{k}\})$
+$=\Delta_{A_{i}^{s-1}\setminus\{e_{k}\}}^{i}({e_{k}}),$
+where the inequality holds for submodularity.
+Following this induction by repeatedly utilizing submodularity, we have that $\Delta_{A_{i}^{s}\setminus\{e_{k}\}}^{i}({e_{k}})\geq\ldots\geq\Delta_{A_{i}^{k}\setminus\{e_{k}\}}^{i}({e_{k}})=\Delta_{A_{i}^{k-1}}^{i}({e_{k}})=1$.
+Since $v_{i}$ is a submodular binary valuation function, it holds that $\Delta_{A_{i}^{s}\setminus\{e_{k}\}}^{i}({e_{k}})\leq 1$.
+Thus, we can conclude that $\Delta_{A_{i}^{s}\setminus\{e_{k}\}}^{i}({e_{k}})=1$.
+∎
+Lemma 2.
+For submodular valuation functions, there exists an instance that no deterministic online algorithm can compute an NW allocation, even for two agents with identical valuation functions.
+Proof.
+We construct an instance with two agents, $v_{1}=v_{2}=v$.
+When first item $e_{1}$ arrives, $v(e_{1})=3$.
+W.l.o.g., we allocate $e_{1}$ to agent 1.
+When the second item $e_{2}$ arrives, $v(e_{2})=3,v(\{e_{1},e_{2}\})=5$.
+Case (1): we allocate $e_{2}$ to agent 2, for the third item $e_{3}$$v(e_{3})=5,v(\{e_{1},e_{3}\})=v(\{e_{2},e_{3}\})=v(\{e_{1},e_{2},e_{3}\})=5$.
+We have that $v(\{e_{1},e_{3}\})-v(e_{3})=v(\{e_{2},e_{3}\})-v(e_{3})=0$.
+So, no matter who is assigned $e_{3}$, it cannot satisfy NW.
+Case (2): we allocate $e_{2}$ to agent 1, for the third item $e_{3}$ with $v(e_{3})=5,v(\{e_{1},e_{3}\})=v(\{e_{2},e_{3}\})=v(\{e_{1},e_{2},e_{3}\})=5$.
+To ensure NW, we must allocate $e_{3}$ to agent 2.
+For the fourth item $e_{4}$, $v(e_{4})=v(\{e_{1},e_{4}\})=v(\{e_{2},e_{4}\})=v(\{e_{3},e_{4}\})=v(\{e_{1},e_{2},e_{4}\})=v(\{e_{1},e_{3},e_{4}\})=v(\{e_{2},e_{3},e_{4}\})=v(\{e_{1},e_{2},e_{3},e_{4}\})=6$.
+We have that $v(\{e_{1},e_{2},e_{4}\})-v(\{e_{2},e_{4}\})=v(\{e_{3},e_{4}\})-v(e_{4})=0$.
+So, no matter who is assigned $e_{4}$, it cannot be NW.
+∎
 3.1 The Impossibilities of Approximate EF1 and MMS
+(Aleksandrov et al., 2015) show that in the online setting, an EF1 allocation exists for additive binary valuations, where for each agent $i\in N$, $v_{i}$ is additive and $v_{i}(e)\in\{0,1\}$ for any $e\in T$.
 In this part, we present a strong negative result regarding the approximations of EF1 and MMS for general additive valuations.
-The valuation function $v_{i}$ is tri-valued if $v_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
+We say that the valuation function $v_{i}$ is additive personalized tri-valued if $v_{i}$ is additive and $v_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
 Theorem 1.
-For general additive functions, no deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any constant $\alpha>0$, even for two agents with additive tri-valued valuation functions.
+No deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any $\alpha>0$, even for two agents with additive personalized tri-valued valuation functions.
 Proof.
+$e_{1}$ $e_{2}$ $e_{3}$
+agent $1$ 1 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
+agent 2 11 $\epsilon$ $\frac{1}{\epsilon}$
+Table 4: The impossibility result of EF1 and MMS.
+Given any $\alpha>0$, let $\epsilon$ be a constant arbitrarily close to 0 such that $0<\epsilon<\alpha$.
 Consider an online instance with two agents.
-For any agent $i\in N$ and item $e\in T$, $v_{i}(e)\in\{\epsilon,1,\frac{1}{\epsilon}\}$, where $10<\epsilon<$ 1.
+For any agent $i\in N$ and item $e\in T$, $v_{i}(e)\in\{\epsilon,1,\frac{1}{\epsilon}\}$, where $0<\epsilon<1$.
 The items arrive online, and the valuation of each item is shown in Table 4.
-When $t=1$, the arriving item has the value of 1 for both agents.
-Without loss of generality, agent 1 picks it.
-When $t=2$, the item that has the value of $\frac{1}{\epsilon}$ for agent 1 and $\epsilon$ for agent 2 arrives.
-In order to avoid 00-EF1 or 00-MMS, this item must be allocated to agent 2.
-In round $t=3$, the arriving item has the value of $\frac{1}{\epsilon}$ for both agents.
-In this case, if this item is allocated to agent 1, for agent 2, the approximation ratio of EF1 is $\epsilon$, and the approximation ratio of MMS is $\frac{\epsilon}{1+\epsilon}$; if this item is allocated to agent 2, for agent 1, the approximation ratio of EF1 is $\frac{1}{\frac{1}{\epsilon}}=\epsilon$ and the approximation ratio of MMS is $\frac{1}{1+\frac{1}{\epsilon}}=\frac{\epsilon}{1+\epsilon}$.
-Therefore, for any constant $\alpha>0$, let $\epsilon<\alpha$, and we can conclude that no algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS for any given $\alpha>0$.
+When $t=1$, the item $e_{1}$ arrives, $T^{1}=\{e_{1}\}$.
+Without loss of generality, suppose it is allocated to agent 1, $A^{1}_{1}=\{e_{1}\},A^{1}_{2}=\emptyset$.
+When $t=2$, the item $e_{2}$ arrives, $T^{2}=\{e_{1},e_{2}\}$.
+To maintain the fairness of allocation, we must allocate it to agent 2, i.e., $A^{2}_{1}=\{e_{1}\},A^{2}_{2}=\{e_{2}\}$.
+Otherwise, agent 1 gets $\{e_{1},e_{2}\}$ and agent 2 gets an empty set, $v_{2}(\emptyset)=0<v_{2}(\{e_{1},e_{2}\}\setminus\{e\})$ for any $e\in\{e_{1},e_{2}\}$ and $MMS_{2}(T^{2})=\epsilon$.
+It is 0-EF1 and 0-MMS.
+When $t=3$, item $e_{3}$ arrives, $T^{3}=\{e_{1},e_{2},e_{3}\}$.
+$MS_{1}(T^{3})=\frac{1}{\epsilon}, MM$$MS_{2}(T^{3})=1+\epsilon$.
+If we allocate it to agent 1, i.e., $A^{3}_{1}=\{e_{1},e_{3}\},A^{3}_{2}=\{e_{2}\}$, we have $v_{2}(A^{3}_{2})=\epsilon$, $v_{2}(A^{3}_{1}\setminus\{e_{1}\})=\frac{1}{\epsilon}$, and $v_{2}(A^{3}_{1}\setminus\{e_{3}\})=1$.
+Thus, $\bf A^{3}$} is $\epsilon$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS.
+If $e_{3}$ is allocated to agent 2, i.e., $A^{3}_{1}=\{e_{1}\},A^{3}_{2}=\{e_{2},e_{3}\}$, we have $v_{1}(A^{3}_{1})=1=\epsilon\cdot v_{1}(A^{3}_{2}\setminus\{e\})$ for any $e\in A^{3}_{2}$.
+Thus, $\bf A^{3}$} is $\epsilon$-EF1 and $\epsilon$-MMS.
+Therefore, in the above instance, no algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS.
 This result can be generalized to more than two agents, and we defer the proof to Proposition 1.
-$t=1$ $t=2$ $t=3$
-Agent $1$ $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
-Agent 2 $1$ $\epsilon$ $\frac{1}{\epsilon}$
-Table 3: The impossibility results of EF1 and MMS for additive tri-valued valuations
-By Theorem 1, we cannot make further progress in the approximation of EF1 and MMS for tri-valued valuation functions and beyond in the online allocation problem.
-This means that we can only focus our efforts on bi-values.
-In the following parts, we focus on two classes of valuation functions, including submodular binary and additive bi-valued valuation functions, and study whether there are positive results about the approximation of EF1 or MMS.
+∎
+Proposition 1.
+For general additive functions, no deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any $\alpha>0$ when there are more than two agents with additive personalized tri-valued valuation functions.
+Proof.
+$e_{1}$ $e_{2}$ $\ldots$ $e_{n-2}$ $e_{n-1}$ $e_{n}$ $e_{n+1}$
+agent 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
+agent 22 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
+$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
+agent n2n-2 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
+agent n1n-1 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
+agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
+Table 5: The impossibility results of EF1 and MMS for additive personalized tri-valued valuations when there are more than two agents.
+We further extend the impossibility results of approximate EF1 and MMS for $n\geq 3$.
+Consider the following instance with $n\geq 3$ agents, where the value of each item is shown in Table 5.
+In the first (n2)(n-2) rounds, to avoid 0-EF1 or 0-MMS, each item is allocated to the agent with the empty bundle.
+For the item that arrives before round t=n2t=n-2, it has the value of $\frac{1}{\epsilon}$ for any agent.
+Without loss of generality, we assume that agent $i$ picks the item in the round $t=i$, where $i\leq n$.
+In round t=n1t=n-1, the arriving item has the value of 1 for any agent, and we assume that it is allocated to agent n1n-1.
+In round $t=n$, the arriving item has the value of $\epsilon$ for agent $i\in N\setminus\{n-1\}$ and $\frac{1}{\epsilon}$ for agent n1n-1, and it should be allocated to agent $n$ to avoid 0-EF1 or 0-MMS.
+Finally, in the round $t=n+1$, the item with the value of $\frac{1}{\epsilon}$ for each agent arrives.
+Due to non-wastefulness, this item must be allocated to one agent.
+If this item is allocated to agent $i\in N\setminus\{n-1,n\}$, it is not hard to see that the allocation is $\epsilon^{2}$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS.
+If this item is allocated to agent n1n-1, this allocation is $\epsilon$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS.
+If this item is allocated to agent $n$, we can derive that the competitive ratio of EF1 is $\epsilon$ and MMS is $\epsilon$.
+Combining the above cases, we can conclude that the upper bound of EF1 is $\epsilon$ and MMS is $\epsilon$.
+For given $\alpha>0$, let $\epsilon<\alpha$, we can conclude that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS for any given $\alpha>0$.
+∎
+By Theorem 1, we cannot make further progress in the approximations of EF1 and MMS for additive personalized tri-valued valuation functions and beyond in the online setting.
+In the following parts, we focus on two classes of valuation functions that are more general than additive binary functions, including submodular binary and additive personalized bi-valued valuation functions, and study whether there are positive results about the approximations of EF1 and MMS.
 3.2 Submodular Binary Valuations
-This part focuses on submodular binary valuation functions, where for each agent $i\in N$, her valuation function $v_{i}$ is submodular and the marginal value of any item is 0 or 1.
-For a subclass of submodular binary valuation functions based on maximum (unweighted) bipartite matching, which is also called assignment [19] or OXS [20] valuations, Hosseini et al.[11] show that online allocations satisfying $\frac{1}{2}$-EF1 and $\frac{1}{2}$-MMS exist.
-In their model, there are some groups, each of which includes some members, and each member can only receive one item at most.
-When an item arrives, they greedily match the item to some member who likes the item and has never received any items.
-Leaving the online matching model, we consider the general online allocation problem, greedily assigning each arriving item to the agent that brings the maximum marginal utility.
-Deriving from the greedy technique of the submodular optimization problem and the sequential allocating technique to guarantee fairness in [11], we propose an algorithm named Marginal-Greedy Algorithm (Algorithm 2) for general binary valuation functions.
-We prove that input submodular binary valuations into Algorithm 2, we can get an NW allocation that is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW.
-Additionally, we conjecture that the Algorithm 2 will also perform well when we input subadditive binary valuation functions into Algorithm 2, and leave this conjecture as an open question.
-Input: An instance $(T,N,(v_{i})_{i\in N})$ with general binay valuation functions
-Output: A non-wasteful approximate EF1 and approximate MMS allocation $A$
-1 Initialize: let $\pi=(\pi_{1},\cdots,\pi_{n})$ be an arbitrary order of $n$ agents;
-2Let $A=(\emptyset,\ldots,\emptyset)$;
+This part focuses on submodular binary valuation functions, where for each agent $i\in N$, $v_{i}$ is submodular and the marginal value of any item is 0 or 1.
+For a subclass of submodular binary valuation functions based on maximum (unweighted) bipartite matching, also known as assignment Munkres (1957) or OXS Lehmann et al.
+(2001) valuations, Hosseini et al.
+(2024) show that online allocations satisfying $\frac{1}{2}$-EF1 and $\frac{1}{2}$-MMS exist.
+Limitation of Hosseini et al.
+(2024): Their $\frac{1}{2}$-EF1/MMS guarantees are restricted to OXS valuations, where items must be assigned within predefined groups under rigid matching constraints.
+Each member in the group can only receive at most one item, and once someone gets an item, it cannot be changed.
+Thus, their algorithm can only handle the OXS valuation function and its subclasses.
+Moving beyond the online allocation model based on matching, we consider the general online allocation problem.
+Building upon the effective greedy technique in the submodular optimization and the sequential allocating technique to guarantee fairness in Hosseini et al.
+(2024), we propose the Marginal-Greedy Algorithm (Algorithm 2), greedily assigning each arriving item to the agent who receives the maximum marginal value, for submodular binary valuation functions, and then show that the Marginal-Greedy Algorithm computes an NW allocation that is $\frac{1}{2}$-EF1, $\frac{1}{2}$-MMS, and $\frac{1}{2}$-max-USW.
+Input: An instance $(T,N,(v_{i})_{i\in N})$ with submodular binary valuation functions
+Output: An NW approximate EF1 and approximate MMS allocation $\bf A$}
+1 Initialize: let $\pi=(\pi_{1},\ldots,\pi_{n})$ be an arbitrary order of $n$ agents;
+2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
 3 when item $e\in T$ arrives do
@@
 9        Let $A_{\pi_{i}}=A_{\pi_{i}}\cup\{e\}$;
-10                  $\pi=(\pi_{1},\cdots,\pi_{i-1},\pi_{i+1},\cdots,\pi_{n},\pi_{i})$;
+10       $\pi=(\pi_{1},\ldots,\pi_{i-1},\pi_{i+1},\ldots,\pi_{n},\pi_{i})$;
 11        break;
@@
 17
-18return $A$;
+return $\bf A$};
 Algorithm 2 Marginal-Greedy Algorithm
-In the analysis of results about submodular binary valuation functions, we skillfully use submodularity and the properties of the matroid.
-Then, we introduce the relationship of submodularity and matroid: a function is submodular binary if and only if it is a rank function of a matroid.
+In the analysis of results about submodular binary valuation functions, we effectively leverage submodularity and the properties of the matroid.
+There exists a relationship of submodularity and matroid: a function is submodular binary iff it is a rank function of a matroid Schrijver et al.
+(2003).
 Definition 5.
@@
 Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, where $v_{i}$ is submodular and binary for all $i\in N$.
-Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ be the set of items that have already arrived, denoted by ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ and $A^{k}_{0}$ the allocation and the set of all unallocated items at the end of round $k$ respectively.
-In the algorithm procedure, the arriving item is allocated to an agent who thinks that the marginal utility of this item is $1$ if such an agent exists.
-By Claim 1, $A^{k}_{i}$ is non-wasteful.
-For each submodular binary valuation function $v_{i}$, there exists a matroid such that $v_{i}$ is the rank function of this matroid.
-let $\mathcal{M}_{i}=(T,\mathcal{F}_{i})$ be the corresponding matroid of $v_{i}$.
+Fix an arbitrary round $k$, let $T^{k}=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by ${\bf A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$ and $A^{k}_{0}$ the allocation and the set of all unallocated items at the end of round $k$ respectively.
+In Algorithm 2, the arriving item is allocated to an agent who thinks that the marginal value of this item is $1$ if there exists such an agent.
+By Lemma 1, $\textbf{A}^{k}$ is NW.
+For each $v_{i}$, there exists a matroid $\mathcal{M}_{i}=(T^{k},\mathcal{F}_{i})$ such that $v_{i}$ is the rank function of $\mathcal{M}_{i}$.
 We have $v_{i}(A^{k}_{i})=|A^{k}_{i}|$, and $A^{k}_{i}\in\mathcal{F}_{i}$.
-Firstly, we show that ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ is $\frac{1}{2}$-EF1.
-Consider any two agents $i,j\in N$, $i\neq j$.
+First, we show that ${\bf A}^{k}$ is $\frac{1}{2}$-EF1.
+Consider any two agents $\in N, iji\neq j$.
 If $v_{i}(A^{k}_{j})>v_{i}(A^{k}_{i})$, by the property (3) of the matroid, there must exist an item $e\in A^{k}_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$.
@@
 We will show $|B_{j}|\leq|A^{k}_{i}|+1$.
-Assume that $|B_{j}|\geq|A^{k}_{i}|+2$, $j$ has priority higher than $i$ at least $|B_{j}|\geq|A^{k}_{i}|+2$ times, which contradicts the order in the algorithm.
+Assume that $|B_{j}|\geq|A^{k}_{i}|+2$, $j$ has the priority higher than $i$ at least $|B_{j}|\geq|A^{k}_{i}|+2$ times, which contradicts the order in the algorithm.
 Combining the property (3) of the matroid, we have $v_{i}(A^{k}_{j})=|A^{k}_{i}|+|B_{j}^{*}|\leq|A^{k}_{i}|+|B_{j}|\leq 2|A^{k}_{i}|+1$.
-Thus, ${A^{k}}$ is $\frac{1}{2}$-EF1.
-Secondly, we show that ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ is $\frac{1}{2}$-MMS.
-Consider a MMS partition $X=(X_{1},\cdots,X_{n})$ for agent $i$ of item set $T$, we assume that the share of agent $i$ in ${A^{k}}$ is less than half of the maximin share of $i$, i.e.
+Thus, ${\bf A}^{k}$ is $\frac{1}{2}$-EF1.
+Second, we show that ${\bf A}^{k}$ is $\frac{1}{2}$-MMS.
+Consider a MMS partition ${\bf X}=(X_{1},\ldots,X_{n})$ for agent $i$ of item set $T$, we assume that the share of agent $i$ in ${\bf A}^{k}$ is less than half of the maximin share of $i$, i.e.
 $v_{i}(A^{k}_{i})\leq\frac{1}{2}v_{i}(X_{j})$ for any $j\in N$.
 By the property (3) of the matroid, there exists an item $e\in X_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$.
-Let $C_{j}=\{e:e\in X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$, $C_{j}^{*}\in\arg\max\{|S|:S\subseteq C_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$.
-We have that $v_{i}(X_{j})=v_{i}(A^{k}_{i}\cup C_{j}^{*})=v_{i}(A^{k}_{i})+v_{i}(C_{j}^{*})=|A^{k}_{i}|+|C_{j}^{*}|$.
-Since $v_{i}(A^{k}_{i})=|A^{k}_{i}|\leq\frac{1}{2}v_{i}(X_{j})$, $|C_{j}|\geq|C_{j}^{*}|>|A^{k}_{i}|+1$.
-Then, $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$.
-We have $|\cup_{j\in N}C_{j}|=|\{e:e\in\cup_{j\in N}X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\{e:e\in T,A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\cup_{j\in N}B_{j}|\leq(n-1)(|A^{k}_{i}|+1)$, it contradicts the inequality $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$.
-Thus, ${A^{k}}$ is $\frac{1}{2}$-MMS.
-Finally, we show that ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ is $\frac{1}{2}$-USW.
+Let $C_{j}=\{e:e\in X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$ and $C_{j}^{*}\in\arg\max\{|S|:S\subseteq C_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$.
+Then, we have $v_{i}(X_{j})=v_{i}(A^{k}_{i}\cup C_{j}^{*})=v_{i}(A^{k}_{i})+v_{i}(C_{j}^{*})=|A^{k}_{i}|+|C_{j}^{*}|$.
+Since $v_{i}(A^{k}_{i})=|A^{k}_{i}|\leq\frac{1}{2}v_{i}(X_{j})$ holds, we get $|C_{j}|\geq|C_{j}^{*}|>|A^{k}_{i}|+1$.
+Then, summing up all agents, we have $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$.
+At last, we get $|\cup_{j\in N}C_{j}|=|\{e:e\in\cup_{j\in N}X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\{e:e\in T,A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\cup_{j\in N}B_{j}|\leq(n-1)(|A^{k}_{i}|+1)$, which contradicts the inequality $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$.
+Thus, ${\bf A}^{k}$ is $\frac{1}{2}$-MMS.
+Finally, we show that ${\bf A}^{k}$ is $\frac{1}{2}$-USW.
 We make a slight adjustment to the Algorithm 2: when a new item arrives, if the marginal utility is 0 for all agents, then assign the item to agent $1$ without updating the agents’ order.
-Then, we will show that ${A^{k^{\prime}}}=(A^{k}_{1}\cup A^{k}_{0},\cdots,A^{k}_{n})$ is the output of the adjusted algorithm.
-We only need to prove that in each round $s\in[k]$, if the marginal utility of the arrived item $e_{s}$ to $A^{s-1}_{1}$ is $1$, $v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1$, by the monotonicity of $v_{i}$, $v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0}\cup\{e_{s}\})\geq v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1=v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0})+1$.
-Then, the marginal utility to $A^{s-1}_{1}\cup A^{s-1}_{0}$ is also $1$.
-So, ${A^{k^{\prime}}}=(A^{k}_{1}\cup A^{t}_{0},\cdots,A^{k}_{n})$ is the output of the adjusted algorithm.
-Now, we will show that ${A^{k^{\prime}}}$ is $\frac{1}{2}$-USW.
-The remaining proof part is based on the idea from [20].
+Then, we will show that ${\bf A}^{k^{\prime}}=(A^{k}_{1}\cup A^{k}_{0},\ldots,A^{k}_{n})$ is the output of the adjusted algorithm.
+We only need to prove that in each round $s\in[k]$, if the marginal value of the arrived item $e_{s}$ to $A^{s-1}_{1}$ is 11, we have $v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1$.
+By the monotonicity of $v_{i}$, we get $v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0}\cup\{e_{s}\})\geq v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1=v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0})+1$, implying that the item $e_{s}$’s marginal value to $A^{s-1}_{1}\cup A^{s-1}_{0}$ is $1$.
+Thus, ${\bf A}^{k^{\prime}}=(A^{k}_{1}\cup A^{t}_{0},\ldots,A^{k}_{n})$ is the output of the adjusted algorithm.
+Since $v_{1}(A^{k}_{1}\cup A^{t}_{0})=v_{1}(A^{k}_{1})$, we have $USW({\bf A}^{k})=USW({\bf A}^{k^{\prime}})$.
+In other words, we only need to show that ${\bf A}^{k^{\prime}}$ is $\frac{1}{2}$-max-USW.
+The remaining proof part is based on the idea from Lehmann et al.
+(2001).
 Consider the first item $e_{1}$, we assume that $e_{1}$ is allocated to agent $i^{*}$ during the execution of the algorithm and $v_{i^{*}}(e_{1})=1$, $A_{i^{*}}^{1}=\{e_{1}\}$ and $A_{i}^{1}=\emptyset$ for $i\neq i^{*}$.
-Let ${Y}=(Y_{1},\cdots,Y_{n})$ be an allocation achieving the optimal USW of instance $I=(T=\{e_{1},\cdots,e_{k}\},N,(v_{i})_{i\in N})$, and we assume that $e_{1}$ is allocated to agent $i^{o}$ in $Y$, $e_{1}\in Y_{i^{o}}$.
-Next, we construct a new instance ${I^{(1)}}=(T^{\prime},N,(v_{i}^{(1)})_{i\in N})$, where $v_{i^{*}}^{(1)}={\Delta^{i^{*}}_{e_{1}}}$ and $v_{i}^{(1)}=v_{i}$ for all $i\neq i^{*}$.
-In $I^{(1)}$, $T^{\prime}=T\setminus\{e_{1}\}$, that is, the first item arriving is $e_{2}$.
-Input the instance $I^{(1)}$ into the adjusted algorithm.
-Let initialized order of $N$ be $\pi=(\pi_{1},\cdots,\pi_{n},\pi_{i^{*}})$, the initialized allocation be ${D^{0}}=(D^{0}_{1},\cdots,D^{0}_{n})$ , where $D_{i^{*}}=\{e_{1}\}$ and $D_{i}=\emptyset$ for all $i\neq i^{*}$.
-Assume ${D^{k}}=(D^{k}_{1},\cdots,D^{k}_{n})$ is the allocation at the end of round $k$.
-We have ${D^{k}}={A^{k^{\prime}}}$, then $USW({D^{k}})=USW({A^{k^{\prime}}})=USW({A^{k}})$.
-Now, we costruct an allocation ${S}=(S_{1},\cdots,S_{n})$ for instance $I^{\prime}$ as follows: $S_{i^{o}}=Y_{i^{o}}\setminus\{e_{1}\}$, $S_{i}=Y_{i}$ for all $i\neq{i^{o}}$.
-It holds $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})=v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$.
-By the monotonicity of $v_{i^{*}}$, $v_{i^{*}}^{(1)}(S_{i^{*}})=v_{i^{*}}(Y_{i^{*}}\cup\{e_{1}\})-v_{i^{*}}(\{e_{1}\})\geq v_{i^{*}}(Y_{i^{*}})-v_{i^{*}}(\{e_{1}\})$.
-$v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$, where the first inequality holds because of the submodularity and the second inequality holds because of $v_{i^{o}}(\{e_{1}\})\leq v_{i^{*}}(\{e_{1}\})$, Let ${Y^{(1)}}=(Y_{1}^{(1)},\cdots,Y_{n}^{(1)})$ be an allocation achieving the optimal USW of instance $I^{(1)}$.
-It holds that $USW({Y^{(1)}})\geq USW(S)=\cup_{i\in N}v_{i}^{(1)}(S_{i})=\cup_{i\in N,i\neq i^{*},i\neq i^{o}}v_{i}^{(1)}(S_{i})+v_{i^{o}}^{(1)}(S_{i^{o}})+v_{i^{*}}^{(1)}(S_{i^{*}})\geq USW(Y)-2v_{i^{*}}(\{e_{1}\})$.
-By lemma 1 in [20], $v_{i^{*}}^{(1)}$ is also submodular valuation.
-So, by constructing the new instances $I^{(2)},\cdots,I^{(k-1)}$ and inducting each round based on the items that arrive, we have $USW(A^{k})=USW(A^{k^{\prime}})=USW(D^{k})=USW(Y^{k})\geq\frac{1}{2}USW(Y).$
+Let ${\bf Y}=(Y_{1},\ldots,Y_{n})$ be an allocation achieving the optimal USW of instance $I=(T^{k},N,(v_{i})_{i\in N})$, and we assume that $e_{1}$ is allocated to agent $i^{o}$ in $Y$, $e_{1}\in Y_{i^{o}}$.
+Next, we construct a new instance ${I^{(1)}}=(T^{\prime},N,(v_{i}^{(1)})_{i\in N})$, where $v_{i^{*}}^{(1)}={\Delta^{i^{*}}_{e_{1}}}$ and $v_{i}^{(1)}=v_{i}$ for all $i\neq i^{*}$, $T^{\prime}=T^{k}\setminus\{e_{1}\}$, that is, the first item arriving is $e_{2}$, and then input the instance $I^{(1)}$ into the adjusted algorithm.
+Let initialized order of $N$ be $\pi=(\pi_{1},\ldots,\pi_{n},\pi_{i^{*}})$, the initialized allocation be ${\bf D}^{0}=(D^{0}_{1},\ldots,D^{0}_{n})$ , where $D_{i^{*}}=\{e_{1}\}$ and $D_{i}=\emptyset$ for all $i\neq i^{*}$.
+Assume that ${\bf D}^{k}=(D^{k}_{1},\ldots,D^{k}_{n})$ is the allocation at the end of round $k$.
+We have ${\bf D}^{k}={\bf A}^{k^{\prime}}$, then $USW({\bf D}^{k})=USW({\bf A}^{k^{\prime}})=USW({\bf A}^{k})$.
+Now, we construct an allocation ${\bf S}=(S_{1},\ldots,S_{n})$ for instance $I^{\prime}$ as follows: $S_{i^{o}}=Y_{i^{o}}\setminus\{e_{1}\}$, $S_{i}=Y_{i}$ for all $i\neq{i^{o}}$.
+It holds that $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})=v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$.
+By the monotonicity of $v_{i^{*}}$, we have $v_{i^{*}}^{(1)}(S_{i^{*}})=v_{i^{*}}(Y_{i^{*}}\cup\{e_{1}\})-v_{i^{*}}(\{e_{1}\})\geq v_{i^{*}}(Y_{i^{*}})-v_{i^{*}}(\{e_{1}\})$.
+$v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$, where the first inequality follows from the submodularity and the second inequality follows from $v_{i^{o}}(\{e_{1}\})\leq v_{i^{*}}(\{e_{1}\})$.
+Let ${\bf Y}^{(1)}=(Y_{1}^{(1)},\ldots,Y_{n}^{(1)})$ be an allocation achieving the optimal USW of instance $I^{(1)}$.
+It holds that $USW({\bf Y}^{(1)})\geq USW({\bf S})=\cup_{i\in N}v_{i}^{(1)}(S_{i})=\cup_{i\in N,i\neq i^{*},i\neq i^{o}}v_{i}^{(1)}(S_{i})+v_{i^{o}}^{(1)}(S_{i^{o}})+v_{i^{*}}^{(1)}(S_{i^{*}})\geq USW({\bf Y})-2v_{i^{*}}(\{e_{1}\})$.
+By lemma 1 in Lehmann et al.
+(2001), $v_{i^{*}}^{(1)}$ is also submodular.
+So, by constructing the new instances $I^{(2)},\ldots,I^{(k-1)}$ and inducting each round based on the items that arrive, we have $USW({\bf A}^{k})=USW({\bf A}^{k^{\prime}})=USW({\bf D}^{k})=USW({\bf Y}^{k})\geq\frac{1}{2}USW({\bf Y}).$
 Theorem 3.
-For submodular binary valuation functions, no deterministic online algorithm can achieve $(\frac{1}{2}+\epsilon)$-EF1 or $(\frac{1}{2}+\epsilon)$-MMS, and non-wastefulness simultaneously for any $\epsilon>0$.
+For any $\epsilon>0$, with submodular binary valuation functions, no deterministic online algorithm can achieve
+  • $(\frac{1}{2}+\epsilon)$-EF1 and non-wastefulness,
+  • or $(\frac{1}{2}+\epsilon)$-MMS and non-wastefulness,
+  • or $(\frac{1}{2}+\epsilon)$-max-USW and non-wastefulness.
 Proof.
 Figure 1: An instance with $\frac{1}{2}$-EF1 and $\frac{1}{2}$-MMS allocations
-Consider an online allocation instance, where there are two agents $\{1,2\}$ with identical valuations $v$, and the online items belong to several categories $C_{1},\cdots,C_{l}$.
+First, let us consider EF1 or MMS.
+Consider an online allocation instance, where there are two agents $\{1,2\}$ with identical valuations $v$, and the online items belong to several categories $C_{1},\ldots,C_{l}$.
 In Figure 1, categories are distinguished by colours, and edges of different colours represent items of different categories.
-Denote a partition matroid $\mathcal{M}=(T,\mathcal{F})$, =$\mathcal{F}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~{}k\in[l]\}$.
+Denote a partition matroid $\mathcal{M}=(T,\mathcal{F})$, $\mathcal{F}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~k\in[l]\}$.
 The valuation $v$ is the rank function of the matroid $\mathcal{M}$; for any item set $S$, the value of $S$ is the number of items of the maximal independent set of $S$.
-For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are $1$.
+For the first arrived item $e_{1}$, the marginal values for agent $1$ and agent $2$ are $1$.
 Without loss of generality, we allocate item $e_{1}$ to agent $1$.
-When the second item $e_{2}$ arrives, to avoid breaking fairness, we must allocate it to agent $2$.
-For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are $1$.
+When the second item $e_{2}$ arrives, in order not to break fairness, we must allocate it to agent $2$.
+For the third item $e_{3}$, the marginal values for agent $1$ and agent $2$ are $1$.
 We will discuss the following two cases.
-Case 1: Allocate $e_{3}$ to agent $2$, shown in Figure 1(1).
-We must allocate the fourth item, $e_{4}$, to agent $2$ due to the requirement of non-wastefulness.
-Then, $A^{4}_{1}=\{e_{1}\}$,$A^{4}_{2}=\{e_{2},e_{3},e_{4}\}$.
+Case 1: Allocate $e_{3}$ to agent $2$, shown in the left graph in Figure 1 Then, $A^{4}_{1}=\{e_{1}\}$$A^{4}_{2}=\{e_{2},e_{3},e_{4}\}$.
 We have that, $v(A^{4}_{1})=\frac{1}{2}v(A^{4}_{2}\setminus\{e\})$ for any $e\in A^{4}_{2}$, and $v(A^{4}_{1})=\frac{1}{2}\mathsf{MMS}_{1}$.
-Case 2: Allocate $e_{3}$ to agent $1$, shown in Figure 1(2); we must allocate the fourth item, $e_{4}$, to agent $1$ due to the requirement of non-wastefulness.
-Then, $A^{4}_{1}=\{e_{1},e_{3},e_{4}\}$,$A^{4}_{2}=\{e_{2}\}$.
+Case 2: Allocate $e_{3}$ to agent $1$, shown in the right graph in Figure 1; we must allocate the fourth item, $e_{4}$, to agent $1$ due to the requirement of non-wastefulness.
+Then, $A^{4}_{1}=\{e_{1},e_{3},e_{4}\}$$A^{4}_{2}=\{e_{2}\}$.
 We have that, $v(A^{4}_{2})=\frac{1}{2}v(A^{4}_{1}\setminus\{e\})$ for any $e\in A^{4}_{1}$, and $v(A^{4}_{2})=\frac{1}{2}\mathsf{MMS}_{2}$.
-∎
-Theorem 4.
-For submodular binary valuation functions, no deterministic online algorithm can achieve $(\frac{1}{2}+\epsilon)$-max-USW and non-wastefulness for any $\epsilon>0$.
-Proof.
-Consider an online allocation instance with two agents $\{1,2\}$.
-For the first arrived item $e_{1}$, the marginal utilities for agents $1$ and $2$ are both $1$, $v_{1}(\{e_{1}\})=v_{2}(\{e_{1}\})=1$.
+Next, let us consider USW.
+Consider an online allocation instance, where there are two agents $\{1,2\}$.
+For the first arrived item $e_{1}$, the marginal values for agents $1$ and $2$ are both $1$, $v_{1}(\{e_{1}\})=v_{2}(\{e_{1}\})=1$.
 Without loss of generality, we allocate $e_{1}$ to agent $1$.
-When the second item $e_{2}$ arrives, $v_{1}(\{e_{2}\})=v_{1}(\{e_{1},e_{2}\})=1$,$v_{2}(\{e_{2}\})=0$, the marginal utilities for agents $1$ and $2$ are both 00.
+When the second item $e_{2}$ arrives, $v_{1}(\{e_{2}\})=v_{1}(\{e_{1},e_{2}\})=1$$v_{2}(\{e_{2}\})=0$, the marginal values for agents $1$ and $2$ are both 0.
 No matter which agent gets $e_{2}$, $USW(A^{2})=1$.
@@
 ∎
-There is a side result of our algorithm, where the Marginal-Greedy Algorithm (Algorithm 2) can also work for additive binary valuation functions to compute an NW allocation that is EF1, MMS, and max-USW, where for each agent $i\in N$, $v_{i}$ is additive and $v_{i}(e)\in\{0,1\}$ for any $e\in T$.
-3.3 Additive Bi-valued Valuations
-In this part, we consider additive bi-valued valuation functions.
-For any agent $i\in N$, $v_{i}$ is bi-valued if $v_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$.
-Someone may wonder whether a non-wasteful EF1 or MMS allocation can be found with additive bi-valued valuation functions because the additive binary valuation function can be seen as a special additive bi-valued valuation function.
-However, we find that even for two agents with additive bi-valued valuations, a non-wasteful EF1 or MMS allocation cannot be guaranteed.
+For an independent interest, we find that there is a side result of the Marginal-Greedy Algorithm, where it can also work for additive binary valuation functions to compute an NW allocation that is EF1, MMS, and max-USW, which demonstrates the robustness of our algorithm.
+3.3 Additive Personalized Bi-valued Valuations
+In this part, we consider the additive personalized bi-valued valuation function, which is often considered in the literature, such as Ebadian et al.
+(2022); Aziz et al.
+(2023).
+For any agent $i\in N$, her valuation function $v_{i}$ is additive personalized bi-valued if $v_{i}$ is additive and $v_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$.
+It is worth investigating whether an NW allocation satisfying EF1 or MMS is achievable with additive personalized bi-valued valuation functions, as the binary case is a special type of bi-valued function.
+However, we find that even for two agents with additive personalized bi-valued valuations, an NW allocation satisfying EF1 or MMS cannot be guaranteed.
 3.3.1 Two Agents
-We study the setting of additive bi-valued valuations for two agents and establish the best possible result for the approximation of EF1 and MMS simultaneously.
-In order to achieve this goal, we propose the Adapted Envy-Graph Procedures (Algorithm 3).
-Note that there are two technical difficulties for us when designing this approximation algorithm.
-One is that we do not know the value of future items.
-The other is that we cannot do any reallocation.
-The main idea of our algorithm is that we take the Adapted Envy-Graph Procedures as the base algorithm to allocate items.
-If we follow the base algorithm, there will be an envy cycle between two agents, and we temporarily terminate it.
-Then, we consider implementing the Preliminary-Breaking-Cycle Algorithm (Algorithm 4) to eliminate the potential envy cycle manually.
-Then, there are two possible results of Algorithm 4.
-One is that we have eliminated the potential envy cycle successfully.
-The other is that there is still an envy cycle, and we execute the Deep-Breaking-Cycle Algorithm (Algorithm 5) to eliminate it further.
-In whichever case, when there is no envy cycle and the allocation is EF1, we go back to the base algorithm.
-The illustration of Algorithms 3, 4 and 5 is shown in Figure 2.
-Before presenting the theorem, we first define an envy graph.
-Figure 2: The illustration of Algorithms 3, 4 and 5.
+We investigate the setting of additive personalized bi-valued valuations for two agents and establish the optimal non-trivial result for simultaneously approximating both EF1 and MMS.
+To achieve this objective, we propose the Adapted Envy-Graph Procedure (Algorithm 3), which requires sophisticated and subtle case analysis due to the uncertainty of future item values.
+The Adapted Envy-Graph Procedure (AEGP) consists of three sub-algorithms: the Envy-Graph Procedure (EGP), the Preliminary-Breaking-Cycle (PBC), and the Deep-Breaking-Cycle (DBC), which are detailed in the Appendix.
+The core structure of our algorithm employs the EGP as the primary procedure.
+However, the EGP has a critical requirement: the input allocation is EF1 and must not contain any envy cycle.
+When an envy cycle is detected, we execute the PBC and DBC algorithms as auxiliary procedures in subsequent rounds to allocate future items more effectively.
+Once the output allocation from either the PBC or DBC algorithm satisfies the EGP algorithm’s precondition, the EGP algorithm resumes execution.
+The execution flow of this integrated process is illustrated in Figure 2.
+Figure 2: The execution flow of EGP, PBC, and DBC algorithms.
+The dashed line represents the execution of the corresponding algorithm.
+The solid line means that when some specific cases occur, we execute another algorithm.
+The main challenge lies in addressing envy cycles effectively.
+This issue is straightforward to resolve in offline settings but becomes significantly more complex in online environments where reallocation is infeasible.
+To overcome this obstacle, we first introduce the PBC algorithm, which is developed through careful analysis of item values over the next two rounds.
+Specifically, the PBC algorithm can be executed consecutively for at most two rounds.
+After these two rounds, the PBC algorithm successfully addresses most cases, achieving EF1 allocation without an envy cycle.
+However, if the allocation remains be $\frac{1}{2}$-EF1, a more strategic approach is required in subsequent rounds to ensure fairer distribution.
+To address this limitation, we introduce the DBC algorithm.
+In the DBC algorithm, we establish allocation rules based on the parity of consecutive execution counts.
+Unlike the PBC algorithm, the DBC algorithm can be executed consecutively across multiple rounds while maintaining the $\frac{1}{2}$-EF1 approximation guarantee, before its output allocation satisfies the precondition of the EGP.
+In summary, both the PBC and DBC algorithms establish specific allocation rules for future items when an envy cycle emerges.
+These rules not only ensure EF1 approximation throughout their execution but also continuously monitor whether the EGP precondition can be satisfied and actively work to achieve it when feasible.
 Definition 6 (Envy Graph for Goods).
-Given an allocation $A^{k}$, in an envy graph G, each vertex corresponds to an agent, and each directed edge $(i,j)$ represents that agent $i$ envies agent $j$, where we have $v_{i}(A_{i}^{k})<v_{i}(A_{j}^{k})$.
-Input: An instance $(T,N,(v_{1},v_{2}))$ with additive bi-valued valuation functions
-Output: An non-wasteful approximate EF1 and approximate MMS allocation $A$
+Given an allocation ${\bf A}^{k}$, the corresponding envy graph is defined as $G=(V,E)$, where the vertex set $V$ corresponds to agent set $N$, and a directed edge $(i,j)\in E$ iff agent $i$ envies agent $j$, i.e., $v_{i}(A_{i}^{k})<v_{i}(A_{j}^{k})$.
+Additionally, the directed cycle in the envy graph is called an envy cycle.
+Input: An instance $(T,N,(v_{1},v_{2}))$ with additive personalized bi-valued valuation functions
+Output: An NW approximate EF1 and approximate MMS allocation $\bf A$}
 1
-2Let $A=(\emptyset,\ldots,\emptyset)$ and build an envy graph $G$ for two agents;
+2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
+Initialize $\alpha=1$; // Algorithm Indicator
+3
+Initialize $(\lambda_{1},\mu_{1})$ and $(\lambda_{2},\mu_{2})$, where $\lambda_{1}=\lambda_{2}=1$ and $\mu_{1}=\mu_{2}=0$; ;
+  // $\lambda$: the times of consecutive execution for the PBC or DBC algorithm, and $\mu$: the case label in the PBC or DBC algorithm.
+4
+5Build an envy graph $G$ for two agents;
 6 when item $e\in T$ arrives do
-4       if  there is no envy between the two agents  then
-5             if $\forall k\in N$, $v_{k}(e)=a_{k}$ or $\forall k\in N$, $v_{k}(e)=b_{k}$  then
-6                   Arbitrarily choose one agent $i\in N$;
+7 if $\alpha=1$ then
+8
+9    ${\bf A},\alpha\leftarrow\mathsf{EGP}(N,{\bf A},\alpha,e)$;
+10    if $\alpha=2$ then
+       ${\bf A},\alpha\leftarrow\mathsf{PBC}(N,{\bf A},\alpha,e,(\lambda_{1},\mu_{1}))$ ;
+          // Envy cycle will appear.
+11
+12      end if
+13
+14 else if $\alpha=2$ then
+15    ${\bf A},\alpha,(\lambda_{1},\mu_{1})\leftarrow\mathsf{PBC}(N,{\bf A},\alpha,e,(\lambda_{1},\mu_{1}))$;
+16 else if $\alpha=3$ then
+17    ${\bf A},\alpha,(\lambda_{2},\mu_{2})\leftarrow\mathsf{DBC}(N,{\bf A},\alpha,e,(\lambda_{2},\mu_{2}))$;
+18   end if
+19
+20  Update the envy graph $G$;
+21
+return $\bf A$};
+Algorithm 3 Adapted Envy-Graph Procedure
+Theorem 4.
+For the deterministic allocation of indivisible goods for two agents with additive personalized bi-valued valuations, the AEGP algorithm (Algorithm 3) computes an NW allocation that satisfies $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS.
+The following two lemmas are crucial in the proof of Theorem 4.
+Input: $N,A,\alpha,e$
+Output: An NW allocation A and $\alpha$
+1
+2if there is no envy between the two agents  then
+3 if $\exists i\in N$, $v_{i}(e)=b_{i}$  then
+4      Choosing the agent $i$ who values $e$ at $b_{i}$ (breaking ties arbitrarily);
 5 else
-8                   Choosing the agent $i\in N$ who has the valuation of $b_{i}$;
+6      Arbitrarily choose one agent $i$;
 7   end if
-10
-11      else
-12             Let $i\in N$ be an unenvied agent and $j=2-i$;
+8
+9else
+10   Let $i\in N$ be an unenvied agent and j=2ij=2-i;
 11 if $v_{i}(A_{i}\cup\{e\})<v_{i}(A_{j})$ then
-14                   Terminate this algorithm and implement the Preliminary-Breaking-Cycle algorithm (Algorithm 4);
+12    $\alpha=2$;
+13    return A and $\alpha$;
+14   end if
 15
 16 end if
-17      $A_{i}=A_{i}\cup\{e\}$;
-18
-return $A$;
-Algorithm 3 Adapted Envy-Graph Procedures
-Theorem 5.
-For the deterministic allocation of indivisible goods for two agents with additive bi-valued valuations, the Adapted Envy-Graph Procedures (Algorithm 3) computes an NW allocation that satisfies $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS.
-Before we show Theorem 5, we first give the following key observations and lemmas.
+[…]
+return A and $\alpha$;
+Algorithm 4 EGP (Envy-Graph Procedure)
+Lemma 3.
+The EGP algorithm (Algorithm 4) computes an NW allocation that is EF1 without an envy cycle.
+Proof.
+Note that there are three kinds of input allocation.
+The first one is the initial empty allocation, which is EF1 without an envy cycle.
+The second one is the output allocation of the PBC algorithm.
+The last one is the output allocation of the DBC algorithm.
+It suffices to show that the last two kinds of input allocation are EF1 without an envy cycle.
+That is because if the input allocation of the EGP algorithm is EF1 without an envy cycle, by the description of the algorithm and induction, it is easy to check that the property of EF1 without an envy cycle is always maintained.
+By Lemma 4, we can derive that when the PBC or DBC algorithm outputs $\alpha=1$, implying that the EGP algorithm is ready to be executed, the output allocation of the PBC or DBC algorithm is EF1 without an envy cycle.
+Therefore, our claim holds.
+∎
+Lemma 4.
+The PBC or DBC algorithm (Algorithm 5 or 6) computes an NW allocation satisfying
+  • $\frac{1}{2}$-EF1,
+  • or EF1 with an envy cycle,
+  • or EF1 without an envy cycle.
+We split the proof of Lemma 4 into two parts, where we show that the claim holds for both the PBC and DBC algorithms.
 Observation 1.
-When Algorithm 4 is executed, for the first arriving item $e$, we have $v_{i}(e)=a_{i}$.
+When the PBC algorithm (Algorithm 5) is executed, for the first arriving item $e$, we have $v_{i}(e)=a_{i}$.
 Proof.
-The reason why Algorithm 4 is executed is that for the arriving item $e$, if we add it to agent $i$’s bundle, there will be an envy cycle between two agents.
+The reason why the PBC algorithm is executed is that for the arriving item $e$, if we add it to agent $i$’s bundle, there will be an envy cycle between two agents.
 i.e., $v_{i}(A_{i}\cup\{e\})<v_{i}(A_{j})$ and $v_{j}(A_{j})<v_{i}(A_{i}\cup\{e\})$.
@@
 Observation 2.
-When Algorithm 4 is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
+When the PBC algorithm (Algorithm 5) is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
 Proof.
@@
 After allocating $e$, there will be an envy cycle, i.e., $v_{i}(e)=a_{i}<v_{i}(A_{j})=b_{i}$ and $v_{j}(A_{j})<v_{j}(e)$.
-If the second inequality holds, we have $v_{j}(A_{j})=a_{j}$ but $v_{i}(A_{j})=b_{i}$, which contradicts the line 3 in Algorithm 3.
+If the second inequality holds, we have $v_{j}(A_{j})=a_{j}$ but $v_{i}(A_{j})=b_{i}$, which contradicts the line 4 in the EGP algorithm (Algorithm 4).
 Suppose that $|A_{j}|=0$.
@@
 Lemma 5.
-Algorithm 4 computes an NW allocation that is EF1 with or without an envy cycle, or $\frac{1}{2}$-EF1.
+The PBC algorithm (Algorithm 5) computes an NW allocation satisfying:
+  • $\frac{1}{2}$-EF1,
+  • or EF1 with an envy cycle,
+  • or EF1 without an envy cycle.
 Proof.
-Before Algorithm 4 is implemented, the output allocation $A=(A_{1},A_{2})$ by Algorithm 3 is EF1. According to the algorithm description, it is executed in at most two rounds.
-$\lambda=1$.
+When the PBC algorithm is implemented ($\alpha=2$), the input allocation $\textbf{A}=(A_{1},A_{2})$ output by the EGP algorithm (Algorithm 4) is EF1 without an envy cycle.
+By the description of the PBC algorithm, it is executed in at most two consecutive rounds.
+$\lambda_{1}=1$.
 For the first arriving item $e_{1}$, by Observation 1, we have $v_{i}(e_{1})=a_{i}$.
@@
 If $\lvert A_{j}\rvert>1$, let us consider two cases about the valuation of item $e_{1}$.
-  • If $v_{j}(e_{1})=b_{j}$, item $e_{1}$ is allocated to agent $j$ and we set $\mu=1$.
+  • If $v_{j}(e_{1})=b_{j}$, item $e$ is allocated to agent $j$.
+Then, we set $\lambda_{1}=2$ and $\mu_{1}=1$.
 In this case, for agent $j$, her new bundle is EF since $v_{j}(A_{j}\cup\{e_{1}\})>v_{j}(A_{j})\geq v_{j}(A_{i})$ and for agent $i$, her bundle is $\frac{1}{2}$-EF1, i.e, $v_{i}(A_{i})\geq\frac{1}{2}v_{i}(A_{j}\cup\{e_{1}\}\setminus\{g\})$ holds for $g\in A_{j}\cup\{e_{1}\}$, since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(A_{i})\geq a_{i}=v_{i}(e_{1})$.
 Thus, this allocation is $\frac{1}{2}$-EF1.
-  • If $v_{j}(e_{1})=a_{j}$, item $e_{1}$ is allocated to agent $i$ and we set $\mu=2$.
+  • If $v_{j}(e_{1})=a_{j}$, item $e$ is allocated to agent $i$.
+Then, we set $\lambda_{1}=2$ and $\mu_{1}=2$.
 In this case, for agent $i$, her new bundle is still EF1 since she is an unenvied agent in Algorithm 3 and the input allocation $A$ is EF1, and for agent $j$, her bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ but $v_{j}(A_{j})<v_{j}(A_{j}\cup\{e_{1}\})$.
 Thus, this allocation is EF1 with an envy cycle.
-$\lambda=2$. The second item $e_{2}$ arrives. Let us consider two cases about $\mu$.
-  • If $\mu=1$, agent $i$ picks it.
+It is clear that any possible allocation from the above cases does not satisfy EF1 without an envy cycle. Thus, we keep $\alpha=2$.
+$\lambda_{1}=2$. The second item $e_{2}$ arrives. Let us consider two cases about $\mu_{1}$.
+  • If $\mu_{1}=1$, agent $i$ picks it.
 In this case, for agent $j$, her bundle is still EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{1})=b_{j}\geq v_{j}(e_{2})$.
 For agent $i$, her new bundle is EF1 since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(e_{2})\geq v_{i}(e_{1})=a_{i}$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
-  • If $\mu=2$, let us consider two cases.
+Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$.
+  • If $\mu_{1}=2$, let us consider two cases.
 When $v_{i}(e_{2})=a_{i}$, agent $j$ picks it.
@@
 For agent $j$, her new bundle is EF because $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{2})=v_{j}(e_{1})=a_{j}$.
-Hence, the allocation is EF1 without an envy cycle, and the algorithm is terminated.
+Hence, the allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$.
 When $v_{i}(e_{2})=b_{i}$, it is allocated to agent $i$.
@@
 For agent $j$, her bundle is $\frac{1}{2}$-EF1, i.e., $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\cup\{e_{1},e_{2}\}\setminus\{g\})$ holds for some $e\in A_{i}\cup\{e_{1},e_{2}\}$, since $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(A_{j})\geq v_{j}(e_{1})=\min\{v_{j}(e_{1}),v_{j}(e_{2})\}$.
-Thus, this allocation is $\frac{1}{2}$-EF1 and the algorithm is terminated.
+Thus, this allocation is $\frac{1}{2}$-EF1, and we set $\alpha=3$, $\lambda_{1}=1$, and $\mu_{1}=0$.
 Combining the above cases, it is not hard to see that our claim is correct. ∎
 Observation 3.
-When Algorithm 5 is executed, there are two properties about input allocation $A$:
+When the DBC algorithm (Algorithm 6) is executed, there are two properties about input allocation $\bf A$:
   • $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$;
-  • Let $e_{1}$ and $e_{2}$ denote two new allocated items in Algorithm 4, we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$, where $e\in\arg\max_{g\in\{e_{1},e_{2}\}}v_{j}(g)$.
+  • Let $e_{1}$ and $e_{2}$ denote two new allocated items in the PBC algorithm, we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$, where $e\in\arg\max_{g\in\{e_{1},e_{2}\}}v_{j}(g)$.
 Proof.
-When Algorithm 5 is ready to be executed, Algorithm 4 is terminated for $\lambda=\mu=2$. Then, by the proof of Algorithm 4, it is not hard to see that the statements are correct. ∎
+When the DBC algorithm is ready to be executed, the PBC algorithm is terminated for $\lambda_{1}=\mu_{1}=2$. Then, by the proof of Lemma 5, it is easy to see that the statements are correct. ∎
 Lemma 6.
-Algorithm 5 computes an NW allocation that is EF1 with or without an envy cycle, or $\frac{1}{2}$-EF1.
+The DBC algorithm (Algorithm 6) computes an NW allocation satisfying
+  • $\frac{1}{2}$-EF1,
+  • or EF1 with an envy cycle,
+  • or EF1 without an envy cycle.
 Proof.
-By Observation 3, the input allocation $A=(A_{1},A_{2})$ is $\frac{1}{2}$-EF1, where $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$ for some $e\in A_{i}$.
-Unlike Algorithm 4, this algorithm may be implemented in many rounds.
-Let $g_{1}$ and $g_{2}$ denote the two online items that are allocated to agent $i$ in Algorithm 4 and we have $\min\{v_{j}(g_{1}),v_{j}(g_{2})\}=a_{j}$.
-$\lambda=1$. For the arriving item $e_{1}$ when $\lambda=1$, let us consider four cases of agents’ valuations.
-  • When $v_{i}(e_{1})=a_{i}$ and $v_{j}(e_{1})=b_{j}$, agent $j$ picks it and we set $\mu=1$.
-For agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$.
+By Observation 3, the input allocation $\textbf{A}=(A_{1},A_{2})$ is $\frac{1}{2}$-EF1, where $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$ for some $e\in A_{i}$.
+Unlike the PBC algorithm, this algorithm may be executed in many rounds.
+Let $g_{1}$ and $g_{2}$ denote the two online items that are allocated to agent $i$ in the PBC algorithm and we have $\min\{v_{j}(g_{1}),v_{j}(g_{2})\}=a_{j}$.
+Currently, we have $\alpha=3$.
+$\lambda_{2}=1$. For the arriving item $e_{1}$ when $\lambda_{2}=1$, let us consider three cases of agents’ valuations.
+  • When $v_{i}(e_{1})=a_{i}$, agent $j$ picks it.
+There are two subcases about $v_{j}(e_{1})$.
+If $v_{j}(e_{1})=b_{j}$, for agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$.
 For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$.
 Thus, this allocation is EF1 without an envy cycle.
-  • When $v_{i}(e_{1})=b_{i}$ and $v_{j}(e_{1})=b_{j}$, it is allocated to agent $j$ and we set $\mu=2$.
+If $v_{j}(e_{1})=a_{j}$, for agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$.
+For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=a_{i}=\min\{v_{j}(g_{1}),v_{j}(g_{2})\}$.
+Thus, this allocation is EF1 without an envy cycle.
+Combining the above two cases, we set $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=1$, implying that the EGP algorithm is ready to be executed.
+  • When $v_{i}(e_{1})=b_{i}$ and $v_{j}(e_{1})=b_{j}$, it is allocated to agent $j$ and we set $\mu_{2}=1$.
 For agent $i$, her bundle is EF1 since after removing $e_{1}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$.
 For agent $j$, her new bundle is also EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$.
-Thus, this allocation is EF1 with an envy cycle.
-  • When $v_{i}(e_{1})=a_{i}$ and $v_{j}(e_{1})=a_{j}$, it is allocated to agent $j$ and we set $\mu=3$.
-For agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$.
-For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=a_{i}=\min\{v_{j}(g_{1}),v_{j}(g_{2})\}$.
-Thus, this allocation is EF1 without an ency cycle.
+Thus, this allocation is EF1 with an envy cycle, and we set $\lambda_{2}=2$.
 […]
-$\lambda=2$. For the arriving item $e_{2}$ when $\lambda=2$, let us consider four cases about $\mu$.
-  • When $\mu=1$, item $e_{2}$ is directly allocated to agent $j$.
-For agent $i$, her bundle is EF1 since after removing $e_{2}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j}\cup\{e_{1}\})=v_{i}(A_{j})+a_{i}$.
-For agent $j$, her new bundle is EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{1})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{2})\geq v_{j}(g_{1})=a_{j}$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
-  • When $\mu=2$, we need to consider two subcases.
+$\lambda_{2}=2$. For the arriving item $e_{2}$ when $\lambda_{2}=2$, let us consider two cases about $\mu_{2}$.
+  • When $\mu_{2}=1$, we need to consider two subcases.
 If $v_{i}(e_{2})=b_{i}$, agent $i$ picks it.
@@
 Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3.
-Additionally, we set $\lambda=1$ and continue executing this algorithm.
+Additionally, we set $\lambda_{2}=1$, $\mu_{2}=0$, and keep $\alpha=3$, implying that the DBC algorithm continues to be executed.
 If $v_{i}(e_{2})=a_{i}$, it is allocated to agent $j$.
@@
 For agent $j$, her new bundle is trivially EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{1})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{2})\geq a_{i}=v_{j}(g_{1})$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
-  • When $\mu=3$, let us also consider two subcases.
-If $v_{j}(e_{2})=b_{j}$, agent $j$ picks it.
-For agent $i$, her bundle is EF1 since after removing $e_{2}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j}\cup\{e_{1}\})=v_{i}(A_{j})+a_{i}$.
-For agent $j$, her new bundle is EF since $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$, $v_{j}(e_{2})=b_{j}\geq v_{j}(g_{2})$ and $v_{j}(e_{1})=a_{j}=v_{j}(e_{2})$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
-If $v_{j}(e_{2})=a_{j}$, agent $i$ picks it.
-For agent $i$, her new bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{2})=a_{i}=v_{i}(e_{1})$.
-For agent $j$, her bundle is $\frac{1}{2}$-EF1 since we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{g\})$ for some $g\in A_{i}$ and $v_{j}(e_{1})=a_{j}=v_{j}(e_{2})$.
-Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3.
-Additionally, we set $\lambda=1$ and continue executing this algorithm.
-  • When $\mu=4$, it is directly allocated to agent $j$.
+Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$, implying the EGP algorithm is ready to be executed.
+  • When $\mu=2$, it is directly allocated to agent $j$.
 For agent $i$, her bundle is EF since $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{i}(e_{1})=b_{i}\geq v_{i}(e_{2})$.
@@
 Thus, this allocation is $\frac{1}{2}$-EF1 and satisfies the properties in Observation 3.
-Additionally, we set $\lambda=1$ and continue executing this algorithm.
+Additionally, we set $\lambda_{2}=1$, $\mu_{2}=0$, and keep $\alpha=3$, implying that the DBC algorithm continues to be executed.
 Combining the above cases, our claim holds. ∎
-Proof of Theorem 5.
-In the main body of Algorithm 3, we follow a similar way of offline envy graph procedures to allocate items.
-The precondition of Algorithm 3 is that there is no envy cycle between two agents.
-Initially, the empty allocation is EF1 without an envy cycle.
-When there is no envy between two agents, we allocate the arriving item to one agent according to different conditions.
-When there exists an envy between two agents, we have two cases.
-One is that if we allocate it to the unenvied agent, the envy cycle does not appear.
-We will follow this procedure.
-The other is that if we allocate it to the unenvied agent, the envy cycle will appear, and we terminate Algorithm 3.
-In the online setting, we cannot swap bundles to eliminate the envy cycle.
-Thus, to enable the Algorithm 3 to be executed normally, Algorithms 4 and 5 help us to manually break the possible envy cycle, where, by Lemmas 1 and 2, we guarantee that $\frac{1}{2}$-EF1 always holds.
-When the allocation is EF1 and there is no envy cycle, we continue executing Algorithm 3.
-By the induction, we can conclude that the output allocation of Algorithm 3 is always EF1 without an envy cycle.
-Combining Lemmas 1, 2, and the description of Algorithm 3, it can be seen that we can guarantee an NW allocation that is $\frac{1}{2}$-EF1.
-As for MMS, by the implication between EF1 and MMS [21], i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{\alpha}{(n-1)\alpha+1}$-MMS, where $n$ is the number of agents, it is clear that the output allocation is $\frac{\frac{1}{2}}{\frac{1}{2}+1}=\frac{1}{3}$-MMS.
-∎
-Input: An instance $(T,N,(v_{1},v_{2}))$ with additive bi-valued valuation functions
-Output: An non-wasteful approximate EF1 and approximate MMS allocation $A$
+Input: $N,A,\alpha,e,(\lambda_{1},\mu_{1})$
+Output: An NW allocation A, $\alpha$, and $(\lambda_{1},\mu_{1})$
 1
-2Initialize: $\lambda=1$ and $\mu=0$;
-3 when item $e\in T$ arrives do
-4
-5      if $\lambda=1$ then
-6
+2if $\lambda_{1}=1$ then
+3
 4 if $\lvert A_{j}\rvert=1$ then
+5
+6    $A_{j}=A_{j}\cup\{e\}$;
+7    $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$;
+8 else
+9
+10    if $v_{j}(e)=b_{j}$ then
+11
+12       $A_{j}=A_{j}\cup\{e\}$;
+13       $\lambda_{1}=2$ and $\mu_{1}=1$;
+14    else
+15       $A_{i}=A_{i}\cup\{e\}$;
+16       $\lambda_{1}=2$ and $\mu_{1}=2$;
+17      end if
+18
+19   end if
+20
+21else
+22
+23 if $\mu_{1}=1$  then
+24
+25    $A_{i}=A_{i}\cup\{e\}$;
+26 else
+27
+28    if $v_{i}(e)=a_{i}$ then
+29
+30       $A_{j}=A_{j}\cup\{e\}$;
+31    else
+32
+33       $A_{i}=A_{i}\cup\{e\}$;
+34       $\alpha=3$, $\lambda_{1}=1$, and $\mu_{1}=0$;
+35      end if
+36
+37   end if
+38
+39 $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$;
+40 end if
+41
+return A, $\alpha$ and $(\lambda_{1},\mu_{1})$;
+Algorithm 5 PBC (Preliminary-Breaking-Cycle)
+Input: $N,A,\alpha,e,(\lambda_{2},\mu_{2})$
+Output: An NW allocation A, $\alpha$, and $(\lambda_{2},\mu_{2})$
+1
+2if $\lambda_{2}\mod 2=1$ then
+3
+4 if $v_{i}(e)=a_{i}$ then
+5    $A_{j}=A_{j}\cup\{e\}$;
+6    $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$;
+7 else if $v_{i}(e)=b_{i}$ and $v_{j}(e)=b_{j}$ then
 8
 9    $A_{j}=A_{j}\cup\{e\}$;
-10                  Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
+10    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=1$;
 11 else
-12
-13                  if $v_{j}(e)=b_{j}$ then
-14
-15                        $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
+12    $A_{i}=A_{i}\cup\{e\}$;
+13    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=2$;
+14   end if
+15
+16 else if $\lambda_{2}\mod 2=0$ then
+17 if  $\mu_{2}=1$  then
+18    if $v_{i}(e)=b_{i}$ then
+19
+20       $A_{i}=A_{i}\cup\{e\}$;
+21       $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
 22    else
-17                         $A_{i}=A_{i}\cup\{e\}$ and $\mu=2$;
+23       $A_{j}=A_{j}\cup\{e\}$;
+24       $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$;
 25      end if
-19                  $\lambda=2$;
-20             end if
-21
-22      else
-23
-24            if $\mu=1$  then
-25
-26                  $A_{i}=A_{i}\cup\{e\}$;
+26
 27 else
 28
-29                  if $v_{i}(e)=a_{i}$ then
-30
 29    $A_{j}=A_{j}\cup\{e\}$;
-32                  else
-33
-34                        $A_{i}=A_{i}\cup\{e\}$;
-35                        Terminate this algorithm and execute the deep-breaking-cycle algorithm (Algorithm 5);
+30    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
 31   end if
-37
+32
 33 end if
-39
-40            Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
-41       end if
-42
-43
-return $A$;
-Algorithm 4 Preliminary-Breaking-Cycle Algorithm
-Input: An instance $(T,N,(v_{1},v_{2}))$ with additive bi-valued valuation functions
-Output: An non-wasteful approximate EF1 and approximate MMS allocation $A$
-1 Initialize: $\lambda=1$ and $\mu=0$;
-2 when item $e\in T$ arrives do
-3
-4      if $\lambda=1$ then
-5
-6            if $v_{i}(e)=a_{i}$ and $v_{j}(e)=b_{j}$ then
-7                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
-8             else if $v_{i}(e)=b_{i}$ and $v_{j}(e)=b_{j}$ then
-9
-10                  $A_{j}=A_{j}\cup\{e\}$ and $\mu=2$;
-11             else if  $v_{i}(e)=a_{i}$ and $v_{j}(e)=a_{j}$  then
-12                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=3$;
-13             else
-14                   $A_{i}=A_{i}\cup\{e\}$ and $\mu=4$;
-15             end if
-16            $\lambda=2$;
-17      else
-18             if $\mu=1$ then
-19                   $A_{j}=A_{j}\cup\{e\}$;
-20             else if  $\mu=2$  then
-21                   if $v_{i}(e)=b_{i}$ then
-22
-23                        $A_{i}=A_{i}\cup\{e\}$;
-24                        $\lambda=1$;
-25                  else
-26                         $A_{j}=A_{j}\cup\{e\}$;
-27                   end if
-28
-29             else if  $\mu=3$  then
-30                   if $v_{j}(e)=b_{j}$ then
-31
-32                        $A_{j}=A_{j}\cup\{e\}$;
-33                  else
-34
-35                        $A_{i}=A_{i}\cup\{e\}$;
-36                        $\lambda=1$;
-37                   end if
-38
-39             else
-40
-41                  $A_{j}=A_{j}\cup\{e\}$;
-42                  $\lambda=1$;
-43             end if
-44
-45            if $\lambda=2$ then
-46                   Terminate this algorithm and execute the adapted envy-graph procedures (Algorithm 3);
-47
-48       end if
-49
-50
-51
-52return $A$;
-Algorithm 5 Deep-Breaking-Cycle Algorithm
-Then, we present the impossible results to show that our approximation guarantees for EF1 and MMS are tight, and there is no compatibility between EF1 or MMS, and USW.
+34return A, $\alpha$, $(\lambda_{2},\mu_{2})$;
+Algorithm 6 DBC (Deep-Breaking-Cycle)
+Proof of Theorem 4.
+In the AEGP, it can be seen that for different values of the algorithm indicator $\alpha$, we need to choose the corresponding algorithm, i.e., $\alpha=1$: EGP, $\alpha=2$: PBC, or $\alpha=3$: DBC, to execute.
+By Lemmas 3 and 4, it can be seen that the output allocation is NW and $\frac{1}{2}$-EF1.
+Regarding MMS, by the implication between EF1 and MMS Amanatidis et al.
+(2018), i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{\alpha}{(n-1)\alpha+1}$-MMS, where $n$ is the number of agents, it is clear that the output allocation is $\frac{\frac{1}{2}}{\frac{1}{2}+1}=\frac{1}{3}$-MMS.∎
+For an independent interest, in the Appendix, we show that if each online item has a deadline of one period, i.e., each arriving item can wait for at most one round to be allocated, EF1 allocation can be found in the above setting.
+At last, we present the impossible results to show that our approximation guarantees for EF1 and MMS are tight, and the approximate EF1 (or MMS) is incompatible with USW.
 Theorem 5.
-For two agents with additive bi-valued valuation functions, no deterministic online algorithm can compute an NW allocation that is $(\frac{1}{2}+\epsilon)$-EF1 or $(\frac{1}{3}+\epsilon)$-MMS for any $\epsilon>0$.
+For two agents with additive personalized bi-valued valuation functions, no deterministic online algorithm can compute an NW allocation that is $(\frac{1}{2}+\epsilon)$-EF1 or $(\frac{1}{3}+\epsilon)$-MMS for any $\epsilon>0$.
 Proof.
-$t=1$ $t=2$ $t=3$ $t=4$
-agent $1$ $\epsilon$ $1$ $\epsilon$ 1
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
+agent 11 $\epsilon$ 11 $\epsilon$ 1
 agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1
 Case 1
-$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
-agent $1$ $\epsilon$ $1$ $\epsilon$ $\epsilon$ $1$
-agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1 $1$
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
+agent 11 $\epsilon$ 11 $\epsilon$ $\epsilon$ 11
+agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1 11
 Case 2-1
-$t=1$ $t=2$ $t=3$ $t=4$
-agent $1$ $\epsilon$ $1$ $\epsilon$ $\epsilon$
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
+agent 11 $\epsilon$ 11 $\epsilon$ $\epsilon$
 agent 2 $\epsilon$ $\epsilon$ $\epsilon$ 1
 Case 2-2
-Table 4: The impossibility results of EF1 and MMS for additive bi-valued valuations for two agents.
-Consider an online allocation instance with two agents $\{1,2\}$.
-For each agent $i\in\{1,2\}$, $v_{i}(e)\in\{\epsilon,1\}$, where $10<\epsilon<$ 1.
-The value of each arrived item is shown in Table 4.
+Table 6: The impossibility results of EF1 and MMS for additive personalized bi-valued valuations for two agents.
+Let $a_{1}=a_{2}=\epsilon$ and $b_{1}=b_{2}=1$, where $0<\epsilon<1$.
+Consider the following online instance for $n=2$, where the value of each item is shown in Table 6.
 When $t=1$, the item with the value of $\epsilon$ for both agents arrives.
-Without loss of generality, we allocate $e_{1}$ to agent 1.
+Without loss of generality, agent 1 picks it.
 When $t=2$, the arriving item has the value of 1 for agent 1 and $\epsilon$ for agent 2.
-To avoid 0-EF1 and 0-MMS, it should be allocated to agent 2.
+If it is allocated to agent 1, agent 2’s bundle is empty, resulting in 0-EF1 and 0-MMS.
+Thus, it should be allocated to agent 2.
 Then, when $t=3$, the arriving item has the value of $\epsilon$ for both agents.
@@
 Proposition 2.
-For the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS, and $\beta$-max-USW for any $\alpha,\beta>0$, even for two agents with additive bi-valued valuation functions.
+Given any $\alpha,\beta>0$, for the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation satisfying
+  • $\alpha$-EF1 and $\beta$-max-USW,
+  • or $\alpha$-MMS and $\beta$-max-USW,
+even for two agents with additive personalized bi-valued valuations.
 Proof.
-Consider an online allocation instance with two agents $\{1,2\}$.
-Assume that, for any item $e$, $v_{1}(e)\in\{1,\frac{1}{\epsilon}\}$ and $v_{2}(e)\in\{\epsilon^{2},\epsilon\}$, where $10<\epsilon<$ 1.
-When $t=1$, the first item $e_{1}$ with the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives.
+Without loss of generality, we assume that $v_{i}\in\{1,\frac{1}{\epsilon}\}$ and $v_{2}\in\{\epsilon^{2},\epsilon\}$, where $0<\epsilon<1$.
+When $t=1$, the first item $e_{1}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives.
 If it is allocated to agent 2, no further items arrive.
-Hence, this allocation is EF1, MMS, and $\epsilon$-max-USW.
+Hence, this allocation is clearly EF1, MMS, and $\epsilon$-max-USW.
 If not, agent $1$ picks it, and this allocation is EF1, MMS, and max-USW.
@@
 If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\epsilon^{2}+1}{1+\frac{1}{\epsilon}}=\frac{\epsilon^{3}+\epsilon}{1+\epsilon}$-max-USW.
-For any $\beta>0$, let $\epsilon<\beta^{2}$, we can derive that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1/MMS and $\beta$-max-USW for any $\alpha,\beta>0$
+For any $\beta>0$, let $\epsilon<\beta^{2}$, we can derive that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1 (or $\alpha$-MMS) and $\beta$-max-USW for any $\alpha,\beta>0$
 3.3.2 The Combination with Additive Binary Valuations
-In this part, we consider the combination of additive bi-valued and additive binary valuation functions and study the setting where only one agent (agent $n$) has the additive bi-valued valuation function, and the remaining agents have identical additive binary valuation functions.
+In this part, we focus on the setting where only one agent (agent $n$) has an additive personalized bi-valued valuation function, while all other agents share identical additive binary valuation functions, formed by $(T,N,(v_{i})_{i\in N\setminus\{n\}},v_{n})$.
 We show that an NW allocation that is EF1 and MMS always exists and can be computed by the Adapted-Picking Algorithm (Algorithm 7).
-The high level idea of our algorithm is that we allocate the appearing items that all agents value at non-zero values evenly with respect to the requirement that agent $n$ has the top priority of picking the item with the value of $b_{n}$ and the lowest priority of picking the item with the value of $a_{n}$.
-In that case, agent $n$ takes advantage of the option of items, and then we can guarantee that EF1 and MMS hold simultaneously.
-Theorem 7.
-For the deterministic allocation of indivisible goods for $(n-1)$ agents with identical additive binary valuation functions and one agent with the additive bi-valued valuation function, the Adapted-Picking Algorithm (Algorithm 6) computes an NW allocation that satisfies EF1 and MMS.
-Input: An instance $(T,N,(v_{i})_{i\in N})$ with identical $(n-1)$ additive binary valuation functions $(v_{1},\ldots,v_{n})$ and one additive bi-valued valuation function $(v_{n})$
-Output: An non-wasteful EF1 and MMS allocation $A$
+The high-level idea of our algorithm is to allocate the arriving items that all agents value at non-zero values “evenly”, according to the requirement that agent $n$ has the top priority of picking the item with the value of $b_{n}$ and the lowest priority of picking the item with the value of $a_{n}$.
+In that case, agent $n$ takes advantage of the item options, and then we can guarantee that EF1 and MMS hold simultaneously.
+In Table 7, we illustrate how our algorithm works for two agents.
+Input: An instance $(T,N,(v_{i})_{i\in N\setminus\{n\}},v_{n})$ with identical (n1)(n-1) additive binary valuation functions $(v_{1},\ldots,v_{n})$ and one additive personalized bi-valued valuation function $v_{n}$
+Output: An NW, EF1, and MMS allocation $\bf A$}
 1
-2Let $A=(\emptyset,\ldots,\emptyset)$;
+2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
 3 when item $e\in T$ arrives do
@@
 12
-13                  if |$|\{g~{}|~{}g\in A_{n}~{}and~{}v_{i_{min}}(g)>0\}|\leq\min_{i\in[n-1]}|A_{i}|$ then
+13       if $v_{i_{min}}(A_{n})\leq v_{i_{min}}(A_{i_{min}})$ then
 14          $A_{n}=A_{n}\cup\{e\}$;
@@
 20    else
-21                   if |$|\{g~{}|~{}g\in A_{n}~{}and~{}v_{i_{min}}(g)>0\}|\geq\min_{i\in[n-1]}|A_{i}|$ then
+21       if $v_{i_{min}}(A_{n})>v_{i_{min}}(A_{i_{min}})$ then
 22
@@
 32
-return $A$;
+return $\bf A$};
 Algorithm 7 Adapted-Picking Algorithm
-$t=1$ $t=2$ $t=3$ $t=4$ $t=5$ldots
-agent $1$ $1$ $1$ 00 1 1 $\ldots$
-agent 2 $a_{2}$ $a_{2}$ $a_{2}$ $b_{2}$ $a_{2}$ldots
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $\ldots$
+agent 11 11 11 0 1 1 $\ldots$
+agent 2 $a_{2}$ $a_{2}$ $a_{2}$ $b_{2}$ $a_{2}$ $\ldots$
 Table 7: A simple illustration for two agents of Algorithm 7.
+Theorem 6.
+For the deterministic allocation of indivisible goods for (n1)(n-1) agents with identical additive binary valuation functions and one agent with the additive personalized bi-valued valuation function, the Adapted-Picking Algorithm (Algorithm 7) computes an NW, EF1, and MMS allocation.
 Proof.
-Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$.
-Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 6.
-Firstly, we show that the allocation in any round is EF1 by induction.
+Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$, the allocation at the end of round $t$ in Algorithm 7.
+EF1.
+We show that the allocation in any round is EF1 by induction.
 For the base case, the empty allocation is trivially EF1.
-For the induction step, after allocating item $e_{k}$, the allocation $A^{k}$ is EF1.
-Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k}$ is still EF1.
+For the induction step, after allocating item $e_{k}$, the allocation $\textbf{A}^{k}$ is EF1.
+Next, we show that after allocating item $e_{k+1}$, the allocation $\textbf{A}^{k}$ is still EF1.
 When item $e_{k+1}$ arrives, let us consider several cases about the allocation of this item.
@@
 For agent $n$, the value of her bundle increases, so her bundle is also EF1 to her.
-  • If $v_{n}(e_{k+1})=b_{n}$, let us consider two subcases.
-When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|\leq\min_{i\in[n-1]}|A_{i}^{k}|$, agent $n$ picks it.
-For agent $n$, her new bundle is EF.
-That is because she has the priority of picking the item with the value of $b_{n}$ and before allocating item $e_{k+1}$, if she envies other agents $j\in[n-1]$, in her perspective, the gap between the value of agent $j$’s bundle and her bundle is at most $a_{n}$, i.e.,$v_{n}(A_{j})-v_{n}(A_{n})\leq a_{n}$.
-Now, fix an arbitrary agent $i\in[n-1]$.
-Since agent $i$ has an identical binary valuation function and she only picks the item with the non-zero value, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{n}^{k})=v_{i}(A_{n}^{k+1}\setminus\{e_{k+1}\})$, implying her bundle is also EF1.
-When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|>\min_{i\in[n-1]}|A$ , $i_{min}$ picks it.
-Clearly, for agent $i_{min}$, her new bundle is still EF1.
-For any other agent $i\in N\setminus\{i_{min},n\}$, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{i_{min}}^{k})=v_{i}(A_{i_{min}}^{k+1}\setminus\{e_{k+1}\})$, implying that her bundle is still EF1.
-For agent $n$, based on the description of our algorithm, she must already pick one item that has the value of $b_{n}$, and in her perspective, the number of items in her bundle is at least one more than that of agent $i_{min}$’s bundle, which means that agent $n$ does not envy agent $i_{min}$ and in her perspective, the value of her bundle is at least $b_{n}$ greater than agent $i_{min}$’s bundle.
-Therefore, allocating $e_{k+1}$ to agent $i_{min}$ will not make agent $n$ envy agent $i_{min}$.
-  • If $v_{n}(e_{k+1})=a_{n}$, let us consider two subcases.
-When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|\geq\min_{i\in[n-1]}|A_{i}^{k}|$, agent $i_{min}$ picks it.
-It is trivial that for agent $i_{min}$, her new bundle is EF1.
-For any other agent $i\in N\setminus\{i_{min},n\}$, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{i_{min}}^{k})=v_{i}(A_{i_{min}}^{k+1}\setminus\{e_{k+1}\})$, implying that her bundle is still EF1.
-For agent $n$, before allocating $e_{k+1}$, she does not envy agent $i_{min}$ based on the fact that the number of items in her bundle is at least the same as that of agent $i_{min}$’s bundle and our algorithm description.
-When |$|\{g~{}|~{}g\in A_{n}^{k}~{}and~{}v_{i_{min}}(g)>0\}|<\min_{i\in[n-1]}|A_{i}^{k}|$, agent $n$ picks it.
-For agent $n$, her new bundle is actually EF since she has the priority of picking the item with the value of $b_{n}$ and before allocating $e_{k+1}$, she envies other agent $j\in[n-1]$; in her perspective, the gap between the value of agent $j$’s bundle and her bundle is at most $a_{n}$, i.e.,$v_{n}(A_{j})-v_{n}(A_{n})\leq a_{n}$.
-Fix an arbitrary agent $i\in[n-1]$.
-Since agent $i$ has the same binary valuation function and she only picks the item with the non-zero value, we have $v_{i}(A_{i}^{k+1})=v_{i}(A_{i}^{k})\geq v_{i}(A_{n}^{k})=v_{i}(A_{n}^{k+1}\setminus\{e_{k+1}\})$, implying EF1 also holds.
+[…]
+[…]
 Combining the above cases, it can be concluded that in round $k+1$, the allocation is EF1. This completes the induction and establishes the correctness of our claim.
-Secondly, we will show that $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ is MMS.
+MMS.
+Fix an arbitrary round $k$.
 Pick an arbitrary agent $i\in[n-1]$.
 For agent $i$, assume that the number of items with the value of one that arrive between round $t=1$ and $k$ is $s_{i}^{k}$.
-By the description of our algorithm, items are almost evenly allocated to agents with identical binary valuations.
+By the description of our algorithm, items are almost evenly allocated to agents with identical additive binary valuations.
 Thus, it is trivial that $v_{i}(A_{i})=|v_{i}(A_{i})|\geq\mathsf{MMS}_{i}=\lfloor\frac{s_{i}^{k}}{n}\rfloor$.
@@
 ∎
-Then, we complement the above result by constructing the following example to show the incompatibility of EF1 or MMS, and USW in the combination of additive binary and additive bi-valued valuation functions.
+At last, we complement the above result by showing the incompatibility of EF1 (or MMS) and USW.
 Proposition 3.
-For the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation that is EF1 or MMS, and $\alpha$-max-USW for any $\alpha>0$, even for two agents where one agent has the additive binary valuation function, and the other agent has the additive bi-valued valuation function.
+Given any $\alpha>0$, for the deterministic allocation of indivisible goods, no online algorithm can compute an NW allocation satisfying
+  • EF1 and $\alpha$-max-USW,
+  • or MMS and $\alpha$-max-USW,
+even for two agents, where one agent has the additive binary valuation function, and the other agent has the additive personalized bi-valued valuation function.
 Proof.
-Consider an online allocation instance with two agents $\{1,2\}$.
-Assume that, for any item $e$, $v_{1}(e)\in\{0,1\}$ and $v_{2}(e)\in\{\epsilon^{2},\epsilon\}$, where $10<\epsilon<$ 1.
+Without loss of generality, we assume that $v_{1}\in\{0,1\}$ and $v_{2}\in\{\epsilon^{2},\epsilon\}$, where $0<\epsilon<1$.
 When $t=1$, the first item $e_{1}$ that has the value of 1 for agent 1 and $\epsilon$ for agent 2 arrives.
@@
 In other words, no allocation is EF1 or MMS.
-Combining the above cases, it can be seen that if an online algorithm outputs an NW allocation that is EF1 or MMS in the setting of binary and bi-valued valuations, it may not guarantee $\alpha$-max-USW for any given $\alpha>0$.
+Combining the above cases, it can be seen that if an online algorithm outputs an NW allocation that is EF1 or MMS in the setting of additive binary and additive personalized bi-valued valuations, it may not guarantee $\alpha$-max-USW for any given $\alpha>0$.
 ∎
@@
 In other words, for each arriving chore, there must be some agent taking it.
-Therefore, we first introduce the following constraint, and then give the fairness and efficiency definitions in the chores setting.
-Definition 7 (Completeness for Chores).
-An allocation $A^{k}$ is complete if, for any chore $e$ that has arrived by round $k$, it is not thrown away.
+Therefore, we first introduce the following constraint.
+Definition 7 (Completeness).
+An allocation ${\bf A}^{k}$ is complete if, for any chore $e$ that has arrived by round $k$, it is not discarded.
+Here are the fairness and efficiency definitions used in the chores setting.
 Definition 8 (Envy-freeness up to One Chore).
-For any $\alpha\in[1,+\infty)$, an allocation $A^{k}$ is $\alpha$-approximate envy-free up to one chore ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i}^{k}\neq\emptyset$, it holds that $c_{i}(A_{i}^{k}\setminus\{e\})\leq\alpha\cdot c_{i}(A_{i^{\prime}}^{k})$ for some $e\in A_{i}^{k}$.
+For any $\alpha\in[1,+\infty)$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate envy-free up to one chore ($\alpha$-EF1) if, for every pair of agents $i,i^{\prime}\in N$ with $A_{i}^{k}\neq\emptyset$, it holds that $c_{i}(A_{i}^{k}\setminus\{e\})\leq\alpha\cdot c_{i}(A_{i^{\prime}}^{k})$ for some $e\in A_{i}^{k}$.
 Definition 9 (Minimax Share Fairness).
 For any agent $i$, her minimax share $\mathsf{MMS}_{i}^{k}$ by round $k$ is defined as:
-$\mathsf{MMS}_{i}^{k}=\min_{B^{k}\in\Pi_{n}(T^{k})}\max_{i^{\prime}\in N}c_{i}(B_{i^{\prime}}^{k}),$
+$\mathsf{MMS}_{i}^{k}=\min_{{\bf B}^{k}\in\Pi_{n}(T^{k})}\max_{i^{\prime}\in N}c_{i}(B_{i^{\prime}}^{k}),$
 where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
-For any $\alpha\in[1,+\infty)$, an allocation $A^{k}$ is $\alpha$-approximate minimiax share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $c_{i}(A_{i}^{k})\leq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
+For any $\alpha\in[1,+\infty)$, an allocation ${\bf A}^{k}$ is $\alpha$-approximate minimiax share fair ($\alpha$-MMS) if for any agent $i\in N$, it holds that $c_{i}(A_{i}^{k})\leq\alpha\cdot\mathsf{MMS}_{i}^{k}$.
 Definition 10 (Utilitarian Social Cost).
-The utilitarian social cost $USC$ of allocation $A^{k}$ is given by $USC(A^{k})=\sum_{i\in N}c_{i}(A_{i}^{k})$.
-For any $\alpha\in[1,+\infty)$, an allocation $A^{k}$ is $\alpha$-min-USC if it holds that $USC(A^{k})\leq\alpha\cdot\min_{B^{k}\in\Pi_{n}(T^{k})}USC(B^{k})$, where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
+The utilitarian social cost $USC$ of allocation ${\bf A}^{k}$ is given by $USC({\bf A}^{k})=\sum_{i\in N}c_{i}(A_{i}^{k})$.
+For any $\alpha\in[1,+\infty)$, an allocation ${\bf A}^{k}$ is $\alpha$-min-USC if it holds that $USC({\bf A}^{k})\leq\alpha\cdot\min_{{\bf B}^{k}\in\Pi_{n}(T^{k})}USC({\bf B}^{k})$, where $\Pi_{n}(T^{k})$ is the set of allocations of $T^{k}$.
 4.1 The Impossibility of Approximate EF1
-In this part, we show the following negative result about the approximation of EF1 for tri-valued cost functions.
-For any agent $i\in N$, her cost function $c_{i}$ is additive tri-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
+In this part, we show the following negative result about the approximation of EF1 for additive personalized tri-valued cost functions.
+For any agent $i\in N$, her cost function $c_{i}$ is additive personalized tri-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i},z_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}\leq z_{i}$.
 Theorem 7.
-For general additive cost functions, no deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$, even for two agents with additive tri-valued cost functions.
+No deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$, even for two agents with additive personalized tri-valued cost functions.
 Proof.
-$t=1$ $t=2$ $t=3$
-agent $1$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
-agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
-Table 6: The impossibility result of EF1 for additive tri-valued cost functions.
-Consider an online allocation instance with two agents $\{1,2\}$.
-Assume that, for any agent $i\in\{1,2\}$ and item $e$, $v_{i}(e)\in\{\epsilon,1,\frac{1}{\epsilon}\}$, where $10<\epsilon<$ 1.
-The cost of each arrived item is shown in Table 6.
-When $t=1$, the arriving item has a cost of 1 for both agents.
+$e_{1}$ $e_{2}$ $e_{3}$
+agent 11 11 $\epsilon$ $\frac{1}{\epsilon}$
+agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
+Table 8: The impossibility result of EF1 for additive personalized tri-valued cost functions.
+Let $a_{1}=a_{2}=\epsilon$, $b_{1}=b_{2}=1$, and $z_{1}=z_{2}=\frac{1}{\epsilon}$, where $0<\epsilon<1$.
+Consider the following online instance for $n=2$, where the cost of each item is shown in Table 8.
+When $t=1$, the arriving item has the cost of 1 for both agents.
 Without loss of generality, agent 1 picks it.
@@
 Therefore, for any $\alpha>0$, let $\epsilon<\frac{1}{\alpha}$, and we can conclude that no algorithm can guarantee a complete allocation that is $\alpha$-EF1 for any given $\alpha>0$.
-This impossibility result can be generalized to more than two agents and we defer the proof to Proposition 6.
+This impossibility result can be generalized to more than two agents, and we defer the proof to Proposition 4.
 ∎
+Proposition 4.
+For general additive cost functions, no deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$ for more than two agents with additive personalized tri-valued cost functions.
+Proof.
+$e_{1}$ $e_{2}$ $\ldots$ $e_{n-2}$ $e_{n-1}$ $e_{n}$ $e_{n+1}$
+agent 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
+agent 22 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
+$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
+agent n2n-2 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
+agent n1n-1 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\epsilon$ $\frac{1}{\epsilon}$
+agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\ldots$ $\frac{1}{\epsilon}$ 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
+Table 9: The impossibility result of EF1 for additive personalized tri-valued cost functions when there are more than two agents.
+We extend the impossibility result of approximate EF1 for chores to $n\geq 3$ agents.
+Consider the following instance with $n\geq 3$ agents, where the cost of each item is shown in Table 9.
+Before each agent’s bundle has one agent, we should allocate the arriving item to the agent with an empty bundle to avoid the approximation ratio of EF1 being infinite.
+For $t\leq n-2$, each arriving item has the cost of $\frac{1}{\epsilon}$ for each agent.
+Without loss of generality, assume that agent $i\in[n-2]$ picks the item that arrives at round $i$.
+When t=n1t=n-1, the arriving item has the cost of 1 for all agents.
+Assume that it is allocated to agent n1n-1.
+When $t=n$, the arriving item has the cost of $\frac{1}{\epsilon}$ for agent $n$ and $\epsilon$ for the remaining agents.
+It is clear that agent $n$ picks it.
+In the last round $t=n+1$, the item that has the cost of $\frac{1}{\epsilon}$ for all agents arrives.
+If it is allocated to agent $i\in[n-2]$, we can derive that the approximation ratio of EF1 is $\frac{1}{\epsilon}$.
+If it is allocated to agent n1n-1 or $n$, the allocation is also $\frac{1}{\epsilon}$-EF1.
+Thus, the allocation in this round is $\frac{1}{\epsilon}$-EF1.
+Given any $\alpha>1$, let $\epsilon>\frac{1}{\alpha}$, we can conclude that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha>1$.
+∎
+By the above theorem, no approximation of EF1 can be guaranteed for additive personalized tri-valued cost functions.
+Like the negative result of goods setting, we cannot make further progress in the approximation of EF1 beyond additive personalized tri-valued cost functions if there is no prior information about future items.
+Therefore, in the following parts, we focus on additive binary, supermodular binary, and additive personalized bi-valued cost functions, and study whether there are positive results about the approximation of EF1.
 4.2 Additive/Supermodular Binary Costs
@@
 Input: An instance $(T,N,(c)_{i\in N})$ with additive binary cost functions
-Output: An complete EF1 allocation $A$
-1 Initialize: let $\pi=(\pi_{1},\cdots,\pi_{n})$ be an arbitrary order of $n$ agents;
-2Let $A=(\emptyset,\ldots,\emptyset)$;
+Output: An complete EF1 allocation $\bf A$}
+1 Initialize: let $\pi=(\pi_{1},\ldots,\pi_{n})$ be an arbitrary order of $n$ agents;
+2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
 3 when item $e\in T$ arrives do
@@
 13 $A_{\pi_{1}}=A_{\pi_{1}}\cup\{e\}$;
-14      $\pi=(\pi_{2},\cdots,\pi_{i-1},\pi_{i+1},\cdots,\pi_{n},\pi_{1})$;
+14 $\pi=(\pi_{2},\ldots,\pi_{i-1},\pi_{i+1},\ldots,\pi_{n},\pi_{1})$;
 15
 16
-17return $A$;
+17return $\bf A$};
 Algorithm 8 Compelled-Greedy Algorithm
@@
 Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, where $c_{i}$ is additive and binary for all $i\in N$.
-Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ be the set of items that have already arrived, denoted by ${\bf A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $k$.
-For every item $e\in T$, if there exists one agent such that the cost of $e$ to some agent is 00, the algorithm must allocate $e$ to some agent $a$ with $v_{a}(e)=0$.
+Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by ${\bf A^{k}}=(A^{k}_{1},\ldots,A^{k}_{n})$ the allocation at the end of round $k$.
+For every item $e\in T$, if there exists one agent such that the cost about $e$ of some agent is 0, the algorithm must allocate $e$ to some agent $a$ with $v_{a}(e)=0$.
 Otherwise, the costs of all agents are $1$.
@@
 For supermodular binary cost functions, where for any agent $i\in N$, $c_{i}$ is supermodular and the marginal cost of any item is 0 or 1, we first obtain a significant negative result regarding the approximation of fairness and efficiency.
-This contrasts starkly with the findings in offline settings, where it is possible to efficiently obtain an allocation satisfying EF1, MMS, or min-USC [18].
+This contrasts starkly with the findings in offline settings, where it is possible to efficiently obtain an allocation satisfying EF1, MMS, or min-USC Barman et al.
+(2023).
 Lemma 7.
-(Lemma 2 in [18]) A cost function $c$ is supermodular and binary if and only if the function $g$, $g(S)=|S|-c(S)$ for any set $S$, is a matroid rank function.
+(Lemma 2 in Barman et al. (2023)) A cost function $c$ is supermodular and binary if and only if the function $g$, g(S)=|S|c(S)g(S)=|S|-c(S) for any set $S$, is a matroid rank function.
 Theorem 9.
-For supermodular binary cost functions, no deterministic online algorithms can achieve $\alpha$-EF1 and completeness for any $\alpha\geq 1$.
+For any $\alpha\geq 1$, with supermodular binary cost functions, no deterministic online algorithms can achieve
+  • $\alpha$-EF1 and completeness,
+  • or $\alpha$-MMS and completeness,
+  • or $\alpha$-min-USC and completeness.
 Proof.
 Figure 3: An instance with $\frac{1}{0}$-EF1 allocations
-Consider an online allocation instance, there are two agents $\{1,2\}$ with supermodular binary cost functions $c_{1}$ and $c_{2}$.
+First, let us consider EF1. Consider an online allocation instance; there are two agents $\{1,2\}$ with supermodular binary cost functions $c_{1}$ and $c_{2}$.
 Based on the relationship between the supermodular and binary functions and the mathematical rank function in Lemma 7, we construct cost functions for the instance by constructing a matroid rank function.
-For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $\{C^{i}_{1},\cdots,C^{i}_{l}\}$.
+For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $\{C^{i}_{1},\ldots,C^{i}_{l}\}$.
 In Figure 3, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents.
@@
 Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$.
-For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 00.
+For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 0.
 W.l.o.g., we allocate item $e_{1}$ to agent $1$.
-When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 00.
+When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 0.
 We will discuss the following two cases.
 Case 1: we allocate $e_{2}$ to agent $1$.
-For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00.
+For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0.
 Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases.
 Case 1-1: we allocate $e_{3}$ to agent $1$.
-When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00.
+When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0.
 To not break the requirement of EF1, we must allocate item $e_{4}$ to agent $2$.
-When the fifth item $e_{5}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00.
+When the fifth item $e_{5}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0.
 To avoid breaking the requirement of EF1, we must allocate item $e_{5}$ to agent $2$.
@@
 Case 1-2: we allocate $e_{3}$ to agent $2$.
-When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00.
+When the fourth item $e_{4}$ arrives, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0.
 Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases.
@@
 Case 2: we allocate $e_{2}$ to agent $2$.
-For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 00.
+For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 0.
 Regardless of how it is allocated, it satisfies EF1, so we continue to discuss it in two sub-cases.
@@
 Else, we allocate $e_{5}$ to agent $2$, $A^{5}_{1}=\{e_{1},e_{4}\}$, $A^{5}_{2}=\{e_{2},e_{3},e_{5}\}$, then we have $c_{2}(A^{5}_{1})=0<1=c_{2}(A^{5}_{2}\setminus\{e\})$.
-∎
-Theorem 11.
-For supermodular binary cost functions, no deterministic online algorithms can achieve $\alpha$-MMS or $\alpha$-min-USC, and completeness for any $\alpha\geq 1$.
-Proof.
 Figure 4: An instance with $\frac{1}{0}$-MMS allocations
-Consider an online allocation instance, there are two agents $\{1,2\}$ with supermodular binary cost functions $c_{1},c_{2}$.
-For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $C^{i}_{1},\cdots,C^{i}_{l}$.
+Next, let us consider MMS or USC.
+Consider an online allocation instance; there are two agents $\{1,2\}$ with supermodular binary cost functions $v_{1}v_{2}$.
+For $i\in\{1,2\}$, the online items in $T$ belongs to several categories $C^{i}_{1},\ldots,C^{i}_{l}$.
 In Figure 4, categories are distinguished by colours, and edges of different colours represent items of different categories for adjacent agents.
@@
 Let the function $r_{i}$ be the rank function of matroid $\mathcal{M}_{i}$, the cost function $c_{i}(\cdot)$ is defined as $c_{i}(S)=|S|-r_{i}(S)$ for any item set $S\in T$.
-For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 00.
+For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are 0.
 We allocate item $e_{1}$ to agent $1$.
-When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 00.
+When the second item $e_{2}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 0.
 We will discuss the following two cases.
 Case 1: we allocate $e_{2}$ to agent $1$.
-For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 00.
+For the third item $e_{3}$, the marginal utility for agent $1$ is $1$, and the marginal utility for agent $2$ is 0.
 In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{3}$ to agent $2$.
@@
 Case 2: we allocate $e_{2}$ to agent $2$.
-For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 00.
-Continue to discuss in two sub-cases.
-Case 2-1: We allocate $e_{3}$ to agent $2$.
+For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are 0.
+Continue to discuss in two sub-cases Case 2-1: we allocate $e_{3}$ to agent $2$.
 For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are $1$.
@@
 Case 2-2: we allocate $e_{3}$ to agent $1$.
-For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are 00.
+For the fourth item $e_{4}$, the marginal utilities for agent $1$ and agent $2$ are 0.
 In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{4}$ to agent $2$.
-When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 00.
+When the fifth item $e_{5}$ arrives, the marginal utilities for agent $1$ and agent $2$ are 0.
 In order not to break the requirement of $\mathsf{MMS}_{1}(S)$, we must allocate item $e_{5}$ to agent $2$.
@@
 ∎
-4.3 Additive Bi-valued Costs
-In this part, we consider additive bi-valued cost functions.For any agent $i\in N$, her cost function $c_{i}$ is additive bi-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$.Similar to the results of goods setting, even for two agents, an EF1 or MMS allocation cannot be guaranteed.
+4.3 Additive Personalized Bi-valued Costs
+In this part, we consider additive personalized bi-valued cost functions.
+For any agent $i\in N$, her cost function $c_{i}$ is additive personalized bi-valued if $c_{i}$ is additive and $c_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$.
+Similar to the results of goods setting, even for two agents, an EF1 or MMS allocation cannot be guaranteed.
 4.3.1 Two Agents
-First, we study the additive bi-valued cost functions for two agents and propose the Adapted Chores Envy-Graph Procedures (Algorithm 8) to return an allocation that achieves the tight approximation guarantee for EF1 and has a good approximation for MMS.
-Using techniques similar to the case of goods, we take the chores adapted envy-graphs procedures as the base algorithm, and then design two auxiliary algorithms (Algorithms 9 and 10) to remove the possible envy cycle by complicated case analysis to ensure the approximation of EF1 and MMS.
-Definition 11 (Envy Graph for Chores).
-Given an allocation $A^{k}$, in an envy graph, each vertex corresponds to an agent, and each directed edge $(i,j)$ represents that agent $i$ envies agent $j$, where we have $c_{i}(A_{i}^{k})>c_{i}(A_{j}^{k})$.
-Input: An instance $(T,N,(c_{1},c_{2}))$ with additive bi-valued cost functions
-Output: A complete approximate EF1 and MMS allocation $A$
+First, we study the additive personalized bi-valued cost functions for two agents and propose the Adapted Chores Envy-Graph Procedure (Algorithm 9) to return an allocation that achieves the tight approximation guarantee for EF1 and has a good approximation for MMS.
+Using techniques similar to the case of goods, we take the Chores Envy-Graphs procedure (Algorithm 10) as the base algorithm.
+When it may not be further executed (an envy cycle will appear), we skip to the Chores Preliminary-Breaking-Cycle algorithm (Algorithm 11), where we carefully allocate the future items and guarantee the desired approximation.
+During the execution of the Chores Preliminary-Breaking-Cycle (CPBC) algorithm, in one complicated case ($(\lambda_{1},\mu_{1})=(2,2)$), we introduce the Chores Deep-Breaking-Cycle algorithm (Algorithm 12) to address it.
+The Chores Deep-Breaking-Cycle (CDBC) algorithm may be consecutively executed in many iterations, but the CPBC algorithm is at most consecutively executed in two iterations.
+Note that the CPBC and CDBC algorithms are used to remove the possible envy cycle and guarantee that in some further iteration, we can return to the CEGP algorithm.
+Input: An instance $(T,N,(c_{1},c_{2}))$ with additive personalized bi-valued cost functions
+Output: A complete approximate EF1 and approximate MMS allocation $\bf A$}
 1
-2Let $A=(\emptyset,\ldots,\emptyset)$ and build an envy graph $G$ for two agents;
+2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
+Initialize $\alpha=1$; // Algorithm Indicator
+3
+Initialize $(\lambda_{1},\mu_{1})$ and $(\lambda_{2},\mu_{2})$, where $\lambda_{1}=\lambda_{2}=1$ and $\mu_{1}=\mu_{2}=0$; ;
+  // $\lambda$: the times of consecutive execution for the CPBC or CDBC algorithm, and $\mu$: the case label in the CPBC or CDBC algorithm.
+4
+5Build an envy graph $G$ for two agents;
 6 when item $e\in T$ arrives do
-4
-5      if there is no envy between the two agents then
-6             if $\forall k\in N$, $c_{k}(e)=a_{k}$ or $\forall k\in N$, $c_{k}(e)=b_{k}$  then
-7                   Arbitrarily choose one agent $i\in N$;
+7 if $\alpha=1$ then
+8
+9    $A,\alpha\leftarrow\mathsf{CEGP}(N,A,\alpha,e)$;
+10    if $\alpha=2$ then
+       $A,\alpha\leftarrow\mathsf{CPBC}(N,A,\alpha,e,(\lambda_{1},\mu_{1}))$; ;
+          // Envy cycle will appear.
+11
+12      end if
+13
+14 else if $\alpha=2$ then
+15    $A,\alpha,(\lambda_{1},\mu_{1})\leftarrow\mathsf{CPBC}(N,A,\alpha,e,(\lambda_{1},\mu_{1}))$;
 16 else
-9                   Choosing the agent $i\in N$ who values it at $a_{i}$;
+17    $A,\alpha,(\lambda_{2},\mu_{2})\leftarrow\mathsf{CDBC}(N,A,\alpha,e,(\lambda_{2},\mu_{2}))$;
 18   end if
+19
+20  Update the envy graph $G$;
+21
+return $\bf A$};
+Algorithm 9 (ACEGP) Adapted Chores Envy-Graph Procedure
+Input: $N,{\bf A},\alpha,e$
+Output: A complete allocation $\bf A$} and $\alpha$
+1
+2if  there is no envy between the two agents  then
+3 if $\exists i\in N$, $c_{i}(e)=a_{k}$  then
+4      Choosing the agent $i$ who values it at $a_{i}$ (breaking ties arbitrarily);
+5 else
+6      Arbitrarily choose one agent $i$;
+7   end if
+8
+9else
+10   Let $i\in N$ be an agent who does not envy the other agent and j=2ij=2-i;
+11 if $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ then
+12    $\alpha=2$;
+13    return $\bf A$} and $\alpha$;
+14   end if
+15
+16 end if
+[…]
+return $\bf A$} and $\alpha$;
+Algorithm 10 (CEGP) Chores Envy-Graph Procedure
+Input: $N,{\bf A},\alpha,e,(\lambda_{1},\mu_{1})$
+Output: A complete allocation $\bf A$}, $\alpha$ and $(\lambda_{1},\mu_{1})$
+1
+2if $\lambda_{1}=1$ then
+3
+4 if $|A_{j}|=1$ then
+5
+6    $A_{j}=A_{j}\cup\{e\}$;
+7    $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$;
+8 else
+9
+10    if $c_{i}(e)=b_{i}$ then
 11
+12       $A_{j}=A_{j}\cup\{e\}$;
+13       $\lambda_{1}=2$ and $\mu_{2}=1$;
 14    else
-13             Let $i\in N$ be an agent who does not envy the other agent and $j=2-i$;
-14            if $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ then
-15                   Terminate this algorithm and implement the chores preliminary-breaking-cycle algorithm (Algorithm 9);
-16
+15       $A_{i}=A_{i}\cup\{e\}$;
+16       $\lambda_{1}=2$ and $\mu_{2}=2$;
 17      end if
+18
+19   end if
+20
+21else
+22
+23 if $\mu_{2}=1$  then
+24
 25    $A_{i}=A_{i}\cup\{e\}$;
-19
-return $A$;
-Algorithm 8 Adapted Chores Envy-Graph Procedures
-Input: An instance $(T,N,(c_{1},c_{2}))$ with additive bi-valued cost functions
-Output: A complete approximate EF1 and MMS allocation $A$
+26 else
+27
+28    if $c_{j}(e)=a_{j}$ then
+29
+30       $A_{j}=A_{j}\cup\{e\}$;
+31    else
+32
+33       $A_{i}=A_{i}\cup\{e\}$;
+34       $\alpha=3$, $\lambda_{1}=1$, and $\mu_{2}=0$;
+35      end if
+36
+37   end if
+38
+39 $\alpha=1$, $\lambda_{1}=1$, and $\mu_{2}=0$;
+40 end if
+41
+return $\bf A$}, $\alpha$ and $(\lambda_{1},\mu_{1})$;
+Algorithm 11 (CPBC) Chores Preliminary-Breaking-Cycle
+Input: $N,{\bf A},\alpha,e,(\lambda_{2},\mu_{2})$
+Output: A complete allocation $\bf A$}, $\alpha$, and $(\lambda_{2},\mu_{2})$
 1
-2Initialize: $\lambda=1$ and $\mu=0$;
-3 when item $e\in T$ arrives do
-4
-5      if $\lambda=1$ then
-6
-7            if $|A_{j}|=1$ then
+2if $\lambda_{2}\mod 2=1$ then
+3
+4 if $c_{j}(e)=a_{j}$ then
+5    $A_{j}=A_{j}\cup\{e\}$;
+6    $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$;
+7 else if $c_{j}(e)=b_{j}$ and $c_{i}(e)=b_{i}$ then
 8
 9    $A_{j}=A_{j}\cup\{e\}$;
-10                  Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
+10    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=1$;
 11 else
-12
-13                  if $c_{i}(e)=b_{i}$ then
-14
-15                        $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
-16                  else
-17                         $A_{i}=A_{i}\cup\{e\}$ and $\mu=2$;
+12    $A_{i}=A_{i}\cup\{e\}$;
+13    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=2$;
 14   end if
-19                  $\lambda=2$;
+15
 16 end if
-21
-22      else
-23
-24            if $\mu=1$  then
-25
+17else if $\lambda_{2}\mod 2=0$ then
+18 if  $\mu_{2}=1$  then
+19    if $c_{j}(e)=b_{j}$ then
+20
 21       $A_{i}=A_{i}\cup\{e\}$;
-27            else
-28
-29                  if $c_{j}(e)=a_{j}$ then
-30
-31                        $A_{j}=A_{j}\cup\{e\}$;
-32                  else
-33
-34                        $A_{i}=A_{i}\cup\{e\}$;
-35                        Terminate this algorithm and execute the chores deep-breaking-cycle algorithm (Algorithm 10);
-36                   end if
-37
-38             end if
-39
-40       end if
-41      Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
-42
-return $A$;
-Algorithm 9 Chores Preliminary-Breaking-Cycle Algorithm
-Input: An instance $(T,N,(c_{1},c_{2}))$ with additive bi-valued cost functions
-Output: A complete approximate EF1 and MMS allocation $A$
-1 Initialize: $\lambda=1$ and $\mu=0$;
-2 when item $e\in T$ arrives do
-3
-4      if $\lambda=1$ then
-5
-6            if $c_{j}(e)=a_{j}$ and $c_{i}(e)=b_{i}$ then
-7                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=1$;
-8             else if $c_{j}(e)=b_{j}$ and $c_{i}(e)=b_{i}$ then
-9
-10                  $A_{j}=A_{j}\cup\{e\}$ and $\mu=2$;
-11             else if  $c_{j}(e)=a_{j}$ and $c_{i}(e)=a_{i}$  then
-12                   $A_{j}=A_{j}\cup\{e\}$ and $\mu=3$;
-13             else
-14                   $A_{i}=A_{i}\cup\{e\}$ and $\mu=4$;
-15             end if
-16            $\lambda=2$;
-17      else
-18             if $\mu=1$ then
-19                   $A_{j}=A_{j}\cup\{e\}$;
-20             else if  $\mu=2$  then
-21                   if $c_{j}(e)=b_{j}$ then
-22
-23                        $A_{i}=A_{i}\cup\{e\}$;
-24                        $\lambda=1$;
+22       $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
 23    else
 24       $A_{j}=A_{j}\cup\{e\}$;
+25       $\alpha=1$, $\lambda_{2}=1$ and $\mu_{2}=0$;
 26      end if
-28
-29             else if  $\mu=3$  then
-30                   if $c_{i}(e)=b_{i}$ then
-31
+27
+28 else
+29
 30    $A_{j}=A_{j}\cup\{e\}$;
-33                  else
+31    $\lambda_{2}=\lambda_{2}+1$ and $\mu_{2}=0$;
+32   end if
+33
 34
-35                        $A_{i}=A_{i}\cup\{e\}$;
-36                        $\lambda=1$;
-37                   end if
-38
-39             else
-40
-41                  $A_{j}=A_{j}\cup\{e\}$;
-42                  $\lambda=1$;
-43             end if
-44
-45            if $\lambda=2$ then
-46                   Terminate this algorithm and execute the adapted chores envy-graph procedures (Algorithm 8);
-47
-48       end if
-49
-50
-51return $A$;
-Algorithm 10 Chores Deep-Breaking-Cycle Algorithm
+35return $\bf A$}, $\alpha$, and $(\lambda_{2},\mu_{2})$;
+Algorithm 12 (CDBC) Chores Deep-Breaking-Cycle
+Definition 11 (Envy Graph for Chores).
+Given an allocation ${\bf A}^{k}$, the corresponding envy graph is defined as $G=(V,E)$, where the vertex set $V$ corresponds to agent set $N$, and a directed edge $(i,j)\in E$ iff agent $i$ envies agent $j$, i.e., $c_{i}(A_{i}^{k})>c_{i}(A_{j}^{k})$.
+Additionally, the directed cycle in the envy graph is called an envy cycle.
 Theorem 10.
-For the deterministic allocation of indivisible chores for two agents with additive bi-valued cost functions, the Adapted Chores Envy-Graph Procedures (Algorithm 8) computes a complete allocation that satisfies $2$-EF1 and $\frac{5}{3}$-MMS.
-Similar to the goods setting, we first show the following key observations and lemmas before the proof of Theorem 12.
+For the deterministic allocation of indivisible chores for two agents with additive personalized bi-valued cost functions, the Adapted Chores Envy-Graph Procedures (Algorithm 9) computes a complete allocation that satisfies $2$-EF1 and $\frac{5}{3}$-MMS.
+Before we show the above Theorem, we first list some lemmas used in the proof.
+Lemma 8.
+The CEGP algorithm (Algorithm 10) computes an EF1 allocation without an envy cycle.
+Proof.
+Note that there are three kinds of input allocation.
+The first one is the initial empty allocation, which is EF1 without an envy cycle.
+The second one is the output allocation of the CPBC algorithm.
+The last one is the output allocation of the CDBC algorithm.
+It suffices to show that the last two kinds of input allocation are EF1 without an envy cycle.
+That is because if the input allocation of the CEGP algorithm is EF1 without an envy cycle, by the description of the algorithm and induction, it is easy to check that the property of EF1 without an envy cycle is always maintained.
+By Lemmas 9 and 10, we can derive that when the CPBC or CDBC algorithm outputs $\alpha=1$, implying that the CEGP algorithm is ready to be executed, the output allocation of the CPBC or CDBC algorithm is EF1 without an envy cycle.
+Therefore, our claim holds.
+∎
 Observation 4.
-When Algorithm 9 is executed, for the first arriving item $e$, we have $c_{j}(e)=a_{j}$.
+When the CPBC algorithm (Algorithm 11) is executed, for the first arriving item $e$, we have $c_{j}(e)=a_{j}$.
 Proof.
-When Algorithm 9 is ready to be executed, it means that if the arriving item $e$ is allocated to agent $i$, there will be an envy cycle, i.e, $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ and $c_{i}(A_{i}\cup\{e\})>c_{i}(A_{j})$ For agent $j$, her bundle is EF1.
+When the CPBC algorithm is ready to be executed, it means that if the arriving item $e$ is allocated to agent $i$, there will be an envy cycle, i.e, $c_{j}(A_{j})>c_{j}(A_{i}\cup\{e\})$ and $c_{i}(A_{i}\cup\{e\})>c_{i}(A_{j})$ For agent $j$, her bundle is EF1.
 Suppose that $c_{j}(e)=b_{j}$.
@@
 Observation 5.
-When Algorithm 9 is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
+When the CPBC algorithm (Algorithm 11) is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
 Proof.
@@
 Lemma 9.
-Algorithm 9 computes a complete allocation that is EF1 with or without an envy cycle, or 2-EF1.
+The CPBC algorithm (Algorithm 11) computes a complete allocation satisfying:
+  • 2-EF1,
+  • or EF1 with an envy cycle,
+  • or EF1 without an envy cycle.
 Proof.
-When Algorithm 9 is executed, the input allocation $A=(A_{1},A_{2})$ is EF1.
+The input allocation $\textbf{A}=(A_{1},A_{2})$ of the CPBC algorithm is EF1.
 According to the description of our algorithm, it is executed in at most two rounds.
 Then, let us consider different rounds.
-$\lambda=1$.
+$\lambda_{1}=1$.
 When the first item $e_{1}$ arrives, we have $c_{j}(e_{1})=a_{j}$ by Observation 4.
@@
 For agent $j$, her new bundle is EF1 since $|A_{j}|=1$ and $c_{j}(e_{1})=a_{j}$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
-When $|A_{j}|>1$, let us consider two subcases about the valuation of $e_{1}$.
-  • If $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\mu=1$.
+Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that the CEGP algorithm is ready to be implemented.
+When $|A_{j}|>1$, let us consider two subcases about the cost of $e_{1}$.
+  • If $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\lambda_{2}=2$ and $\mu_{1}=1$.
 For agent $j$, her new bundle is $2$-EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(A_{i})\geq c_{j}(e_{1})=a_{j}$.
@@
 Thus, this allocation is $2$-EF1.
-  • If $c_{j}(e_{1})=a_{i}$, agent $i$ picks it and we set $\mu=2$.
+  • If $c_{j}(e_{1})=a_{i}$, agent $i$ picks it and we set $\lambda_{2}=2$ and $\mu_{2}=2$.
 For agent $i$, her new bundle is EF1 since after removing $e_{1}$ from her bundle, we have $c_{j}(A_{j})\leq c_{j}(A_{i})$.
@@
 Thus, this allocation is EF1 with an envy cycle.
-$\lambda=2$. When the second item $e_{2}$ arrives, let us consider two cases about $\mu$.
-  • If $\mu=1$, agent $i$ picks it.
+In the above cases, the CPBC algorithm continues to be executed.
+$\lambda_{1}=2$. When the second item $e_{2}$ arrives, let us consider two cases about $\mu_{1}$.
+  • If $\mu_{1}=1$, agent $i$ picks it.
 For agent $i$, her new bundle is EF since we have $c_{i}(A_{i}\cup\{e_{2}\})\leq c_{i}(A_{i})+b_{i}\leq c_{i}(A_{j}\cup\{e_{1}\})$.
 For agent $j$, her bundle is EF1 since we have $v_{j}(A_{j}\setminus\{f\})\leq v_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(e_{2})$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
-  • If $\mu=2$, let us consider two subcases about the cost of $e_{2}$.
+Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that the CEGP algorithm is ready to be executed.
+  • If $\mu_{1}=2$, let us consider two subcases about the cost of $e_{2}$.
 When $c_{j}(e_{2})=a_{j}$, agent $j$ picks it.
@@
 For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i})$ for some $f\in A_{j}$ and $c_{j}(e_{2})=a_{j}\leq c_{j}(e_{1})$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
+Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that the CEGP algorithm is ready to be executed.
 When $c_{j}(e_{2})=b_{j}$, agent $i$ picks it.
@@
 For agent $j$, her bundle is EF1 since she does not pick any item in these two rounds.
-Thus, this allocation is 2-EF1, and the algorithm is terminated.
+Thus, this allocation is 2-EF1, and we set $\alpha=3$, $\lambda_{1}=1$, and $\mu_{1}=0$, implying that we need to implement the CDBC algorithm.
 Combining the above cases together, it is not hard to see that the claim holds. ∎
 Observation 6.
-When Algorithm 10 is executed, there are two properties about allocation $A$:
+When the CDBC algorithm (Algorithm 12) is executed, there are two properties about allocation A:
   • $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$;
-  • Let $e_{1}$ and $e_{2}$ denote the two new allocated items in Algorithm 9, we have $c_{i}(A_{i}\setminus\{e\}\leq 2c_{i}(A_{j})$, where $e\in\arg\max_{f\in\{e_{1},e_{2}\}}c_{i}(f)$.
+  • Let $e_{1}$ and $e_{2}$ denote the two new allocated items in the CPBC algorithm, we have $c_{i}(A_{i}\setminus\{e\}\leq 2c_{i}(A_{j})$, where $e\in\arg\max_{f\in\{e_{1},e_{2}\}}c_{i}(f)$.
 Proof.
-When Algorithm 10 is executed, in Algorithm 9 we have $\lambda=2$ and $\mu=2$. By the proof of Lemma 4, it is easy to check the statements are correct. ∎
+When the CDBC algorithm is executed, the CPBC algorithm is terminated for $\lambda_{1}=\mu_{1}=2$. Then, by the proof of Lemma 9, it is easy to check that statements are correct. ∎
 Lemma 10.
-Algorithm 10 computes a complete allocation that is EF1 with or without an envy cycle, or 2-EF1.
+The CDBC algorithm (Algorithm 12) computes a complete allocation satisfying:
+  • 2-EF1,
+  • or EF1 with an envy cycle,
+  • or EF1 without an envy cycle.
 Proof.
-When Algorithm 10 is executed, by Observation 6, the input allocation $A=(A_{1},A_{2})$ is 2-EF1, where $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$ and $c_{i}(A_{i}\setminus\{e\})\leq 2c_{i}(A_{j})$ for some $e\in A_{i}$.
-Let $f_{1}$ and $f_{2}$ denote the two online items that are allocated to agent $j$ in Algorithm 9 and we have $\min\{c_{j}(e_{1}),c_{j}(e_{2})\}=a_{j}$.
-$\lambda=1$. When the item $e_{1}$ arrives at $\lambda=1$, let us consider four cases of agent’s costs functions.
-  • When $c_{j}(e_{1})=a_{j}$ and $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\mu=1$.
+When the CDBC algorithm is executed, by Observation 6, the input allocation $A=(A_{1},A_{2})$ is 2-EF1, where $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$ and $c_{i}(A_{i}\setminus\{e\})\leq 2c_{i}(A_{j})$ for some $e\in A_{i}$.
+Let $f_{1}$ and $f_{2}$ denote the two items that are allocated to agent $j$ in the CPBC algorithm and we have $\min\{c_{j}(e_{1}),c_{j}(e_{2})\}=a_{j}$.
+$\lambda_{2}=1$. When the item $e_{1}$ arrives at $\lambda_{2}=1$, let us consider three cases of cost functions.
+  • When $c_{j}(e_{1})=a_{j}$, agent $j$ picks it and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$.
 For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$.
-For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(f_{1})$.
-Thus, this allocation is EF1 with an envy cycle.
-  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=b_{i}$, agent $j$ picks it and we set $\mu=2$.
+For agent $j$, her new bundle is EF since we have $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$.
+Thus, this allocation is EF1 without an envy cycle, implying that the CEGP algorithm is ready to be executed.
+  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=b_{i}$, agent $j$ picks it.
+Then, we set $\lambda_{2}=1$ and $\mu_{2}=1$.
 For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$.
@@
 Thus, this allocation is EF1 with an envy cycle.
-  • When $c_{j}(e_{1})$ and $c_{i}(e_{1})=a_{i}$, agent $j$ picks it and we set $\mu=3$.
-For agent $i$, her bundle is EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{1})$.
-For agent $j$, her new bundle is 2-EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$ for some $f\in A_{j}$ and $c_{j}(e_{1})=a_{i}\leq c_{j}(f_{1})$.
-Thus, this allocation is EF1 with an envy cycle.
-  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=a_{j}$, agent $i$ picks it and we set $\mu=4$.
+  • When $c_{j}(e_{1})=b_{j}$ and $c_{i}(e_{1})=a_{j}$, agent $i$ picks it.
+Then, we set $\lambda_{2}=2$ and $\mu_{2}=2$.
 For agent $i$, her new bundle is still 2-EF1 since we have $\frac{c_{i}(A_{i}\setminus\{f\}\cup\{e_{1}\})}{c_{i}(A_{j})}\leq\frac{c_{i}(A_{j})+c_{i}(e_{1})}{c_{i}(A_{j})}\leq 2$ for some $f\in\{f_{1},f_{2}\}$ where the inequality follows from the fact that $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}$ and $|A_{j}|\geq 2$.
 Thus, this allocation is 2-EF1.
-$\lambda=2$. When the item $e_{2}$ arrives at $\lambda=2$, let us consider four cases about $\mu$.
-  • When $\mu=1$, agent $j$ picks it.
-For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{2})\leq c_{i}(e_{1})=b_{i}$.
-For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j})+a_{j}\leq c_{j}(A_{i})$, $c_{j}(e_{1})=a_{j}$ and $c_{j}(e_{2})\leq c_{j}(f_{2})=b_{j}$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
-  • When $\mu=2$, we need to consider two subcases.
+It can be seen that in the last two cases, the CDBC algorithm continues to be executed.
+$\lambda_{2}=2$. When the item $e_{2}$ arrives at $\lambda_{2}=2$, let us consider two cases about $\mu_{2}$.
+  • When $\mu_{2}=1$, we need to consider two subcases.
 If $c_{j}(e_{2})=b_{j}$, agent $i$ picks it.
@@
 Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6.
-Additionally, we set $\lambda=1$ and continue executing this algorithm.
+Additionally, we set $\lambda_{2}=1$ and $\mu_{2}=0$, implying that the CDBC algorithm continues to be executed.
 If $c_{j}(e_{2})=a_{j}$, agent $j$ picks it.
@@
 For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{j}(e_{1})=b_{j}=c_{j}(f_{2})$ and $c_{j}(e_{2})=a_{j}=c_{j}(f_{1})$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
-  • When $\mu=3$, let us consider two subcases.
-If $c_{i}(e_{2})=b_{i}$, agent $j$ picks it.
-For agent $i$, her bundle is EF since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(f_{1})=a_{i}=c_{i}(e_{1})$ and $c_{i}(f_{2})=b_{i}=c_{i}(e_{2})$.
-For agent $j$, her new bundle is EF1 since we have $c_{j}(A_{j}\setminus\{f\})\leq c_{j}(A_{i}\setminus\{f_{1},f_{2}\})$, $c_{j}(e_{1})=a_{j}=c_{j}(f_{1})$ and $c_{j}(e_{2})\leq c_{j}(f_{2})=b_{i}$.
-Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
-If $c_{i}(e_{2})=a_{i}$, agent $i$ picks it.
-For agent $i$, her new bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(e_{2})=a_{i}=c_{i}(e_{1})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$.
-For agent $j$, her bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{1})=a_{j}\leq c_{j}(e_{2})$.
-Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6.
-Additionally, we set $\lambda=1$ and continue executing this algorithm.
-  • When $\mu=4$, agent $j$ picks it, For agent $i$, her bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$.
+Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=0$, implying that the CEGP algorithm is ready to be executed.
+  • When $\mu_{2}=1$, agent $j$ picks it, For agent $i$, her bundle is 2-EF1 since we have $c_{i}(A_{i}\setminus\{f_{1},f_{2}\})\leq c_{i}(A_{j})$, $c_{i}(e_{1})=a_{i}\leq c_{i}(e_{2})$ and $c_{i}(f_{1})=a_{i}\leq c_{i}(A_{j})$.
 For agent $j$, her new bundle is EF since we have $c_{j}(A_{j})\leq c_{j}(A_{i})$ and $c_{j}(e_{2})\leq c_{j}(e_{1})=b_{j}$.
 Thus, this allocation is 2-EF1 and satisfies the properties in Observation 6.
-Additionally, we set $\lambda=1$ and continue executing this algorithm.
+Additionally, we set $\lambda_{2}=1$ and $\mu_{2}=0$, implying that the CDBC algorithm continues to be executed.
 Combining the above cases, it can be seen that our statement holds. ∎
 Proof of Theorem 10.
-In Algorithm 8, we follow a similar way of offline envy graph procedures to allocate items.
-The precondition of Algorithm 8 is that there is no envy cycle between two agents.
-Initially, the empty allocation is EF1 without an envy cycle.
-When there is no envy between two agents, we allocate the arriving item to one agent according to different conditions.
-When there exists an envy between two agents, we have two cases.
-One is that we allocate it to the unenvied agent when the envy cycle does not appear.
-The other is that if we allocate it to the unenvied agent, the envy cycle will appear, and we terminate Algorithm 8.
-Similar to the allocation of goods, we need to ensure that there is no envy cycle between two agents.
-Because we can not revoke the previous allocation in the online setting, Algorithms 9 and 10 help us eliminate the possible envy cycle.
-When the allocation $A$ satisfies the precondition and is EF1, we will continue to execute Algorithm 8.
-By the induction, it is not hard to see that the output allocation of Algorithm 8 is always EF1 without an envy cycle.
-Combining Lemmas 4, 5, and the description of Algorithm 8, we can conclude that the output complete allocation is $2$-EF1.
-For MMS, by the implication between EF1 and MMS [22], i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{n\alpha+n-1}{n-1+\alpha}$-MMS, where $n$ is the number of agents, it is easy to see that the output allocation is $\frac{5}{3}$-MMS.
+In the main body of the CAGEP algorithm 9, it can be seen that for different values of the algorithm indicator $\alpha$, we need to choose the corresponding algorithm to execute.
+By Lemmas 8, 9, and 10, it can be concluded that the output allocation is complete and $2$-EF1.
+Regarding MMS, by the implication between EF1 and MMS Sun et al.
+(2021), i.e., $\alpha$-EF1 $\Rightarrow$ $\frac{n\alpha+n-1}{n-1+\alpha}$-MMS, where $n$ is the number of agents, it is easy to see that the output allocation is $\frac{5}{3}$-MMS.
 ∎
+Next, we complement our approximation results by providing the lower bounds below.
 Theorem 11.
-For two agents with additive bi-valued cost functions, no deterministic online algorithm can compute a complete allocation that is $(2-\epsilon)$-EF1 or $(\frac{3}{2}-\epsilon)$-MMS for any $\epsilon>0$.
+For two agents with additive personalized bi-valued cost functions, no deterministic online algorithm can compute a complete allocation that is $(2-\epsilon)$-EF1 or $(\frac{3}{2}-\epsilon)$-MMS for any $\epsilon>0$.
 Proof.
-$t=1$ $t=2$ $t=3$
-agent $1$ $1$ $1$ $\frac{1}{\epsilon}$
-agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
+$e_{1}$ $e_{2}$ $e_{3}$
+agent 11 11 11 $\frac{1}{\epsilon}$
+agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
 Case 1
-$t=1$ $t=2$ $t=3$ $t=4$
-agent $1$ $1$ $1$ $\frac{1}{\epsilon}$ $1$
-agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
+agent 11 11 11 $\frac{1}{\epsilon}$ 11
+agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
 Case 2.2
-$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
-agent $1$ $1$ $1$ $\frac{1}{\epsilon}$ $1$ $\frac{1}{\epsilon}$
-agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
+agent 11 11 11 $\frac{1}{\epsilon}$ 11 $\frac{1}{\epsilon}$
+agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
 Case 2.3
-Table 7: The impossibility results of EF1 for bi-valued cost functions for two agents.
-$t=1$ $t=2$
-agent $1$ $1$ $1$
-agent 2 $1$ $1$
+Table 10: The impossibility results of EF1 for additive personalized bi-valued cost functions for two agents.
+$e_{1}$ $e_{2}$
+agent 11 11 11
+agent 2 11 11
 Case 1
-$t=1$ $t=2$ $t=3$
-agent $1$ $1$ $1$ 2
-agent 2 $1$ $1$ $2$
+$e_{1}$ $e_{2}$ $e_{3}$
+agent 11 11 11 2
+agent 2 11 11 22
 Case 2
-Table 8: The impossibility results of MMS for additive bi-valued cost functions for two agents.
-Consider an online allocation instance with two agents $\{1,2\}$.
-For any agent $i\in\{1,2\}$ and any item $e$, $c_{i}(e)\in\{1,\frac{1}{\epsilon}\}$, where $10<\epsilon<$ 1.
-The cost of each arrived item is shown in Table 9.
-To guarantee the completeness of chores allocation, we should guarantee that each arriving item is allocated to one agent.
-The impossibility of EF1.
-When $t=1$, the arriving item has a cost of 1 for both agents.
+Table 11: The impossibility results of MMS for additive personalized bi-valued cost functions for two agents.
+Let $a_{1}=a_{2}=1$ and $b_{1}=b_{2}=\frac{1}{\epsilon}$ where $0<\epsilon<1$.
+Consider the following online instance for $n=2$, where the cost of each item is shown in Table 12.
+To guarantee the completeness of the chores allocation, we should guarantee that each arriving item is allocated to one agent.
+EF1.
+When $t=1$, the arriving item has the cost of 1 for both agents.
 Without loss of generality, assume that agent 1 picks it.
@@
 Combining the above cases together, we can derive that the lower bound of EF1 is 2.
-The impossibility of MMS.
+MMS.
 The constructed instance is below.
@@
 ∎
+At last, we examine the compatibility between approximate EF1 or MMS and USC, and provide the following negative results.
 Proposition 5.
-For the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation that is $\alpha$-EF1 or $\beta$-MMS, and $\gamma$-min-USC for any $\alpha,\gamma\geq 1$ and $\beta<2$, even for two agents with additive bi-valued cost functions.
+Given any $\alpha,\gamma\geq 1$ and $\beta<2$, for the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation satisfying
+  • $\alpha$-EF1 and $\gamma$-min-USC,
+  • or $\beta$-MMS and $\gamma$-min-USC,
+even for two agents with additive personalized bi-valued cost functions.
 Proof.
-Consider an online allocation instance with two agents $\{1,2\}$, $c_{1}\in\{1,2\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $10<\epsilon<$ 1.
+Without loss of generality, we assume that $c_{1}\in\{1,2\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $0<\epsilon<1$.
 When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives.
@@
 If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+\epsilon}{\epsilon+1}\rightarrow\infty$-min-USC.
-Combining the above cases, it can be seen that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha\geq 1$ or $\beta$-MMS for any $\beta<2$, and $\gamma$-min-USC for any given $\gamma\geq 1$ when two agents have the bi-valued cost functions.
+Combining the above cases, it can be seen that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha\geq 1$ or $\beta$-MMS for any $\beta<2$, and $\gamma$-min-USC for any given $\gamma\geq 1$ when two agents have the additive personalized bi-valued cost functions.
 ∎
 4.3.2 The Combination with Additive Binary Costs
-Next, we consider the setting where one agent (agent $n$) has the additive bi-valued cost function, and the remaining agents have additive binary cost functions and propose the Adapted-Chores-Picking Algorithm (Algorithm 11) to compute a complete allocation that satisfies EF1 and MMS.
+Next, we consider the setting where one agent (agent $n$) has the additive personalized bi-valued cost function, and the remaining agents have additive binary cost functions and propose the Adapted-Chores-Picking Algorithm (Algorithm 13) to compute a complete allocation that satisfies EF1 and MMS.
 The intuition of this algorithm is similar to the goods setting.
@@
 Regarding efficiency, we show that EF1 or MMS is incompatible with USC in this setting.
--Input: An instance $(T,N,(c_{i})_{i\in N})$ with $(n-1)$ additive binary cost functions $(c_{1}.\ldots,c_{n-1})$ and one additive bi-valued cost function $(c_{n})$.
-Output: A complete EF1 and MMS allocation $A$
+Input: An instance $(T,N,(c_{i})_{i\in N})$ with (n1)(n-1) additive binary cost functions $(c_{1}.\ldots,c_{n-1})$ and one additive personalized bi-valued cost function $(c_{n})$.
+Output: A complete EF1 and MMS allocation $\bf A$}
 1
-2Let $A=(\emptyset,\ldots,\emptyset)$;
+2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
 3 when item $e\in T$ arrives do
@@
 12
-13                  if |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\geq|A_{n}|$ then
+13       if $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\geq|A_{n}|$ then
 14          $A_{n}=A_{n}\cup\{e\}$;
@@
 20    else
-21                   if |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\leq|A_{n}|$ then
+21       if $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\leq|A_{n}|$ then
 22
@@
 32
-return $A$;
+return $\bf A$};
 Algorithm 13 Adapted-Chores-Picking Algorithm
-$t=1$ $t=2$ $t=3$ $t=4$ $t=5$ldots
-agent $1$ $1$ 00 $1$ 1 $1$ldots
-agent $2$ $1$ 1 1 $1$ 1 $\ldots$
-agent 3 $a_{3}$ $a_{3}$ $a_{3}$ $b_{3}$ $b_{3}$ldots
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $\ldots$
+agent 11 11 0 11 1 11 $\ldots$
+agent 22 11 1 1 11 1 $\ldots$
+agent 3 $a_{3}$ $a_{3}$ $a_{3}$ $b_{3}$ $b_{3}$ $\ldots$
 Table 12: A simple illustration for three agents of Algorithm 13.
 Theorem 12.
-For the deterministic allocation of indivisible chores for $(n-1)$ agents with additive binary cost functions and one agent with the additive bi-valued cost function, the Adapted-Chores-Picking Algorithm (Algorithm 11) computes a complete allocation that satisfies EF1 and MMS.
+For the deterministic allocation of indivisible chores for (n1)(n-1) agents with additive binary cost functions and one agent with the additive personalized bi-valued cost function, the Adapted-Chores-Picking Algorithm (Algorithm 13) computes a complete allocation that satisfies EF1 and MMS.
 Proof.
-Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 11.
-Firstly, we show that $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ is EF1 by induction.
+Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ is the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 13.
+EF1.
+We show that the computed allocation is EF1 by induction.
 For the base case, the empty allocation is trivially EF1.
@@
   • If $c_{n}(e_{k+1})=a_{n}$, then let us consider two subcases.
-When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\geq|A_{n}|$, agent $n$ picks it.
+When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\geq|A_{n}|$, agent $n$ picks it.
 For agent $i\in[n-1]$, her bundle is still EF1.
-For agent $n$, her bundle is EF1 since she has the priority of picking the item with the cost of $a_{n}$ and the number of items in her bundle is no more than that of items with non-zero costs in agent $i_{min}$’s bundle.
-When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|<|A_{n}|$, agent $i_{min}$ picks it.
+For agent $n$, her bundle is EF1 since she has the priority of picking the item with the cost of $a_{n}$, and the number of items in her bundle is no more than that of items with non-zero costs in agent $i_{min}$’s bundle.
+When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|<|A_{n}|$, agent $i_{min}$ picks it.
 For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1.
-For agent $i_{min}$, her bundle is EF1 since she has the binary cost valuation and the lowest number of items with non-zero costs.
+For agent $i_{min}$, her bundle is EF1 since she has the additive binary cost valuation and the lowest number of items with non-zero costs.
   • If $v_{n}(e_{k+1})=b_{n}$, the let us consider two subcases.
-When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|\leq|A_{n}|$, agent $i_{min}$ picks it.
+When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|\leq|A_{n}|$, agent $i_{min}$ picks it.
 For any agent $i\in N\setminus\{i_{min}\}$, her bundle is still EF1.
-For agent $i_{min}$, her bundle is EF1 since she has the binary cost valuation and the lowest number of items with non-zero costs.
+For agent $i_{min}$, her bundle is EF1 since she has the additive binary cost valuation and the lowest number of items with non-zero costs.
 For any agent $i\in N\setminus\{i_{min}\}$, her bundle remains untouched and is still EF1.
-When |$|\{f~{}|f\in A_{i_{min}}~{}and~{}c_{i_{min}}(f)>0\}|>|A_{n}|$, agent $n$ picks it.
+When $|\{f~|f\in A_{i_{min}}~and~c_{i_{min}}(f)>0\}|>|A_{n}|$, agent $n$ picks it.
 For agent $i\in[n-1]$, her bundle is trivially EF1.
-For agent $n$, her bundle is EF since she has the priority of picking the item with the cost of $a_{n}$, and in this case, the extra one item in another agent’s bundle must have the cost of $b_{n}$ for agent $n$.
-Combining the above cases, it can be seen that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is still EF1.
-Next, we will show that $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ is MMS.
-For arbitrary agent $i\in N$, if her bundle is EF, we are done.
+For agent $n$, her bundle is EF since she has the priority of picking the item with the cost of $a_{n}$, and in this case, the extra item in another agent’s bundle must have the cost of $b_{n}$ for agent $n$.
+Combining the above cases, it can be seen that after allocating item $e_{k+1}$, the allocation $\textbf{A}^{k+1}$ is still EF1.
+MMS.
+Fix an arbitrary agent $i\in N$ and an arbitrary round $k$.
+If her bundle is EF, we are done.
 If not, then let us consider two cases.
-When $i\in[n-1]$, she has a binary valuation function, and by our algorithm description, the number of items with non-zero costs in her bundle is at most one more than that of the other agent’s bundle.
+When $i\in[n-1]$, she has the additive binary valuation function, and by our algorithm description, the number of items with non-zero costs in her bundle is at most one more than that of the other agent’s bundle.
 Thus, it is not hard to see that her bundle is MMS.
 When $i=n$, currently, the number of items in her bundle is one more than that of the other agent’s bundle, and the extra item has the cost of $a_{n}$.
-After removing this item, the cost of her bundle is no more than that of the other agent’s bundle in her perspective, so we can conclude that her bundle is MMS.
+After removing this item, the cost of her bundle is no more than that of the other agent’s bundle from her perspective, so we can conclude that her bundle is MMS.
 ∎
 Proposition 6.
-For the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation that is $\alpha$-EF1 or MMS, and $\beta$-min-USC for any $\alpha,\beta\geq 1$, even for two agents where one agent has the additive binary cost function and the other agent has the additive bi-valued cost function.
+Given any $\alpha,\beta\geq 1$, for the deterministic allocation of indivisible chores, no online algorithm can compute a complete allocation satisfying
+  • $\alpha$-EF1 and $\beta$-min-USC,
+  • or MMS and $\beta$-min-USC,
+even for two agents where one agent has the additive binary cost function and the other agent has the additive personalized bi-valued cost function.
 Proof.
-Without loss of generality, we assume that $c_{1}\in\{0,1\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $10<\epsilon<$ 1.
+Without loss of generality, we assume that $c_{1}\in\{0,1\}$ and $c_{2}\in\{\frac{1}{\epsilon},\frac{1}{\epsilon^{2}}\}$, where $0<\epsilon<1$.
 When $t=1$, the first item $e_{1}$ that has the cost of 1 for agent 1 and $\frac{1}{\epsilon}$ for agent 2 arrives.
@@
 If it is allocated to agent 2, this allocation is EF1, MMS, and $\frac{\frac{1}{\epsilon^{2}}+1}{2}=\frac{1}{2\epsilon^{2}}+\frac{1}{2}\rightarrow\infty$-min-USC.
-Combining the above cases, it can be seen that no online algorithm outputs a complete allocation that is $\alpha$-EF1 or MMS in the setting of binary and bi-valued cost functions, and guarantees $\beta$-min-USC for any given $\alpha,\beta\geq 1$.
+Combining the above cases, it can be seen that no online algorithm outputs a complete allocation that is $\alpha$-EF1 or MMS in the setting of additive binary and additive personalized bi-valued cost functions, and guarantees $\beta$-min-USC for any given $\alpha,\beta\geq 1$.
 ∎
 5 Conclusion and Future Work
-In our paper, we study the fair online allocation of indivisible items and present a series of possible and impossible results regarding the existence and approximation results in fairness and efficiency with various valuation or cost functions, which give a picture of what can be achieved.
-For future work, one interesting direction is to fill the gap of the approximation of EF1 and MMS for bi-valued valuation or cost functions when there are more than three agents.
+In our paper, we study the fair online allocation of indivisible items and present a series of positive and negative results regarding the existence and approximation results in fairness and efficiency with various valuation or cost functions, which provide a comprehensive view of what can be achieved.
+For future work, the most interesting direction is to fill the gap in the approximation of EF1 and MMS for bi-valued valuation or cost functions when there are more than two agents, because it involves the complicated envy cycle, and the current technique cannot be directly generalized.
 Additionally, based on our results in other online settings, it would also be interesting to study the items with deadline setting in a more complicated valuation or cost function setting.
-Appendix A Missing Proofs
-A.1 Section 3
-A.1.1 Proof of Proposition 5
-Proposition 5.
-For general additive functions, no deterministic online algorithm can compute an NW allocation that guarantees $\alpha$-EF1 or $\alpha$-MMS for any $\alpha>0$ when there are more than two agents with additive tri-valued valuation functions.
-Proof.
-$t=1$ $t=2$ldots $t=n-2$ $t=n-1$ $t=n$ $t=n+1$
-agent $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
-agent $2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
-$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
-agent $n-2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
-agent $n-1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
-agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
-More than two agents
-Table 10: The impossibility results of EF1 and MMS for additive tri-valued valuations.
-We further extend the impossibility results of approximate EF1 and MMS for $n\geq 3$.
-Consider the following instance with $n\geq 3$ agents, where the value of each item is shown in Table 10.
-In the first $(n-2)$ rounds, to avoid 0-EF1 or 0-MMS, each item is allocated to the agent with the empty bundle.
-For the item that arrives before round $t=n-2$, it has the value of $\frac{1}{\epsilon}$ for any agent.
-Without loss of generality, we assume that agent $i$ picks the item in the round $t=i$, where $i\leq n$.
-In round $t=n-1$, the arriving item has the value of 1 for any agent, and we assume that it is allocated to agent $n-1$.
-In round $t=n$, the arriving item has the value of $\epsilon$ for agent $i\in N\setminus\{n-1\}$ and $\frac{1}{\epsilon}$ for agent $n-1$, and it should be allocated to agent $n$ to avoid 0-EF1 or 0-MMS.
-Finally, in the round $t=n+1$, the item with the value of $\frac{1}{\epsilon}$ for each agent arrives.
-Due to non-wastefulness, this item must be allocated to one agent.
-If this item is allocated to agent $i\in N\setminus\{n-1,n\}$, it is not hard to see that the allocation is $\epsilon^{2}$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS.
-If this item is allocated to agent $n-1$, this allocation is $\epsilon$-EF1 and $\frac{\epsilon}{1+\epsilon}$-MMS.
-If this item is allocated to agent $n$, we can derive that the competitive ratio of EF1 is $\epsilon$ and MMS is $\epsilon$.
-Combining the above cases together, we can conclude that the upper bound of EF1 is $\epsilon$ and MMS is $\epsilon$.
-For given $\alpha>0$, let $\epsilon<\alpha$, we can conclude that no online algorithm can guarantee an NW allocation that is $\alpha$-EF1 or $\alpha$-MMS for any given $\alpha>0$.
-∎
-A.2 Section 4
-A.2.1 Proposition 6
-Proposition 6.
-For general additive cost functions, no deterministic online algorithm can compute a complete allocation that guarantees $\alpha$-EF1 for any $\alpha\geq 1$ for more than two agents with additive tri-valued cost functions.
-Proof.
-$t=1$ $t=2$ldots $t=n-2$ $t=n-1$ $t=n$ $t=n+1$
-agent $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
-agent $2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
-$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
-agent $n-2$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
-agent $n-1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\epsilon$ $\frac{1}{\epsilon}$
-agent $n$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ldots $\frac{1}{\epsilon}$ $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
-More than two agents
-Table 11: The impossibility result of EF1 for additive tri-valued cost functions.
-We extend the impossibility result of approximate EF1 for chores to $n\geq 3$ agents.
-Consider the following instance with $n\geq 3$ agents, where the cost of each item is shown in Table 11.
-Before each agent’s bundle has one agent, we should allocate the arriving item to the agent with an empty bundle to avoid the approximation ratio of EF1 being infinite.
-For $t\leq n-2$, each arriving item has the cost of $\frac{1}{\epsilon}$ for each agent.
-Without loss of generality, assume that agent $i\in[n-2]$ picks the item that arrives at round $i$.
-When $t=n-1$, the arriving item has the cost of 1 for all agents.
-Assume that it is allocated to agent $n-1$.
-When $t=n$, the arriving item has the cost of $\frac{1}{\epsilon}$ for agent $n$ and $\epsilon$ for the remaining agents.
-It is clear that agent $n$ picks it.
-In the last round $t=n+1$, the item that has the cost of $\frac{1}{\epsilon}$ for all agents arrives.
-If it is allocated to agent $i\in[n-2]$, we can derive that the approximation ratio of EF1 is $\frac{1}{\epsilon}$.
-If it is allocated to agent $n-1$ or $n$, the allocation is also $\frac{1}{\epsilon}$-EF1.
-Thus, the allocation in this round is $\frac{1}{\epsilon}$-EF1.
-Given any $\alpha>1$, let $\epsilon>\frac{1}{\alpha}$, we can conclude that no online algorithm can compute a complete allocation that is $\alpha$-EF1 for any $\alpha>1$.
-∎
-Appendix B Other Online Settings
+Appendix A Other Online Settings
 Online algorithms face two roadblocks: unknown future information and the inability to revoke previous decisions.
@@
 In this section, we study two special settings.
-One is that each arriving item can wait for one period and does not need to be allocated immediately, which can be found in the online matching literature [23].
-In this setting, we prove that an EF1 allocation for two agents with bi-valued valuation/cost functions exists.
-The other one is knowing some information in advance, which originates from the online MMS allocation with normalized valuations [5].
+One is that each arriving item can wait for one period and does not need to be allocated immediately, which can be found in the online matching literature Ashlagi et al.
+[2018].
+In this setting, we prove that an EF1 allocation for two agents with additive personalized bi-valued valuation/cost functions exists.
+The other one is knowing some information in advance, which originates from the online MMS allocation with normalized valuations Zhou et al.
+[2023].
 In this setting, we show that the Marginal-Greedy Algorithm computes a non-wasteful EF1 allocation when items are goods, and the Round-Robin algorithm computes a complete EF1 allocation when items are chores.
-B.1 Items with Deadline
+A.1 Items with Deadline
 In this part, each item has a deadline $d$ after its arrival.
 After $d$ units of time, we must decide how to allocate it.
-For two agents with bi-valued valuation functions in the goods setting, we design the Adapted Matching Algorithm (Algorithm 12) to compute an NW allocation that satisfies EF1.
-For two agents with bi-valued cost functions in the chores setting, we design the Adapted Chores Matching Algorithm (Algorithm 13) to compute a complete allocation that satisfies EF1.
+For two agents with additive personalized bi-valued valuation functions in the goods setting, we design the Adapted Matching Algorithm (Algorithm 14) to compute an NW allocation that satisfies EF1.
+For two agents with additive personalized bi-valued cost functions in the chores setting, we design the Adapted Chores Matching Algorithm (Algorithm 15) to compute a complete allocation that satisfies EF1.
 The common intuition of the above two algorithms is that we find a perfect matching between two agents and two items after two consecutive items arrive so that we can guarantee that there is no envy cycle between two agents and EF1 always holds.
 Theorem 13.
-For the deterministic allocation of indivisible goods for two agents with bi-valued valuation functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Matching Algorithm (Algorithm 12) computes an NW allocation that satisfies EF1.
+For the deterministic allocation of indivisible goods for two agents with additive personalized bi-valued valuation functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Matching Algorithm (Algorithm 14) computes an NW allocation that satisfies EF1.
 Lemma 11.
@@
 Proof of Theorem 13.
-Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 12.
+Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$is the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 14.
+EF1.
 We will show that the returned allocation A is always EF1 by induction.
@@
   • $\lambda=1$ and item $e_{k+1}$ is the last arriving item.
-When envy exists between two agents, assume that agent $i$ is the agent who envies agent $j=2-i$.
+When envy exists between two agents, assume that agent $i$ is the agent who envies agent j=2ij=2-i.
 If she picks $e_{k+1}$, we are done.
@@
   • $\lambda=2$ and item $e_{k+1}$ needs to be allocated firstly.
-When envy exists between two agents, assume that agent $i$ is the agent who envies agent $j=2-i$.
+When envy exists between two agents, assume that agent $i$ is the agent who envies agent j=2ij=2-i.
 If agent $i$ picks $e_{k+1}$, we are done.
@@
   • $\lambda=2$ and item $e_{k+1}$ needs to be allocated secondly.
-When envy exists between two agents, assume that agent $i$ is the agent who envies agent $j=2-i$.
+When envy exists between two agents, assume that agent $i$ is the agent who envies agent j=2ij=2-i.
 If she picks it, we are done.
@@
 When no envy exists between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
-Combining the above cases together, we can conclude that the allocation $A^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
+Combining the above case, we can conclude that the allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
 Theorem 14.
-For the deterministic allocation of indivisible chores for two agents with bi-valued cost functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Chores Matching algorithm (Algorithm 13) computes a complete allocation that satisfies EF1.
+For the deterministic allocation of indivisible chores for two agents with additive personalized bi-valued cost functions, if we allow each item to stay $d=1$ period after its arrival, the Adapted Chores Matching algorithm (Algorithm 15) computes a complete allocation that satisfies EF1.
 Lemma 12.
@@
   • $c_{i}(A_{i})<c_{i}(A_{j})$.
-After allocating two items $e_{k+1}$ and $e_{k+2}$, the two properties hold since agent $i$ only picks one item and has a bi-valued cost function i.e., $c_{i}(e_{k+1}),c_{i}(e_{k+2})\in\{a_{i},b_{i}\}$.
+After allocating two items $e_{k+1}$ and $e_{k+2}$, the two properties hold since agent $i$ only picks one item and has a additive personalized bi-valued cost function i.e., $c_{i}(e_{k+1}),c_{i}(e_{k+2})\in\{a_{i},b_{i}\}$.
   • $c_{i}(A_{i})>c_{i}(A_{j})$.
@@
 Proof of Theorem 14.
-Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$is the set of items that have already arrived, denoted by $A^{k}=(A^{k}_{1},\cdots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 13.
+Given an online fair allocation instance $I=(T,N,(c_{i})_{i\in N})$, Fix an arbitrary round $k$, let $T=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by $\textbf{A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$ the allocation at the end of round $t$ in Algorithm 15.
+EF1.
 We will show that the output allocation A is always EF1 by induction.
 For the base cases, the empty bundle is EF1 without an envy cycle.
-For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $A^{t}$ is EF1 without an envy cycle.
-Next, we show that after allocating item $e_{k+1}$, the allocation $A^{k+1}$ is EF1 without an envy cycle.
+For the induction step, after allocating item $e_{t}$ for any $t\in[k]$, the allocation $\textbf{A}^{t}$ is EF1 without an envy cycle.
+Next, we show that after allocating item $e_{k+1}$, the allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle.
 Let us consider three cases about item $e_{k+1}$.
   • $\lambda=1$ and item $e_{k+1}$ is the last arriving item.
-When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent $j=2-i$.
+When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent j=2ij=2-i.
 If she picks it, we are done.
@@
 For agent $j$, by Lemma 12, we have $c_{j}(A_{j}^{k})-c_{j}(A_{i}^{k})=b_{j}-a_{j}$.
-Then, we have $c_{j}(A_{j}^{k}\setminus\{f\}\cup\{e_{k+1}\})-c_{j}(A_{i})=b_{j}-a_{j}-b_{j}+a_{j}=0$ where $A_{j}^{k}$ is EF1 and there must exist an item $f$ such that $c_{j}(f)=b_{j}$, implying that allocation $A^{k+1}$ is EF1 without an envy cycle.
+Then, we have $c_{j}(A_{j}^{k}\setminus\{f\}\cup\{e_{k+1}\})-c_{j}(A_{i})=b_{j}-a_{j}-b_{j}+a_{j}=0$ where $A_{j}^{k}$ is EF1 and there must exist an item $f$ such that $c_{j}(f)=b_{j}$, implying that allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle.
 When there is no envy between two agents, allocating $e_{k+1}$ to any agent will not violate EF1.
   • $\lambda=2$ and item $e_{k+1}$ needs to allocated firstly.
-When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent $j=2-i$.
+When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent j=2ij=2-i.
 If agent $i$ picks it, we are done.
@@
   • $\lambda=2$ and item $e_{k+2}$ needs to be allocated secondly.
-When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent $j=2-i$.
+When envy exists between two agents, assume that agent $i$ is the agent who does not envy agent j=2ij=2-i.
 If agent $i$ picks it, we are done.
@@
 When there is no envy between two agents, no matter which agent picks $e_{k+1}$, the new allocation is EF1 without an envy cycle.
-Combining the above cases together, we can conclude that the allocation $A^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
-B.2 Partial Information in Advance
-In this part, agents’ preferences over items are the same, and the arriving order of items is known.
-With this information, we can obtain a non-wasteful EF1 allocation by the Marginal-Greedy (Algorithm 2) in the goods setting and a complete EF1 allocation by the Round-Robin algorithm in the chores setting.
-Definition 12.
-An online goods/chores instance is monotone, if for any agent $i\in N$, we have $f_{i}(e_{1})\geq\cdots\geq f_{i}(e_{T})$ or for any agent $i\in N$, we have $f_{i}(e_{1})\leq\cdots\leq f_{i}(e_{T})$, where $f_{i}\in\{v_{i},c_{i}\}$
-Theorem 17.
-For the deterministic allocation of indivisible goods from a monotone instance, the Marginal-Greedy Algorithm (Algorithm 2) computes an NW allocation that is EF1.
-Proof.
-Without loss of generality, we assume that the picking sequence is $1,\ldots,n$.
-First, let us consider the case of $v_{i}(e_{1})\leq\ldots\leq v_{i}(e_{T})$ for any $i\in N$.
-Note that, for any agent $i\in N$, if she values some items at zero, these items will arrive before the items with non-zero values.
-In this phase, she does not pick anything until the item with non-zero values arrives.
-Additionally, when these zero items are allocated to other agents, from the perspective of agent $i$, the value of the other agent’s bundle remains unchanged.
-Thus, without loss of generality, we assume that $v_{i}(e_{1})\neq 0$ for any agent $i\in N$.
-Fix an arbitrary round $k$ and an arbitrary agent $i\in N$.
-Let $A^{k}$ denote the allocation after the allocation of $e_{k}$.
-Suppose that she picks item $e_{k}$.
-For agent $i$, we have $v_{i}(A^{k}_{i})\geq v_{i}(A^{k}_{j}\setminus\{e\})$, where $e$ is the last item added to agent $j$’s bundle, since for any $j<i$, agent $i$ is the latter one to pick the item and for any $j>i$, $v_{i}(e_{s})\geq v_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and $s-t=n-i+j$.
-For agent $j\neq i$, we have $v_{j}(A_{j}^{k})\geq v_{j}(A_{i}^{k}\setminus\{e\})$, where $e$ is the last item added to agent $i$’s bundle, since for any $j>i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and $s-t=j-i$, and $j<i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and $s-t=n-j+i$.
-Thus, the allocation $A^{k}$ is EF1.
-As for the case of $v_{i}(e_{1})\geq\ldots\geq v_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it.
-∎
-Theorem 18.
-For the deterministic allocation of indivisible chores from a monotone instance, the Round-Robin Algorithm computes a complete allocation that is EF1.
-Proof.
-Without loss of generality, we assume that the picking sequence is $1,\ldots,n$.
-First, let us consider the case of $c_{i}(e_{1})\leq\ldots\leq c_{i}(e_{T})$ for any $i\in N$.
-Fix an arbitrary round $k$ and an arbitrary agent $i\in N$.
-Let $A^{k}$ denote the allocation after the allocation of $e_{k}$.
-Suppose that she picks item $e_{k}$.
-For agent $i$, we have $c_{i}(A^{k}_{i}\setminus\{e_{k}\})=c_{i}(A^{k-1}_{i})\leq c_{i}(A^{k-1}_{j})=c_{i}(A^{k}_{j})$ since for any $j>i$, agent $i$ has the priority of picking the item and for any $j<i$, $c_{i}(e_{s})\leq c_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and $t-s=n-i+j$.
-For agent $j\neq i$, we have $c_{j}(A_{j}^{k})\leq c_{j}(A_{i}^{k})$ for any $j<i$ agent $j$ has the priority of picking items and $j>i$, $c_{j}(e_{s})\leq c_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and $t-s=n-j+i$.
-Therefore, the allocation $A^{k}$ is EF1.
-As for the case of $c_{i}(e_{1})\geq\ldots\geq c_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it.∎
-In the monotone instance, each agent has the same preference over all items, and the arriving order of items follows the nonincreasing/nondecreasing order in terms of value or cost.
-Then, if we allow the arriving order to be arbitrary and keep the same preference, we have the following proposition.
-Proposition 7.
-For an online goods/chores instance where agents only have identical preferences, there exists an instance for which no EF1 allocation can be found, even for two agents with bi-valued valuation or cost functions.
-Proof.
-$t=1$ $t=2$ $t=3$ $t=4$
-agent $1$ $1$ $1$ $\epsilon$ $1$
-agent 2 $1$ $\epsilon$ $\epsilon$ $1$
-Case 1
-$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
-agent $1$ $1$ $1$ $\epsilon$ $1$ $1$
-agent 2 $1$ $\epsilon$ 1 $\epsilon$ $1$
-Case 2
-Table 12: The nonexistence of EF1 for two agents that have the same preference over goods.
-$t=1$ $t=2$ $t=3$
-agent $1$ $1$ $1$ $\frac{1}{\epsilon}$
-agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
-Case 1
-$t=1$ $t=2$ $t=3$ $t=4$ $t=5$
-agent $1$ $1$ $1$ $\frac{1}{\epsilon}$ 1 $\frac{1}{\epsilon}$
-agent 2 $1$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
-Case 2
-Table 13: The nonexistence of EF1 for two agents that have the same preference over chores.
-First, we consider the online goods instance.
-Without loss of generality, we assume that $v_{1},v_{2}\in\{\epsilon,1\}$, where $10<\epsilon<$ 1 The value of each item is shown in Table 12.
-When $t=1$, the arriving item has the value of $1$ for both agents.
-Without loss of generality, it is allocated to agent 1.
-For the second round $t=2$, the item that has the value of $1$ for agent 1 and $\epsilon$ for agent 2.
-To guarantee EF1, agent 2 picks it.
-In round $t=3$, the arriving item has the value of $\epsilon$ for both agents.
-Then, let us consider two cases related to allocating it.
-  • If agent 2 picks it, this allocation is still EF1.
-In the last round $t=4$, the arriving item has the value of $1$ for both agents.
-In this case, it is not hard to check that no matter which agent picks it, the resulting allocation is not EF1.
-  • If agent 1 picks it, this allocation is still EF1.
-Then, in round $t=4$, the arriving item has the value of $1$ for agent 1 and $\epsilon$ for agent 2.
-To guarantee that EF1 is not violated, it should be allocated to agent 2.
-In this last round $t=5$, the item has the value of $1$ for both agents.
-In this case, it is trivial that as long as one agent picks it, the final allocation is not EF1.
-Next, let us consider the online chores instance.
-We assume that $c_{1},c_{2}\in\{1,\frac{1}{\epsilon}\}$ where $10<\epsilon<$ 1.
-The cost of each item is shown in Table 13.
-When $t=1$, the arriving item has the cost of 1 for both agents, and we assume that agent 1 picks it.
-In round $t=2$, the item that has the cost of 1 for agent $1$ and $\frac{1}{\epsilon}$ for agent 2.
-To maintain EF1, agent 2 picks it.
-For the next round $t=3$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents.
-Then, let us consider two cases related to allocating it.
-  • If it is allocated to agent 2, no further items appear. In this case, it is clear that this allocation is not EF1.
-  • If it is allocated to agent 1, in the following round $t=4$, the arriving item has the cost of 1 for agent $1$ and $\frac{1}{\epsilon}$ for agent 2.
-It should be allocated to agent 2 to ensure that the property of EF1 still holds.
-In the last round $t=5$, this arriving item has the cost of $\frac{1}{\epsilon}$ for both agents.
-In this case, it is easy to verify that no allocation is EF1 when it is allocated to one agent.
-Combining the above cases, we can conclude that our statement holds. ∎
-Input: An instance $(T,N,(v_{1},v_{2}))$ with bi-valued valuation functions and deadline $d=1$
-Output: An non-wasteful EF1 allocation $A$
+Combining the above cases, we can conclude that the allocation $\textbf{A}^{k+1}$ is EF1 without an envy cycle. This completes the induction and establishes the correctness of our claim. ∎
+Input: An instance $(T,N,(v_{1},v_{2}))$ with additive personalized bi-valued valuation functions and deadline $d=1$
+Output: An NW EF1 allocation A
 1
@@
 14
-15                  Let $i\in N$ be the agent who envies the other one and $j=2-i$;
+15        Let $i\in N$ be the agent who envies the other one and j=2ij=2-i;
 16       if  $v_{i}(e_{1})=b_{i}$  then
@@
 52
-53            Let $i\in N$ be the agent who envies the other one and $j=2-i$;
+53     Let $i\in N$ be the agent who envies the other one and j=2ij=2-i;
 54    if $v_{i}(e_{1})=b_{i}$ then
@@
 Algorithm 14 Adapted Matching Algorithm
-Input: An instance $(T,N,(c_{1},c_{2}))$ with bi-valued cost functions and deadline $d=1$
+Input: An instance $(T,N,(c_{1},c_{2}))$ with additive personalized bi-valued cost functions and deadline $d=1$
 Output: A complete EF1 allocation A
@@
 14
-15                  Let $i\in N$ be the agent who does not envy the other one, and $j=2-i$;
+15        Let $i\in N$ be the agent who does not envy the other one, and j=2ij=2-i;
 16       if  $c_{j}(e_{1})=b_{j}$  then
@@
 50     Allocate item $e_{1}$ to an arbitrarily agent;
-51       end if
+51 else
 52
-53      Let $i\in N$ be the agent who does not envy the other one and $j=2-i$;
+53     Let $i\in N$ be the agent who does not envy the other one and j=2ij=2-i;
 54    if $c_{j}(e_{1})=b_{j}$ then
@@
 59
-60
+60   end if
+61
+62
 return A;
 Algorithm 15 Adapted Chores Matching Algorithm
+A.2 Partial Information in Advance
+In this part, agents’ preferences over items are the same, and the arriving order of items is known.
+With this information, we can obtain a non-wasteful EF1 allocation by the Marginal-Greedy (Algorithm 2) in the goods setting and a complete EF1 allocation by the Round-Robin algorithm in the chores setting.
+Definition 12.
+An online goods/chores instance is monotone if for any agent $i\in N$, we have $f_{i}(e_{1})\geq\cdots\geq f_{i}(e_{T})$ or for any agent $i\in N$, we have $f_{i}(e_{1})\leq\ldots\leq f_{i}(e_{T})$, where $f_{i}\in\{v_{i},c_{i}\}$ .
+Theorem 15.
+For the deterministic allocation of indivisible goods from a monotone instance, the Marginal-Greedy Algorithm (Algorithm 2) computes an NW allocation that is EF1.
+Proof.
+Without loss of generality, we assume that the picking sequence is $1,\ldots,n$.
+First, let us consider the case of $v_{i}(e_{1})\leq\ldots\leq v_{i}(e_{T})$ for any $i\in N$.
+Note that, for any agent $i\in N$, if she values some items at zero, these items will arrive before the items with non-zero values.
+In this phase, she does not pick anything until the item with non-zero values arrives.
+Additionally, when these zero items are allocated to other agents, from the perspective of agent $i$, the value of the other agent’s bundle remains unchanged.
+Thus, without loss of generality, we assume that $v_{i}(e_{1})\neq 0$ for any agent $i\in N$.
+Fix an arbitrary round $k$ and an arbitrary agent $i\in N$.
+Let $\textbf{A}^{k}$ denote the allocation after the allocation of $e_{k}$.
+Suppose that she picks item $e_{k}$.
+For agent $i$, we have $v_{i}(A^{k}_{i})\geq v_{i}(A^{k}_{j}\setminus\{e\})$, where $e$ is the last item added to agent $j$’s bundle, since for any $j<i$, agent $i$ is the latter one to pick the item and for any $j>i$, $v_{i}(e_{s})\geq v_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and st=ni+js-t=n-i+j.
+For agent $j\neq i$, we have $v_{j}(A_{j}^{k})\geq v_{j}(A_{i}^{k}\setminus\{e\})$, where $e$ is the last item added to agent $i$’s bundle, since for any $j>i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and st=jis-t=j-i, and $j<i$, $v_{j}(e_{s})\geq v_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and st=nj+is-t=n-j+i.
+Thus, the allocation $A^{k}$ is EF1.
+As for the case of $v_{i}(e_{1})\geq\ldots\geq v_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it.
+∎
+Theorem 16.
+For the deterministic allocation of indivisible chores from a monotone instance, the Round-Robin Algorithm computes a complete allocation that is EF1.
+Proof.
+Without loss of generality, we assume that the picking sequence is $1,\ldots,n$.
+First, let us consider the case of $c_{i}(e_{1})\leq\ldots\leq c_{i}(e_{T})$ for any $i\in N$.
+Fix an arbitrary round $k$ and an arbitrary agent $i\in N$.
+Let $\textbf{A}^{k}$ denote the allocation after the allocation of $e_{k}$.
+Suppose that she picks item $e_{k}$.
+For agent $i$, we have $c_{i}(A^{k}_{i}\setminus\{e_{k}\})=c_{i}(A^{k-1}_{i})\leq c_{i}(A^{k-1}_{j})=c_{i}(A^{k}_{j})$ since for any $j>i$, agent $i$ has the priority of picking the item and for any $j<i$, $c_{i}(e_{s})\leq c_{i}(e_{t})$ holds, where agent $i$ picks $e_{s}$, agent $j$ picks $e_{t}$ and ts=ni+jt-s=n-i+j.
+For agent $j\neq i$, we have $c_{j}(A_{j}^{k})\leq c_{j}(A_{i}^{k})$ for any $j<i$ agent $j$ has the priority of picking items and $j>i$, $c_{j}(e_{s})\leq c_{j}(e_{t})$ holds, where agent $j$ picks $e_{s}$, agent $i$ picks $e_{t}$ and ts=nj+it-s=n-j+i.
+Therefore, the allocation $\textbf{A}^{k}$ is EF1.
+As for the case of $c_{i}(e_{1})\geq\ldots\geq c_{i}(e_{T})$ for any $i\in N$, we can use a similar analysis to show that the returned allocation is EF1 and we omit it.∎
+In the monotone instance, each agent has the same preference over all items, and the arriving order of items follows the nonincreasing/nondecreasing order in terms of value or cost.
+Then, if we allow the arriving order to be arbitrary and keep the same preference, we have the following proposition.
+Proposition 7.
+For an online goods/chores instance where agents only have identical preferences, there exists an instance for which no EF1 allocation can be found, even for two agents with additive personalized bi-valued valuation or cost functions.
+Proof.
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$
+agent 11 11 11 $\epsilon$ 11
+agent 2 11 $\epsilon$ $\epsilon$ 11
+Case 1
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
+agent 11 11 11 $\epsilon$ 11 11
+agent 2 11 $\epsilon$ 1 $\epsilon$ 11
+Case 2
+Table 13: The nonexistence of EF1 for two agents that have the same preference over goods.
+$e_{1}$ $e_{2}$ $e_{3}$
+agent 11 11 11 $\frac{1}{\epsilon}$
+agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
+Case 1
+$e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$
+agent 11 11 11 $\frac{1}{\epsilon}$ 1 $\frac{1}{\epsilon}$
+agent 2 11 $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$ $\frac{1}{\epsilon}$
+Case 2
+Table 14: The nonexistence of EF1 for two agents that have the same preference over chores.
+First, we consider the online goods instance.
+Without loss of generality, we assume that $v_{1},v_{2}\in\{\epsilon,1\}$, where $0<\epsilon<1$ The value of each item is shown in Table 13.
+When $t=1$, the arriving item has the value of $1$ for both agents.
+Without loss of generality, it is allocated to agent 1.
+For the second round $t=2$, the item that has the value of $1$ for agent 1 and $\epsilon$ for agent 2.
+To guarantee EF1, agent 2 picks it.
+In round $t=3$, the arriving item has the value of $\epsilon$ for both agents.
+Then, let us consider two cases related to allocating it.
+  • If agent 2 picks it, this allocation is still EF1.
+In the last round $t=4$, the arriving item has the value of $1$ for both agents.
+In this case, it is not hard to check that no matter which agent picks it, the resulting allocation is not EF1.
+  • If agent 1 picks it, this allocation is still EF1.
+Then, in round $t=4$, the arriving item has the value of $1$ for agent 1 and $\epsilon$ for agent 2.
+To guarantee that EF1 is not violated, it should be allocated to agent 2.
+In this last round $t=5$, the item has the value of $1$ for both agents.
+In this case, it is trivial that as long as one agent picks it, the final allocation is not EF1.
+Next, let us consider the online chores instance.
+We assume that $c_{1},c_{2}\in\{1,\frac{1}{\epsilon}\}$ where $0<\epsilon<1$.
+The cost of each item is shown in Table 14.
+When $t=1$, the arriving item has the cost of 1 for both agents, and we assume that agent 1 picks it.
+In round $t=2$, the item that has the cost of 1 for agent $1$ and $\frac{1}{\epsilon}$ for agent 2.
+To maintain EF1, agent 2 picks it.
+For the next round $t=3$, the arriving item has the cost of $\frac{1}{\epsilon}$ for both agents.
+Then, let us consider two cases related to allocating it.
+  • If it is allocated to agent 2, no further items appear. In this case, it is clear that this allocation is not EF1.
+  • If it is allocated to agent 1, in the following round $t=4$, the arriving item has the cost of 1 for agent $1$ and $\frac{1}{\epsilon}$ for agent 2.
+It should be allocated to agent 2 to ensure that the property of EF1 still holds.
+In the last round $t=5$, this arriving item has the cost of $\frac{1}{\epsilon}$ for both agents.
+In this case, it is easy to verify that no allocation is EF1 when it is allocated to one agent.
+Combining the above cases, we can conclude that our statement holds. ∎